### v2

# On algebraic properties of some non-prime ideals of collections of cells

Abstract.
In this work, we consider a family of coordinate rings obtained as a quotient ring of a binomial ideal associated to a collection of cells. The algebraic properties of Cohen-Macaulayness, Gorensteiness and the Hilbert-Poincaré series have been studied in literature only for some classes of collections of cells whose related binomial ideals are prime. We study these algebraic properties for some classes of collections of cells whose coordinate rings are not domain, considering a new strategy involving the combinatorics of the collections of cells and a recent result of Conca and Varbaro on the radical initial ideals.
Key words and phrases:
Polyominoes, Cohen-Macaulay, Gorenstein, zig-zag walk, Hilbert series, rook-polynomial.
2010 Mathematics Subject Classification:
05B50, 05E40
Introduction
Combinatorial Commutative Algebra has been an intensive area of research since the pioneering work of R. Stanley ([34]). In 2012 a new topic emerged in this field by the work of A. A. Qureshi. In [28] she establishes a connection between collections of cells and Commutative Algebra, assigning to every collection of cells $\mathcal{P}$ the ideal of the inner 2-minors of $\mathcal{P}$, in a suitable polynomial ring $S_{\mathcal{P}}$. This ideal $I_{\mathcal{P}}$ is called the inner 2-minor ideal of $\mathcal{P}$ and $K[\mathcal{P}]=S_{\mathcal{P}}/I_{\mathcal{P}}$ is said the coordinate ring of $\mathcal{P}$. In particular, if $\mathcal{P}$ is a polyomino, that is a collection of cells where the cells are joined edge by edge, then $I_{\mathcal{P}}$ is called the polyomino ideal of $\mathcal{P}$. The aim of the research consists in the studying of the main algebraic properties of $K[\mathcal{P}]$ depending on the shape of $\mathcal{P}$. Since 2012, many algebraic properties of the polyomino ideals have been studied and many open questions have arisen. For example, one of the most interesting is the classification of the polyominoes whose coordinate ring is a domain. As usual, we say that a collection of cells is prime if the related ideal is prime. In [18, 19, 30], it is proved that if $\mathcal{P}$ is simple, which means $\mathcal{P}$ has no hole, then $K[\mathcal{P}]$ is a domain. Nowadays, the study is devoted to non-simple polyominoes and, although there are several articles where it is discussed, a complete characterization is still unknown. We invite the reader to consult [3, 5, 18, 19, 21, 25, 26, 30, 33]. Also, many other algebraic properties, like when $K[\mathcal{P}]$ is Cohen-Macaulay or Gorenstein, are known only for some specific polyominoes and they are discussed just for prime classes of collections of cells so far. Regarding the Cohen-Macaulayness, the demonstrative strategy is based on two well-known theorems of Sturmfels ([36, Chapter 13]) and Hochster ([2, Theorem 6.3.5]), which state that a toric ring whose defining ideal admits a squarefree initial ideal is a normal Cohen-Macaulay ring. See for instance [6, 21, 26, 30]. No class of collections of cells or polyominoes, whose coordinate ring is a Cohen-Macaulay ring but not a domain, has been discussed. Motivated by these considerations, we investigate the Cohen-Macaulayness and the Hilbert-Poincaré series of the zig-zag collections, a class of non-prime collections of cells coming from the so-called zig-zag walks (see [25, Definition 3.2]), and of the closed paths, a kind of non-simple polyominoes introduced in [3]. Taking advantage of some recent results shown in [9], we provide a demonstrative technique involving the combinatorics of the collections of cells, like the combinatorial interpretation of the Krull dimension and the rook polynomial of a polyomino. The rook polynomial is a polynomial whose $k$-th coefficient is the number of ways of placing $k$ non-attacking rooks in a polyomino. The maximum number of non-attacking rooks is called the rook number of $\mathcal{P}$. The connection of the Castelnuovo-Mumford regularity of $K[\mathcal{P}]$ with the rook number is shown in [14] for the class of $L$-convex polyominoes. In general, the study of the $h$-polynomial in terms of the rook polynomial is a very interesting issue and this research has produced a relevant number of articles ([7, 11, 22, 24, 29, 31]). The rook polynomial also plays a crucial role in the study of the Gorensteiness of $K[\mathcal{P}]$ when it is a domain. A well-known result of Stanley ([35]) allows us to characterize the Gorenstein coordinate rings which are domain, checking if the $h$-polynomial is palindromic, which means counting the number of suitable configurations of non-attacking rooks. It is hard to characterize non-prime polyominoes whose coordinate ring is Gorenstein and such a class has never appeared so far. Some partial results are discussed in our paper (look at Corollaries 2.5, 3.9 and Theorem 4.18) and we point out that an approach using the initial ideal is not useful (Remark 3.10). Finally, for other very interesting combinatorial problems on polyomino ideals, we invite the reader to look at [1, 8, 10, 13, 15, 23, 32].
The article is structured as follows. Section 1 is devoted to provide some basic definitions and properties of collections of cells and the associated coordinate ring and to recall some useful known results of Commutative Algebra. In Section 2 we prove that the $h$-polynomial of a simple collection of cells is equal to the product of the $h$-polynomial of each connected component (Theorem 2.2), exploiting the monomial order in [27] and [5, Theorem 3.3]. As a direct consequence, we obtain in Corollary 2.5 that the $h$-polynomial of a simple thin collection of cells coincides with the rook-polynomial. In Section 3 we introduce a new class of non-prime collections of cells, named zig-zag collections. We show that their coordinate rings are Cohen-Macaulay (Theorem 3.6), we compute the Hilbert-Poincaré series in terms of the switching rook polynomial (Corollary 3.8) and we eventually discuss the Gorenstein property (Corollary  3.9). The results achieved in these sections are used in Section 4 to prove our main result, which states that the coordinate ring of a non-prime closed path is Cohen-Macaulay (Theorem 4.13). In the end, Theorem 4.18 summarizes the results about the primality and the Cohen-Macaulayness of this kind of polyominoes and provides a complete characterization of the Gorenstein property in terms of the length of the maximal blocks. We conclude the paper with Section 5, giving some open questions for future works.
1. Collections of cells and known results on graded $K$-algebras
This section is devoted to introduce several preliminaries. In particular, we give the definitions and the notations related to collections of cells and the related binomial ideals (see [28]) and we mention several well-known results in commutative algebra. We firstly point out that, in the whole paper, for each positive integer $n$ we use the standard notation $[n]$ to denote the set $\{1,\ldots,n\}$.
1.1. Collections of cells.
Let $(i,j),(k,l)\in{\mathbb{Z}}^{2}$. We say that $(i,j)\leq(k,l)$ if $i\leq k$ and $j\leq l$. Consider $a=(i,j)$ and $b=(k,l)$ in ${\mathbb{Z}}^{2}$ with $a\leq b$. The set $[a,b]=\{(m,n)\in{\mathbb{Z}}^{2}:i\leq m\leq k,\ j\leq n\leq l\}$ is called an interval of ${\mathbb{Z}}^{2}$. If $i<k$ and $j<l$, then we say that $[a,b]$ is a proper interval. In this case, we call $a$ and $b$ the diagonal corners of $[a,b]$, and $c=(i,l)$ and $d=(k,j)$ the anti-diagonal corners of $[a,b]$. If $j=l$ (or $i=k$), then $a$ and $b$ are in horizontal (or vertical) position. A proper interval $C=[a,b]$ with $b=a+(1,1)$ is called a cell of ${\mathbb{Z}}^{2}$; moreover, the elements $a$, $b$, $c$ and $d$ are called respectively the lower left, upper right, upper left and lower right corners of $C$. The set of vertices of $C$ is $V(C)=\{a,b,c,d\}$ and the set of edges of $C$ is $E(C)=\{\{a,c\},\{c,b\},\{b,d\},\{a,d\}\}$. Let $\mathcal{S}$ be a non-empty collection of cells in $\mathbb{Z}^{2}$. Then $V(\mathcal{S})=\bigcup_{C\in\mathcal{S}}V(C)$ and $E(\mathcal{S})=\bigcup_{C\in\mathcal{S}}E(C)$, while the rank of $\mathcal{S}$ is the number of cells that belong to $\mathcal{S}$. If $C$ and $D$ are two distinct cells of $\mathcal{S}$, then a walk from $C$ to $D$ in $\mathcal{S}$ is a sequence $\mathcal{C}:C=C_{1},\dots,C_{m}=D$ of cells of ${\mathbb{Z}}^{2}$ such that $C_{i}\cap C_{i+1}$ is an edge of $C_{i}$ and $C_{i+1}$ for $i=1,\dots,m-1$. Moreover, if $C_{i}\neq C_{j}$ for all $i\neq j$, then $\mathcal{C}$ is called a path from $C$ to $D$.
We say that $C$ and $D$ are connected in $\mathcal{S}$ if there exists a path of cells in $\mathcal{S}$ from $C$ to $D$. A polyomino $\mathcal{P}$ is a non-empty, finite collection of cells in $\mathbb{Z}^{2}$ where any two cells of $\mathcal{P}$ are connected in $\mathcal{P}$ (see Figure 1 (a)) Let $\mathcal{P}$ be a non-empty, finite collection of cells in $\mathbb{Z}^{2}$, $\mathcal{P}$ is called weakly connected if for any two cells $C$ and $D$ in $\mathcal{P}$ there exists a sequence of cells $\mathcal{C}:C=C_{1},\dots,C_{m}=D$ of $\mathcal{P}$ such that $V(C_{i})\cap V(C_{i+1})\neq\emptyset$ for all $i=1,\dots,m-1$. A subset of cells $\mathcal{P}^{\prime}$ of $\mathcal{P}$ is called a connected component of $\mathcal{P}$ if $\mathcal{P}^{\prime}$ is a polyomino and it is maximal with respect to the set inclusion, that is if $A\in\mathcal{P}\setminus\mathcal{P}^{\prime}$ then $\mathcal{P}^{\prime}\cup\{A\}$ is not a polyomino. For instance, see Figure 1 (b). Observe trivially that every polyomino is a weakly connected collection of cells.
(a) A polyomino.
(b) A collection of cells.
Figure 1. The collection of cells in the right has three connected components.
A sub-polyomino of $\mathcal{P}$ is a polyomino whose cells belong to $\mathcal{P}$. A weakly connected collection of cells is said to be thin if it does not contain the square tetromino (that is, the square obtained as a union of four distinct cells). We say that $\mathcal{P}$ is simple if for any two cells $C$ and $D$ not in $\mathcal{P}$ there exists a path of cells not in $\mathcal{P}$ from $C$ to $D$. A finite collection of cells $\mathcal{H}$ not in $\mathcal{P}$ is an hole of $\mathcal{P}$ if any two cells of $\mathcal{H}$ are connected in $\mathcal{H}$ and $\mathcal{H}$ is maximal with respect to set inclusion. For example, the collections of cells in Figure 1 are not simple. Each hole of $\mathcal{P}$ is a simple polyomino, and $\mathcal{P}$ is simple if and only if it has no hole. An interval $[a,b]$ with $a=(i,j)$, $b=(k,j)$ and $i<k$ is called a horizontal edge interval of $\mathcal{P}$ if the sets $\{(\ell,j),(\ell+1,j)\}$ are edges of cells of $\mathcal{P}$ for all $\ell=i,\dots,k-1$. In addition, if $\{(i-1,j),(i,j)\}$ and $\{(k,j),(k+1,j)\}$ do not belong to $E(\mathcal{P})$, then $[a,b]$ is called a maximal horizontal edge interval of $\mathcal{P}$. We define similarly a vertical edge interval and a maximal vertical edge interval.
Following [25], we call a zig-zag walk of $\mathcal{P}$ a sequence $\mathcal{W}:I_{1},\dots,I_{\ell}$ of distinct inner intervals of $\mathcal{P}$ where, for all $i=1,\dots,\ell$, the interval $I_{i}$ has either diagonal corners $v_{i}$, $z_{i}$ and anti-diagonal corners $u_{i}$, $v_{i+1}$ or anti-diagonal corners $v_{i}$, $z_{i}$ and diagonal corners $u_{i}$, $v_{i+1}$, such that:
  1. (1)
    $I_{1}\cap I_{\ell}=\{v_{1}=v_{\ell+1}\}$ and $I_{i}\cap I_{i+1}=\{v_{i+1}\}$, for all $i=1,\dots,\ell-1$;
  2. (2)
    $v_{i}$ and $v_{i+1}$ are on the same edge interval of $\mathcal{P}$, for all $i=1,\dots,\ell$;
  3. (3)
    for all $i,j\in\{1,\dots,\ell\}$ with $i\neq j$, there exists no inner interval $J$ of $\mathcal{P}$ such that $z_{i}$, $z_{j}$ belong to $J$.
Figure 2. An example of a zig-zag walk of $\mathcal{P}$.
Consider two cells $A$ and $B$ of $\mathbb{Z}^{2}$ with $a=(i,j)$ and $b=(k,l)$ as the lower left corners of $A$ and $B$ with $a\leq b$. A cell interval $[A,B]$ is the set of the cells of $\mathbb{Z}^{2}$ with lower left corner $(r,s)$ such that $i\leqslant r\leqslant k$ and $j\leqslant s\leqslant l$. If $(i,j)$ and $(k,l)$ are in horizontal (or vertical) position, we say that the cells $A$ and $B$ are in horizontal (or vertical) position.
Let $\mathcal{P}$ be a polyomino. Consider two cells $A$ and $B$ of $\mathcal{P}$ in a vertical or horizontal position. The cell interval $[A,B]$, containing $n>1$ cells, is called a block of $\mathcal{P}$ of rank n if all cells of $[A,B]$ belong to $\mathcal{P}$. The cells $A$ and $B$ are called extremal cells of $[A,B]$. We set also $[A,B[=[A,B]\setminus\{B\}$ and $]A,B[=[A,B]\setminus\{A,B\}$. Moreover, a block $\mathcal{B}$ of $\mathcal{P}$ is maximal if there does not exist any block of $\mathcal{P}$ which contains properly $\mathcal{B}$. It is clear that an interval of ${\mathbb{Z}}^{2}$ identifies a cell interval of ${\mathbb{Z}}^{2}$ and vice versa, hence we can associate to an interval $I$ of ${\mathbb{Z}}^{2}$ the corresponding cell interval denoted by $\mathcal{P}_{I}$. A proper interval $[a,b]$ is called an inner interval of $\mathcal{P}$ if all cells of $\mathcal{P}_{[a,b]}$ belong to $\mathcal{P}$. We denote by $\mathcal{I}(\mathcal{P})$ the set of all inner intervals of $\mathcal{P}$.
1.2. Switching rook polynomial.
Let $\mathcal{P}$ be a collection of cells. Two rooks in $\mathcal{P}$ are in non-attacking position if they do not belong to the same vertical or horizontal cell interval contained in $\mathcal{P}$. A $k$-rook configuration in $\mathcal{P}$ is a configuration of $k$ rooks which are arranged in $\mathcal{P}$ in non-attacking positions. Figure 3 shows a 6-rook configuration.
Figure 3. An example of a $6$-rook configuration in $\mathcal{P}$.
The rook number $r(\mathcal{P})$ is the maximum number of rooks that can be arranged in $\mathcal{P}$ in non-attacking positions. For all $k\in\{0,\ldots,r(\mathcal{P})\}$, we denote by $\mathcal{R}_{k}$ the set of all $k$-rook configurations in $\mathcal{P}$, setting conventionally $R_{0}=\{\emptyset\}$. We also set $r_{k}=|\mathcal{R}_{k}|$ for all $k\in\{0,\dots,r(\mathcal{P})\}$. The rook-polynomial of $\mathcal{P}$ is the polynomial in $\mathbb{Z}[t]$ defined as $r_{\mathcal{P}}(t)=\sum_{k=0}^{r(\mathcal{P})}r_{k}t^{k}$. Moreover, the set $\mathcal{R}=\mathcal{R}_{0}\cup\mathcal{R}_{1}\dots\cup\mathcal{R}_{r(\mathcal{P})}$ is a simplicial complex, called rook complex. We call a rectangle of $\mathcal{P}$ the set of all cells belonging to an inner interval of $\mathcal{P}$ and we say that two cells $A,B$, with left lower corners $(i,j),(h,k)$ respectively, are in diagonal position if $(i+1,j+1)\leq(h,k)$, while they are in anti-diagonal) if $(i+1,j+1)$ and $(h,k)$ are not comparable with respect to $\prec$. Two non-attacking rooks in $\mathcal{P}$ are switching rooks if they are placed in diagonal or anti-diagonal cells of a rectangle of $\mathcal{P}$. In such a case we say that the rooks are in a diagonal or anti-diagonal position, respectively. Fix $k\in\{0,\dots,r_{\mathcal{P}}\}$. Let $F\in\mathcal{R}_{k}$ and $R_{1}$ and $R_{2}$ be two switching rooks in $F$ in diagonal (resp. anti-diagonal) position. Let $R_{1}^{\prime}$ and $R_{2}^{\prime}$ be the rooks in anti-diagonal (resp. diagonal) cells in $R$. Then the set $\left(F\backslash\{R_{1},R_{2}\}\right)\cup\{R_{1}^{\prime},R_{2}^{\prime}\}$ belongs to $\mathcal{R}_{k}$. The operation of replacing $R_{1}$ and $R_{2}$ by $R_{1}^{\prime}$ and $R_{2}^{\prime}$ is called switch of $R_{1}$ and $R_{2}$. This induce the following equivalence relation $\sim$ on $\mathcal{R}_{k}$: let $F_{1},F_{2}\in\mathcal{R}_{k}$, then $F_{1}\sim F_{2}$ if $F_{2}$ can be obtained from $F_{1}$ after some switches. We define the quotient set ~$\tilde{\mathcal{R}}_{k}=\mathcal{R}_{k}/\sim$. We set $\tilde{r}_{k}=|\tilde{\mathcal{R}}_{k}|$ for all $k\in[r(\mathcal{P})]$. The switching rook-polynomial of $\mathcal{P}$ is the polynomial in $\mathbb{Z}[t]$ defined as $\tilde{r}_{\mathcal{P}}(t)=\sum_{k=0}^{r(\mathcal{P})}\tilde{r}_{k}t^{k}$. Observe that if $\mathcal{P}$ is a thin polyomino then $\tilde{r}_{\mathcal{P}}(t)=r_{\mathcal{P}}(t).$ Moreover, let $\mathcal{P}$ be a collection of cells. Consider an element $\mathcal{C}$ of $\tilde{R}_{k}$ and we assume that all rooks in $\mathcal{C}$ are in diagonal position. We say that $\mathcal{C}$ is a canonical configuration in $\mathcal{P}$ of $k$ rooks, so $\tilde{r}_{k}$ is the number of canonical positions in $\mathcal{P}$ of $k$ rooks.
Lemma 1.1.
Let $\mathcal{P}$ be a collection of cells consisting of $n$ connected components $\mathcal{P}_{1},\dots,\mathcal{P}_{n}$. Denote by $\tilde{r}_{\mathcal{P}_{i}}(t)$ the switching rook polynomial of $\mathcal{P}_{i}$. Then $\prod_{i=1}^{n}\tilde{r}_{\mathcal{P}_{i}}(t)$ is the switching rook polynomial of $\mathcal{P}$.
Proof.
It is enough to prove the claim when $\mathcal{P}$ is made up of two connected components $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Indeed, if the latter is proved, then $\tilde{r}_{1}(t)\tilde{r}_{2}(t)$ is the switching rook polynomial of $\mathcal{P}_{1}\cup\mathcal{P}_{2}$. Now, we consider $(\mathcal{P}_{1}\cup\mathcal{P}_{2})\cup\mathcal{P}_{3}$, so we have that $(\tilde{r}_{1}(t)\tilde{r}_{2}(t))\tilde{r}_{3}(t)$ is the switching rook polynomial of $(\mathcal{P}_{1}\cup\mathcal{P}_{2})\cup\mathcal{P}_{3}$. We can continue these arguments until $\mathcal{P}_{n}$, getting that $\prod_{i=1}^{n}\tilde{r}_{i}(t)$ is the switching rook polynomial of $\mathcal{P}$.
Assume that $\mathcal{P}$ consists of two connected components $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Denote by $\tilde{r}(t)$ the switching rook polynomial of $\mathcal{P}$ and we set
$\tilde{r}_{1}(t)=\sum_{i=0}^{r(\mathcal{P}_{1})}r_{i}^{(1)}t^{i}\quad\text{and}\quad\tilde{r}_{2}(t)=\sum_{j=0}^{r(\mathcal{P}_{2})}r_{j}^{(2)}t^{j}.$
Hence
$\tilde{r}_{1}(t)\tilde{r}_{2}(t)=\Bigg{(}\sum_{i=0}^{r(\mathcal{P}_{1})}r_{i}^{(1)}t^{j}\Bigg{)}\Bigg{(}\sum_{j=0}^{r(\mathcal{P}_{2})}r_{j}^{(2)}t^{j}\Bigg{)}=\sum_{k=0}^{r(\mathcal{P}_{1})+r(\mathcal{P}_{2})}c_{k}t^{k}.$
where $c_{k}=r_{0}^{(1)}r_{k}^{(2)}+r_{1}^{(1)}r_{k-1}^{(2)}+r_{2}^{(1)}r_{k-2}^{(2)}+\dots+r_{k}^{(1)}r_{0}^{(2)}$. From the structure of $\mathcal{P}$, observe that two rooks placed in different connected components are surely in non-attacking position. So, the number of the canonical configurations of $k$-rooks in $\mathcal{P}$ is obtained considering the sum of the products of the number of the canonical configurations of $k_{1}$-rooks in $\mathcal{P}_{1}$ and that one of the canonical configurations of $k_{2}$-rooks in $\mathcal{P}_{2}$, for all non-negative integers $k_{1},k_{2}$ such that $k_{1}+k_{2}=k$. Hence, it follows that $c_{k}$ represents the number of the canonical configurations of $k$-rooks in $\mathcal{P}$, which is the $k$-th coefficient of $\tilde{r}(t)$. Therefore $\tilde{r}_{1}(t)\tilde{r}_{2}(t)$ is the switching rook polynomial of $\mathcal{P}$. ∎
1.3. The coordinate ring attached to a collection of cells
Following [28], we define a binomial ideal attached to a collection of cells. Let $\mathcal{P}$ be a collection of cells. Set $S_{\mathcal{P}}=K[x_{v}|v\in V(\mathcal{P})]$, where $K$ is a field. If $[a,b]$ is an inner interval of $\mathcal{P}$, with $a$,$b$ and $c$,$d$ respectively diagonal and anti-diagonal corners, then the binomial $x_{a}x_{b}-x_{c}x_{d}$ is called an inner 2-minor of $\mathcal{P}$. We defined the binomial ideal $I_{\mathcal{P}}$ in $S_{\mathcal{P}}$ as the ideal generated by all the inner 2-minors of $\mathcal{P}$ and we call it the inner 2-minors ideal of $\mathcal{P}$. If $\mathcal{P}$ is a polyomino, then $I_{\mathcal{P}}$ is simply said the polyomino ideal of $\mathcal{P}$. We set also $K[\mathcal{P}]=S_{\mathcal{P}}/I_{\mathcal{P}}$, which is the coordinate ring of $\mathcal{P}$. We denote by $G(\mathcal{P})$ the set of the generators of $I_{\mathcal{P}}$. For a generalization of collections of cells and the related algebras look at [8].
Remark 1.2.
Let $\mathcal{P}$ be a collection of cells and assume that it is written as a disjoint union of the collection of cells $\mathcal{P}_{1},\dots,\mathcal{P}_{r}$ such that $V(\mathcal{P})$ is the disjoint union of vertex sets $V(\mathcal{P}_{j})$, with $j\in[r]$. Then $K[\mathcal{P}]\cong\bigotimes_{i=1}^{r}K[\mathcal{P}_{j}]$. It follows that $K[\mathcal{P}]$ is a normal Cohen–Macaulay domain if and only if $K[\mathcal{P}_{j}]$ is a normal Cohen–Macaulay domain for all $j\in[r]$. Hence it is not restrictive to consider just weakly connected collections of cells. For this reason, whenever the term “collection of cells” is introduced in the paper, we mean a weakly connected collection of cells.
For utility, we state the following result about the Cohen-Macaulay property of the simple collection of cells, that in literature we found stated in this form only in the case $\mathcal{P}$ is polyomino (see [18, Theorem 2.1] and [19, Corollary 3.3]). For the sake of completeness, we also provide its proof.
Proposition 1.3.
Let $\mathcal{P}$ be a simple collection of cells. Then $K[\mathcal{P}]$ is a normal Cohen-Macaulay domain of dimension $|V(\mathcal{P})|-|\mathcal{P}|$.
Proof.
By [5, Theorem 3.3] we know that $I_{\mathcal{P}}=J_{\mathcal{P}}$ where $J_{\mathcal{P}}$ is the toric ring of a bipartite weakly chordal graph. In particular $I_{\mathcal{P}}$ is prime and, by [27], $I_{\mathcal{P}}$ has a squarefree quadratic Gröbner basis. A toric ring whose toric ideal admits a squarefree initial ideal is normal, see [16, Corollary 4.26] and by a theorem of Hochster (see [2, Theorem 6.3.5]), a normal toric ring is Cohen–Macaulay. Finally, by [28, Theorem 3.5] and [28, Corollary 3.6] we know that $I_{\mathcal{P}}=I_{\Lambda}$ for a saturated lattice $\Lambda$ such that $\operatorname{rank}_{\mathbb{Z}}(\Lambda)=|\mathcal{P}|$. Hence $\operatorname{height}(I_{\mathcal{P}})=|\mathcal{P}|$ (see [16, Problem 3.12]) and so $\dim K[\mathcal{P}]=|V(\mathcal{P})|-|\mathcal{P}|$ (see also [37, Corollary 3.1.7]). ∎
In accordance to [3], we recall the definition of a closed path polyomino, and the configuration of cells characterizing its primality.
Definition 1.4.
We say that a polyomino $\mathcal{P}$ is a closed path if it is a sequence of cells $A_{1},\dots,A_{n},A_{n+1}$, $n>5$, such that:
  1. (1)
    $A_{1}=A_{n+1}$;
  2. (2)
    $A_{i}\cap A_{i+1}$ is a common edge, for all $i=1,\dots,n$;
  3. (3)
    $A_{i}\neq A_{j}$, for all $i\neq j$ and $i,j\in\{1,\dots,n\}$;
  4. (4)
    For all $i\in\{1,\dots,n\}$ and for all $j\notin\{i-2,i-1,i,i+1,i+2\}$ then $V(A_{i})\cap V(A_{j})=\emptyset$, where $A_{-1}=A_{n-1}$, $A_{0}=A_{n}$, $A_{n+1}=A_{1}$ and $A_{n+2}=A_{2}$.
Figure 4. A closed path with two $L$-configurations and a ladder of four steps.
A path of five cells $C_{1},C_{2},C_{3},C_{4},C_{5}$ of $\mathcal{P}$ is called an L-configuration if the two sequences $C_{1},C_{2},C_{3}$ and $C_{3},C_{4},C_{5}$ go in two orthogonal directions. A set =$\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1,\dots,n}$ of maximal horizontal (or vertical) blocks of rank at least two, with $V(\mathcal{B}_{i})\cap V(\mathcal{B}_{i+1})=\{a_{i},b_{i}\}$ and $a_{i}\neq b_{i}$ for all $i=1,\dots,n-1$, is called a ladder of $n$ steps if $[a_{i},b_{i}]$ is not on the same edge interval of $[a_{i+1},b_{i+1}]$ for all $i=1,\dots,n-2$. For instance, in Figure 4 there is a closed path having two $L$-configurations and a ladder of four steps. It is proved in [3, Section 6] that a closed path has no zig-zag walks if and only if it contains an $L$-configuration or a ladder of at least three steps. Furthermore, in [3, Theorem 6.2] the authors characterize the primality of a closed path $\mathcal{P}$ proving that $I_{\mathcal{P}}$ is prime if and only if $\mathcal{P}$ does not contain any zig-zag walk.
1.4. Some results in Commutative Algebra.
In this subsection we recall some well-known results in Commutative Algebra that we use extensively throughout the paper. We recall, for completeness, that $\operatorname{depth}(M)\leq\dim(M)$ holds for every $R$-module $M$ over a Noetherian ring $R$, and $M$ is Cohen-Macaulay when equality holds. About some notations we use throughout the paper, if $S$ is a polynomial ring, $\prec$ is a monomial order on it and $f\in S$ is a polynomial, we denote by $\operatorname{in}_{\prec}(f)$, or $\operatorname{in}(f)$ if the order is clear in the context, the initial monomial of $f$ with respect to $\prec$. As usual, if $S=K[x_{1},\ldots,x_{n}]$ and $m$ is a monomial, then $\operatorname{supp}(m)=\{x_{i}\in S\mid x_{i}\text{ divides }m\}$.
We start by providing some basics on the Hilbert function and the Hilbert-Poincaré series of a graded $K$-algebra $R/I$. Let $R$ be a graded $K$-algebra and $I$ be an homogeneous ideal of $R$. Then $R/I$ has a natural structure of graded $K$-algebra as $\bigoplus_{k\in\mathbb{N}}(R/I)_{k}$. The numerical function $\mathrm{H}_{R/I}:\mathbb{N}\to\mathbb{N}$ with $\mathrm{H}_{R/I}(k)=\dim_{K}(R/I)_{k}$ is called the Hilbert function of $R/I$. The formal power series $\mathrm{HP}_{R/I}(t)=\sum_{k\in\mathbb{N}}\mathrm{H}_{R/I}(k)t^{k}$ is called the Hilbert-Poincaré series of $R/I$. It is known by Hilbert-Serre theorem that there exists a polynomial $h(t)\in\mathbb{Z}[t]$ with $h(1)\neq 0$ such that $\mathrm{HP}_{R/I}(t)=\frac{h(t)}{(1-t)^{d}}$, where $d$ is the Krull dimension of $R/I$. The polynomial $h(t)$ is called h-polynomial of $R/I$. Moreover, if $R/I$ is Cohen-Macaulay then $\mathrm{reg}(R/I)=\deg h(t)$. Recall also that if $S=K[x_{1},\ldots,x_{n}]$ then $\mathrm{HP}_{S}(t)=\frac{1}{(1-t)^{n}}$. In some parts of the paper, if $K[\mathcal{P}]$ is the coordinate ring of a collection of cells, we use the notation $h_{K[\mathcal{P}]}$ for the $h$-polyonimial of the Hilbert-Poincaré series of $K[\mathcal{P}]$.
Proposition 1.5.
If $I$ is a homogeneous ideal of a graded ring $R$ then:
  1. (1)
    $R/I$ and $R/\mathrm{in}(I)$ have the same Hilbert function for all monomial orders (see [37, Corollary 3.3.15]);
  2. (2)
    If $R/\mathrm{in}(I)$ is Cohen–Macaulay (resp. Gorenstein), then $R/I$ is Cohen–Macaulay (resp. Gorenstein) (see [37, Proposition 9.6.17]).
From [9] we deduce that the converse for the Cohen-Macaulayness in (2) of Proposition 1.5 holds for a homogeneous ideal whose initial ideal is squarefree.
Proposition 1.6 ([9]).
Let $I$ be a homogeneous ideal in a polynomial ring $S$ and suppose $\mathrm{in}(I)$ is radical with respect to some term order. Then $\mathrm{reg}(S/I)=\mathrm{reg}(S/\mathrm{in}(I))$ and $\mathrm{depth}(S/I)=\mathrm{depth}(S/\mathrm{in}(I))$. Furthermore, $S/I$ is Cohen-Macaulay if and only if $S/\mathrm{in}(I)$ is Cohen-Macaulay.
For the following result, see Lemma 5.1.2 and Lemma 5.1.3 of [37].
Proposition 1.7.
Let $R$ be a graded $K$-algebra and $I$ be a graded ideal of $R$. Let $q$ be an homogeneous element of $R$ of degree $m$ and let
$0\longrightarrow R/(I:q)\longrightarrow R/I\longrightarrow R/(I,q)\longrightarrow 0$
be a short exact sequence. Then $\mathrm{HP}_{R/I}(t)=\mathrm{HP}_{R/(I,q)}(t)+t^{m}\mathrm{HP}_{R/(I:q)}(t)$.
For the following result, see [37, Lemma 5.1.11].
Proposition 1.8.
Let $A$ and $B$ be standard graded $K$-algebras over a field $K$. Then $\mathrm{HP}_{A\otimes_{K}B}(t)=\mathrm{HP}_{A}(t)\cdot\mathrm{HP}_{B}(t)$. In particular $\mathrm{dim}(A\otimes_{K}B)=\mathrm{dim}(A)+\mathrm{dim}(B)$.
For the following result, see Lemma 3.1.34, Lemma 3.1.35 and Proposition 11.5.8 of [37].
Proposition 1.9.
Let $A$ and $B$ be standard graded $K$-algebras over a field $K$. Then $\mathrm{depth}(A\otimes_{K}B)=\mathrm{depth}(A)+\mathrm{depth}(B)$. In particular $A\otimes_{K}B$ is Cohen-Macaulay (Gorenstein) if and only if $A$ and $B$ are Cohen-Macaulay (Gorenstein).
For the following result, see [2, Proposition 1.2.9].
Proposition 1.10.
Let $0\longrightarrow N_{1}\longrightarrow N_{2}\longrightarrow N_{3}\longrightarrow 0$ be a short exact sequence of finitely generated $R$-module over a Noetherian ring $R$. Then:
  • $\mathrm{depth}(N_{2})\geq\min\{\mathrm{depth}(N_{3}),\mathrm{depth}(N_{1})\}$.
  • $\mathrm{depth}(N_{1})\geq\min\{\mathrm{depth}(N_{2}),\mathrm{depth}(N_{3})+1\}$.
  • $\mathrm{depth}(N_{3})\geq\min\{\mathrm{depth}(N_{1})-1,\mathrm{depth}(N_{2})\}$.
2. Hilbert-Poincaré series of simple collections of cells
In this section we present some results on simple collections of cells. From [6, Theorem 3.3] and [27] we know that, if $\mathcal{P}$ is a simple collection of cells, the set of generators of $I_{\mathcal{P}}$ forms the reduced Gröbner basis of $I_{\mathcal{P}}$ with respect to a suitable monomial order. The monomial order provided in [27] is fundamental for our purposes because it allows us to write $S_{\mathcal{P}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}})$ as $K$-tensor product of suitable $K$-algebras coming from the connected components. We start proving the following Lemma.
Lemma 2.1.
Let $\mathcal{P}$ be a collection of cells and $[a,b]$ and $[\alpha,\beta]$ be two inner intervals with $\alpha=b$ (see Figure 5a). Let $\prec$ be a monomial order on $S_{\mathcal{P}}$ and $f_{a,b}$ and $f_{\alpha,\beta}$ be the generators of $I_{\mathcal{P}}$ attached to $[a,b]$ and $[\alpha,\beta]$ respectively. Assume that $S(f_{a,b},f_{\alpha,\beta})$ reduces to 0 modulo the set of generators of $I_{\mathcal{P}}$ with respect to $\prec$. Suppose $[c,\gamma]$ and $[d,\delta]$ are not inner intervals of $\mathcal{P}$, then $\gcd(\mathop{\rm in}\nolimits(f_{a,b}),\mathop{\rm in}\nolimits(f_{\alpha,\beta}))=1$.
Figure 5. Arrangements of intervals.
The same holds considering Figure 5b, that is, if $S(f_{\gamma,\delta},f_{c,d}))$ reduces to 00 and $[r,b]$ and $[b,h]$ are not inner intervals, then $\gcd(\mathop{\rm in}\nolimits(f_{c,d}),\mathop{\rm in}\nolimits(f_{\gamma,\delta}))=1$.
Proof.
From Figure 5a it is clear that $f_{a,b}=x_{a}x_{b}-x_{c}x_{d}$ and $f_{\alpha,\beta}=x_{\alpha}x_{\beta}-x_{\gamma}x_{\delta}$. Suppose that $\gcd(\mathop{\rm in}\nolimits(f_{a,b}),\mathop{\rm in}\nolimits(f_{\alpha,\beta}))\neq 1$. Then $\mathop{\rm in}\nolimits(f_{a,b})=x_{a}x_{b}$ and $\mathop{\rm in}\nolimits(f_{\alpha,\beta})=x_{\alpha}x_{\beta}=x_{b}x_{\beta}$. This leads that $S(f_{a,b},f_{\alpha,\beta})=x_{a}x_{\gamma}x_{\delta}-x_{\beta}x_{c}x_{d}$. Assume that $\mathop{\rm in}\nolimits(S(f_{a,b},f_{\alpha,\beta}))=x_{a}x_{\gamma}x_{\delta}$. Since $\mathop{\rm in}\nolimits(f_{\alpha,\beta})=x_{b}x_{\beta}$, if $S(f_{a,b},f_{\alpha,\beta}))$ reduces to 00 then $x_{a}x_{\gamma}-x_{r}x_{d}$ is an inner 2-minor having initial monomial $x_{a}x_{\gamma}$, or $x_{a}x_{\delta}-x_{c}x_{h}$ is an inner 2-minor with initial monomial $x_{a}x_{\delta}$. Hence, we obtain $[c,\gamma]$ or $[d,\delta]$ is an inner interval, a contradiction. Assume that $\mathop{\rm in}\nolimits(S(f_{a,b},f_{\alpha,\beta}))=x_{\beta}x_{c}x_{d}$. Since $\mathop{\rm in}\nolimits(f_{a,b})=x_{a}x_{b}$, if $S(f_{a,b},f_{\alpha,\beta}))$ reduces to 00 then $x_{c}x_{\beta}-x_{r}x_{\delta}$ is an inner 2-minor having initial monomial $x_{c}x_{\beta}$, or $x_{d}x_{\beta}-x_{\gamma}x_{h}$ is an inner 2-minor with initial monomial $x_{d}x_{\beta}$. Hence, we obtain $[c,\gamma]$ or $[d,\delta]$ as an inner interval, a contradiction. The last claim, for Figure 5b, holds using the same arguments. ∎
Theorem 2.2.
Let $\mathcal{P}$ be a simple collection of cells consisting of $n$ connected components $\mathcal{P}_{1},\ldots,\mathcal{P}_{n}$. Let $h_{K[\mathcal{P}_{i}]}(t)$ be the $h$-polynomial of $K[\mathcal{P}_{i}]$ for each $i=1,\ldots,n$, then $\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}(t)$ is the $h$-polynomial of $K[\mathcal{P}]$.
Proof.
Let $\mathcal{P}$ be a simple collection of cells. According to the references [5, Theorem 3.3] and [27], $I_{\mathcal{P}}$ possesses a quadratic Gröbner basis with respect to a certain monomial order. We fix such a monomial order, then for any pair of inner 2-minors $f$ and $g$, their $S$-polynomial $S(f,g)$ reduces to 0. Suppose that $f$ and $g$ are inner 2-minors of $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ respectively, where $i\neq j$. Now we have two cases: when $f$ and $g$ are binomials associated with inner intervals that are disjoint or they share a vertex. For the first, we have $\gcd(\mathop{\rm in}\nolimits(f),\mathop{\rm in}\nolimits(g)=1$ trivially. For the second case, since the inner 2-minors $f$ and $g$ are of the form as shown in Figure 5 and $S(f,g)$ reduces to 00, so by Lemma 2.1 we obtain $\gcd(\mathop{\rm in}\nolimits(f),\mathop{\rm in}\nolimits(g))=1$ as well. So, for all $i,j\in[n]$ with $i\neq j$ and for all $f,g$ inner 2-minors of $I_{\mathcal{P}}$ such that $f$ is related to an inner interval of $\mathcal{P}_{i}$ and $g$ is related to an inner interval of $\mathcal{P}_{j}$, we have $\gcd(\mathop{\rm in}\nolimits(f),\mathop{\rm in}\nolimits(g))=1$. As a consequence, if $v\in V(\mathcal{P}_{i})\cap V(\mathcal{P}_{j})$ for some $i,j\in[n]$ with $i\neq j$, and $x_{v}\in\mathrm{supp}(\mathop{\rm in}\nolimits(f))$, for some inner 2-minor $f$ related to an inner interval of $\mathcal{P}_{i}$, then $x_{v}\notin\mathrm{supp}(\mathop{\rm in}\nolimits(g))$, for all inner 2-minors $g$ related to the inner intervals of $\mathcal{P}_{j}$. Let $A_{\mathcal{P}}=\{v\in V(\mathcal{P})\mid v\in V(\mathcal{P}_{i})\cap V(\mathcal{P}_{j}),\text{ for some }i,j\in[n],i\neq j\}$. Let $C_{\mathcal{P}}$ be the set of vertices $v\in A_{\mathcal{P}}$ such that $x_{v}\notin\mathrm{supp}(\mathop{\rm in}\nolimits(f))$, for all inner 2-minor $f$ related to an inner interval of $\mathcal{P}$. For each $i\in[n]$, define $D_{\mathcal{P}_{i}}$ the set of vertices $v\in V(\mathcal{P}_{i})\cap A_{\mathcal{P}}$ such that $x_{v}\notin\mathrm{supp}(\mathop{\rm in}\nolimits(f)),\text{ for all inner 2-minor }f\text{ related to an inner interval of }\mathcal{P}_{i}$. Denote $d_{i}=|D_{\mathcal{P}_{i}}|$. Set, for all $i\in[n]$, $S^{\prime}_{\mathcal{P}_{i}}=K[x_{a}:a\in V(\mathcal{P}_{i})\setminus D_{\mathcal{P}_{i}}]$. Then we have:
$S_{\mathcal{P}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}})=K[x_{v}\mid v\in C_{\mathcal{P}}]\otimes_{K}\left(\bigotimes_{i=1}^{n}S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})\right).$
Observe that for all $i\in[n]$ we have $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})=S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})\otimes_{K}K[x_{v}\mid v\in D_{\mathcal{P}_{i}}]$. In particular, by Proposition 1.8, $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=(1-t)^{d_{i}}\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)$, and since $\mathcal{P}_{i}$ is a simple polyomino, by Proposition 1.3 we have $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=\frac{h_{\mathcal{P}_{i}}(t)}{(1-t)^{|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-d_{i}}}$. Note that $\sum_{i=1}^{n}|V(\mathcal{P}_{i})|=|V(\mathcal{P})|+|A_{\mathcal{P}}|$ and $\sum_{i=1}^{n}|\mathcal{P}_{i}|=|\mathcal{P}|$. Therefore, by Proposition 1.8, we have:
$\mathrm{HP}_{K[\mathcal{P}]}(t)=\frac{\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|+|A_{\mathcal{P}}|+|C_{\mathcal{P}}|-\sum_{i=1}^{n}d_{i}}}$
Observe that $|A_{\mathcal{P}}|=\sum_{i=1}^{n}d_{i}-|C_{\mathcal{P}}|$. In fact, let $v\in A_{\mathcal{P}}$. Then there exist $j_{1},j_{2}\in[n]$ such that $v\in V(\mathcal{P}_{j_{1}})\cap V(\mathcal{P}_{j_{2}})$. Furthermore, if $v\in C_{\mathcal{P}}$, then $v\in D_{\mathcal{P}_{j_{1}}}\cap D_{\mathcal{P}_{j_{2}}}$, and if $v\notin C_{\mathcal{P}}$, then either $v\in D_{\mathcal{P}_{j_{1}}}$ or $v\in D_{\mathcal{P}_{j_{2}}}$. By this observation, we easily argue the previous equality. So, $\mathrm{HP}_{K[\mathcal{P}]}(t)=\frac{\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|}}$. By Proposition 1.3 we have $\dim K[\mathcal{P}]=|V(\mathcal{P})|-|\mathcal{P}|$, hence $\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}(t)$ is the $h$-polynomial of $K[\mathcal{P}]$. ∎
By the previous result, we argue that, if $\mathcal{P}$ is a simple collection of cells, the study of the $h$-polynomial of $K[\mathcal{P}]$ in terms of the switching rook polynomial of $\mathcal{P}$ can be reduced to investigate the Hilbert-Poincaré series of its connected components. In particular, if for all simple polyominoes [29, Conjecture 3.2] is true, then the same holds for all simple collections of cells.
Corollary 2.3.
Let $\mathcal{P}$ be a simple collection of cells consisting of $n$ connected components $\mathcal{P}_{1},\dots,\mathcal{P}_{n}$. If $h_{K[\mathcal{P}_{i}]}(t)$ is equal to the switching rook polynomial of $\mathcal{P}_{i}$ for all $i\in[n]$, then $h_{K[\mathcal{P}]}(t)$ is equal to the switching rook polynomial of $\mathcal{P}$.
Proof.
It is a consequence of Lemma 1.1 and Theorem 2.2. ∎
Example 2.4.
Let $\mathcal{P}$ be a simple collection of cells where every connected component $\mathcal{P}_{i}$ is a parallelogram polyomino. Then, by [29, Theorem 3.5], $h_{K[\mathcal{P}_{i}]}(t)$ is the switching rook polynomial of $\mathcal{P}_{i}$ for all $i\in[n]$, so $h_{K[\mathcal{P}]}$ is the switching rook polynomial of $\mathcal{P}$ by Corollary 2.3.
As a consequence, we generalize [31, Theorem 3.13 and Theorem 4.2] to simple thin collections of cells (see Figure 6 for an example). This is useful in the next sections.
We recall the definition of S-property given in [31]. Let $\mathcal{P}$ be a thin collection of cells. A cell $C$ is called single if there exists a unique maximal interval of $\mathcal{P}$ containing $C$. $\mathcal{P}$ has the S-property if every maximal interval of $\mathcal{P}$ has only one single cell.
Figure 6. A simple thin polyomino and a simple thin collection of cells.
Corollary 2.5.
Let $\mathcal{P}$ be a simple thin collection of cells. Then $h_{K[\mathcal{P}]}$ is the rook polynomial of $\mathcal{P}$, in particular $\operatorname{reg}(K[\mathcal{P}])=r(\mathcal{P})$. Moreover, $K[\mathcal{P}]$ is Gorenstein if and only if $\mathcal{P}$ satisfies the $S$-property.
Proof.
As a consequence of Corollary 2.3 and [31, Theorem 3.13], we get that $h_{K[\mathcal{P}]}$ is the rook polynomial of $\mathcal{P}$. Moreover, $\operatorname{reg}(K[\mathcal{P}])=r(\mathcal{P})$ since $K[\mathcal{P}]$ is a Cohen-Macaulay ring. Now we prove the result about the Gorensteiness. Let $h_{K[\mathcal{P}]}(t)=r_{\mathcal{P}}(t)=\sum_{k=0}^{s}r_{k}t^{k}$, where $s=r(\mathcal{P})$. Since $K[\mathcal{P}]$ is a Cohen-Macaulay domain, it is known from [35] that $K[\mathcal{P}]$ is Gorenstein if and only if $r_{i}=r_{s-i}$ for all $i=0,\dots,s$, that is $h_{K[\mathcal{P}]}(t)$ is palindromic. Assume that $\mathcal{P}$ consists of $n$ connected components and denote by $h_{K[\mathcal{P}_{i}]}(t)$ the $h$-polynomial of $K[\mathcal{P}_{i}]$ for all $i\in[n]$. Suppose that $\mathcal{P}$ has the $S$-property and we prove that $K[\mathcal{P}]$ is Gorenstein. Since $\mathcal{P}$ has the $S$-property, then each connected component $\mathcal{P}_{i}$ has the $S$-property as well. By [31, Theorem 4.2], $h_{K[\mathcal{P}_{i}]}(t)$ is palindromic for all $i\in[n]$, so $h_{K[\mathcal{P}]}(t)$ is palindromic as well, since $h_{K[\mathcal{P}]}(t)=\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}(t)$ and the product of palindromic polynomials is palindromic. Hence it follows by [35] that $K[\mathcal{P}]$ is Gorenstein. Now, suppose that $K[\mathcal{P}]$ is Gorenstein and we show that $\mathcal{P}$ has the $S$-property. In particular the rook polynomial of $\mathcal{P}$, that is $h_{K[\mathcal{P}]}$, is palindromic and observe also that $r_{0}=1$ and $r_{1}=|\mathcal{P}|$. We assume by contradiction that $\mathcal{P}$ does not have the $S$-property. For all $i\in[n]$, let $h_{K[\mathcal{P}_{i}]}(t)=\sum_{j=0}^{s_{i}}r_{j}^{(i)}t^{j}$ be the $h$-polynomial of $K[\mathcal{P}_{i}]$, where $s_{i}=r(\mathcal{P}_{i})$ and observe that $r_{1}^{(i)}=|\mathcal{P}_{i}|$. Looking at the proof of case (b)$\Rightarrow$(c) of [31, Theorem 4.2], for all $k\in[n]$ such that $\mathcal{P}_{k}$ has not the $S$-property we have $r_{s_{k}}>1$ or $r_{s_{k}-1}>|\mathcal{P}_{k}|$. Therefore, it follows easily that $r_{s}>1$ or $r_{s-1}>\sum_{i=1}^{n}|\mathcal{P}_{i}|=|\mathcal{P}|$, that is a contradiction, since $h_{K[\mathcal{P}]}(t)$ is palindromic. ∎
3. Zig-zag collections
In this section we discuss the algebraic properties of the coordinate ring of a new family of collections of cells called zig-zag collections. Roughly speaking, a zig-zag collection is a collection of cells with just one hole that comes from the intervals related to a zig zag-walk. We start by pointing out that definition.
Definition 3.1.
Let $I_{1},\dots,I_{\ell}$ be a sequence of distinct intervals in $\mathbb{Z}^{2}$. For all $i\in\{1,\ldots,\ell\}$ denote by $\mathcal{P}_{i}$ the collection of cells related to $I_{i}$ and $\mathcal{P}=\bigcup_{i=1}^{\ell}\mathcal{P}_{i}$. $\mathcal{P}$ is a zig-zag collection if satisfies:
  1. (1)
    $I_{1}\cap I_{\ell}=\{v_{1}=v_{\ell+1}\}$ and $I_{i}\cap I_{i+1}=\{v_{i+1}\}$, for all $i=1,\dots,\ell-1$;
  2. (2)
    $v_{i}$ and $v_{i+1}$ are on the same edge interval of $\mathcal{P}$, for all $i=1,\dots,\ell$.
  3. (3)
    $I_{i}\cap I_{j}=\emptyset$ for all $\{i,j\}\subseteq[\ell]$ with $i<j$, such that $j\neq i+1$ and $(i,j)\neq(1,\ell)$.
In such a case, we say that $\mathcal{P}$ is supported by $I_{1},\dots,I_{\ell}$.
In Figure 7 (A) a zig-zag collection is shown. In Figure 7 (B) we figure out an example of a non zig-zag collection, in fact $|I_{1}\cap I_{4}|=4.$
(a) A zig-zag collection.
(b) A non zig-zag collection.
Figure 7. The collection of cells in the right is not a zig-zag collection but contains a zig-zag walk.
Remark 3.2.
Let $\mathcal{P}$ be a zig-zag collection. From Definition 3.1 it follows that $\mathcal{P}$ is a non-simple weakly connected collection of cells having just one hole and $\ell$ is an even number.
Moreover, $I_{\mathcal{P}}$ is not prime because $I_{1},\dots,I_{\ell}$ is a zig-zag walk of $\mathcal{P}$. For more details we refer to [25, Section 3], where the arguments given for polyominoes can be extended easily to collections of cells.
Our first step is to study the Gröbner basis of the inner 2-minors ideal of a zig-zag collection. For this aim, we introduce four total orders on $\mathbb{Z}^{2}$. Let $(i,j),(k,l)\in\mathbb{Z}^{2}$, then we say that:
  • $(k,l)<^{(1)}(i,j)$, if $l<j$, or $l=j$ and $i<k$;
  • $(k,l)<^{(2)}(i,j)$, if $l<j$, or $l=j$ and $k<i$;
  • $(k,l)<^{(3)}(i,j)$, if $l>j$, or $l=j$ and $k<i$;
  • $(k,l)<^{(4)}(i,j)$, if $l>j$, or $l=j$ and $i<k$.
Let $\mathcal{P}$ be a collection of cells. For $i\in\{1,2,3,4\}$, denote by $<_{\mathrm{lex}}^{(i)}$ the lexicographic order on $K[x_{v}\mid v\in V(\mathcal{P})]$, induced by the following total order on the variables: $x_{a}<_{\mathrm{lex}}^{(i)}x_{b}$ if and only if $a<^{(i)}b$ for $a,b\in V(\mathcal{P})$. If $f$ is an inner 2-minor in $I_{\mathcal{P}}$, observe that for $i\in\{1,3\}$ the initial monomial of $f$ with respect to $<_{\mathrm{lex}}^{(i)}$ is related to the anti-diagonal corners of the inner interval associated. While for $i\in\{2,4\}$ the initial monomial of $f$ is related to the diagonal corners. Moreover, [28, Theorem 4.1] and the arguments of its proof hold also considering the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{2,4\}$. The same can be considered for [28, Remark 4.2] with respect to the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{1,3\}$. Similar considerations can be made using the reverse lexicographic order, with reference to [26, Proposition 3.2]. For our aim, we state the result below, which can be proved by the same arguments of the mentioned results in [28] or also using the lemmas in [6, Section 3].
Proposition 3.3.
Let $\mathcal{P}$ be a collection of cells and $f,g\in I_{\mathcal{P}}$ inner 2-minors. Assume $f$ is related to the inner interval $[a,b]$, with anti-diagonal corners $c,d$ and $g$ is related to the inner interval $[\alpha,\beta]$, with anti-diagonal corners $\gamma,\delta$. Without loss of generality, assume the second coordinate of $a$ is smaller or equal to the second coordinate of $\alpha$, $c$ belongs to the same vertical edge interval of $a$ and $\gamma$ belongs to the same horizontal edge interval of $\alpha$. For $i\in\{1,\ldots,4\}$ consider $S^{(i)}(f,g)$ be the S-polynomial of $f,g$ with respect to $<^{(i)}_{\mathrm{lex}}$. Then the following holds:
  • For $i\in\{2,4\}$, $S^{(i)}(f,g)$ does not reduce to 0 modulo $G(\mathcal{P})$ with respect to $<^{(i)}_{\mathrm{lex}}$ if and only $b=\alpha$ and $[a,\gamma],[a,\delta]$ are not inner intervals of $\mathcal{P}$ (see Figure 8a).
  • For $i\in\{1,3\}$, $S^{(i)}(f,g)$ does not reduce to 0 modulo $G(\mathcal{P})$ with respect to $<^{(i)}_{\mathrm{lex}}$ if and only $c=\gamma$ and the sets $\{d,\alpha\},\{d,\beta\}$ are not the set of anti-diagonal corners of inner intervals of $\mathcal{P}$ (see Figure 8b).
Figure 8. Configurations related to conditions of Proposition 3.3.
Lemma 3.4.
Let $\mathcal{P}$ be a zig-zag collection supported by $I_{1},\dots,I_{\ell}$. Then there exists a monomial order $\prec$ on $S_{\mathcal{P}}$ such that:
  1. (1)
    the set of generators $G(\mathcal{P})$ of $I_{\mathcal{P}}$ forms the reduced Gröbner basis of $I_{\mathcal{P}}$ with respect $\prec$;
  2. (2)
    the order $\prec$ induces a monomial order $<_{i}$ on $S_{\mathcal{P}_{i}}$ for all $i\in[\ell]$, such that for all $k\in[\ell/2]$ we have:
    • $\mathrm{in}_{<_{2k-1}}(I_{\mathcal{P}_{2k-1}})=(\{x_{a}x_{b}\mid a,b\ \mbox{are anti-diagonal corners of an inner interval of}\ I_{2k-1}\})$;
    • $\mathrm{in}_{<_{2k}}(I_{\mathcal{P}_{2k}})=(\{x_{c}x_{d}\mid c,d\ \mbox{are diagonal corners of an inner interval of}\ I_{2k}\})$;
    • $\mathrm{in}_{\prec}(I_{\mathcal{P}})=\sum_{i=1}^{\ell}\mathrm{in}_{<_{i}}(I_{\mathcal{P}_{i}})$.
In particular, $I_{\mathcal{P}}$ is radical.
Proof.
We start by defining inductively a suitable total order $<_{V(\mathcal{P})}$ on $V(\mathcal{P})$. It is not restrictive to suppose that $I_{1}$, $I_{2}$ and $I_{\ell}$ are arranged as in Figure 9a, otherwise it is sufficient to do some suitable reflections of $\mathcal{P}$ or to relabel the intervals $\{I_{i}\}_{i\in[\ell]}$.
Figure 9. Arrangement of $I_{1}$, $I_{2}$ and $I_{\ell}$ and the related orders.
In the first step we give a total order on $(I_{1}\setminus\{v_{1}\})\cup I_{2}$. We use $<^{(3)}$ and $<^{(2)}$ respectively to order the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$; look at Figure 9b, where the arrows denote the sense of the orders of the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$ respectively. Moreover, we set that every vertex of $\mathcal{P}$ in $I_{1}\setminus\{v_{1}\}$ is smaller than each one in $I_{2}\setminus\{v_{2}\}$. Let $k\in[\ell/2]$ with $k\geq 2$ and consider the intervals $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$. Assume that a total order in $I_{2k-2}\setminus\{v_{2k-2}\}$ is already defined; we want to define a total order on $(I_{2k-1}\setminus\{v_{2k-1}\})\cup I_{2k}$. First of all, we introduce two total orders on $I_{2k-1}\setminus\{v_{2k-1}\}$ and on $I_{2k}\setminus\{v_{2k}\}$ in the following way.
  1. (1)
    If $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 10a and 10d or as in Figures 11a, 11b and 11d, then we use $<^{(1)}$ and $<^{(2)}$ to order the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ respectively.
  2. (2)
    In the case that $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are as in Figure 10b, then we define an order of the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ using $<^{(3)}$ and $<^{(2)}$ respectively.
  3. (3)
    When $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are in the position of Figure 10c, then use respectively $<^{(1)}$ and $<^{(4)}$ to make order in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$.
  4. (4)
    If $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are in the position of Figure 11c then apply respectively $<^{(3)}$ and $<^{(4)}$ to define an order in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$.
Figure 10. Horizontal arrangements of $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ and the related orders.
Figure 11. Vertical arrangements of $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ and the related orders.
Therefore, depending on the described situations, we have two total orders on $I_{2k-1}\setminus\{v_{2k-1}\}$ and on $I_{2k}\setminus\{v_{2k}\}$; moreover, putting that every vertex of $\mathcal{P}$ in $I_{2k-1}\setminus\{v_{2k-1}\}$ is smaller than everyone in $I_{2k}\setminus\{v_{2k}\}$, we get a total order on $(I_{2k-1}\setminus\{v_{2k-1}\})\cup I_{2k}$. Applying inductively the procedure until $k=\ell/2$, we can get a total order $<_{V(\mathcal{P})}$ for the set of the vertices of $\mathcal{P}$. Now, define the lexicographic order $<_{\mathrm{lex}}$ on $S_{\mathcal{P}}$ induced by the order on the variables: $x_{b}<_{\mathrm{lex}}x_{a}$, where $a,b\in V(\mathcal{P})$, if $b<_{V(\mathcal{P})}a$. We prove that $<_{\mathrm{lex}}$ provides the desired claim.
We start proving (1). Let $f=x_{a}x_{b}-x_{c}x_{d}$ and $g=x_{p}x_{q}-x_{r}x_{s}$ be two generators of $I_{\mathcal{P}}$ attached to the inner intervals $[a,b]$ and $[p,q]$ of $\mathcal{P}$, respectively. We want to show that the $S$-polynomial $S(f,g)$ of $f$ and $g$ reduces to 00 with respect to $G(\mathcal{P})$. Let $k\in[\ell/2]$. We may consider just the case when $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 10b (with the convention that $I_{0}=I_{\ell}$); indeed, all the other situations described in Figures 10 and 11 can be proved similarly. If $[a,b],[p,q]\subseteq I_{2k-1}$ or $[a,b],[p,q]\subseteq I_{2k}$ then $S(f,g)$ reduces to 00 with respect to $G(\mathcal{P})$ by Proposition 3.3. Assume that $[a,b]\subseteq I_{2k-1}$ and $[p,q]\subseteq I_{2k}$. In such a case, observe that $[a,b]\cap[p,q]=\emptyset$ or $[a,b]\cap[p,q]=\{b\}$ with $b=p$. Since in this case we have also $\operatorname{in}_{<}(f)=x_{c}x_{d}$ and $\operatorname{in}_{<}(g)=x_{p}x_{q}$, then in both cases we obtain $\gcd(\operatorname{in}_{<}(f),\operatorname{in}_{<}(g))=1$. Hence $S(f,g)$ reduces to 00 with respect to $G(\mathcal{P})$. It is trivial to see that $\gcd(\operatorname{in}_{<}(f),\operatorname{in}_{<}(g))=1$ occurs when $[a,b]\subseteq I_{i}$ and $[p,q]\subseteq I_{j}$ with $i,j\in[\ell]$ and $j>i+1.$ In conclusion, the claim (1) is completely proved.
Now, observe that $<_{\mathrm{lex}}$ induces in a natural way a monomial order $<_{i}$ on $S_{\mathcal{P}_{i}}$, for $i\in[\ell]$, which is the restriction of $<_{\mathrm{lex}}$ on $S_{\mathcal{P}_{i}}$. In particular, the claim (2) follows from the considerations done for claim (1).
Finally, it is known that if the initial ideal of a homogeneous ideal $I$ of $K[x_{1},\dots,x_{n}]$, with respect to a monomial order, is squarefree then $I$ is radical (see [17, Corollary 2.2]). Since $\mathop{\rm in}\nolimits_{<_{\mathrm{lex}}}(I_{\mathcal{P}})$ is squarefree, then $I_{\mathcal{P}}$ is radical. ∎
Remark 3.5.
Let $\mathcal{P}$ be a zig-zag collection and $\prec$ be the monomial order satisfying the conditions of Lemma 3.4. Denote $\mathcal{E}_{i}=I_{i}\setminus\{v_{i}\}$ for all $i\in[\ell]$. We highlight that, by the proof of Lemma 3.4, the order $\prec$ is the lexicographic order induced by a total order on $V(\mathcal{P})$ and it satisfies the following property: if $a\in\mathcal{E}_{i}$ and $b\in\mathcal{E}_{j}$ with $i<j$ then $x_{a}\prec x_{b}$. This fact is useful further.
For shortness we will write $\mathrm{in}(I_{\mathcal{P}})$ and $\mathrm{in}(I_{\mathcal{P}_{i}})$ for $\mathrm{in}_{\prec}(I_{\mathcal{P}})$ and $\mathrm{in}_{<_{i}}(I_{\mathcal{P}_{i}})$ respectively. Moreover, following the construction in Lemma 3.4, without loss of generality we can always assume that $v_{1}$ is an anti-diagonal corner in $I_{1}$. For all $i\in\{1\ldots,\ell\}$ denote by $S^{\prime}_{i}=K[x_{a}\mid a\in V(I_{i})\setminus\{v_{i}\}]$. Set also $\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)=\frac{h_{i}(t)}{(1-t)^{n_{i}}}$, in particular we know that $n_{i}=|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$.
Theorem 3.6.
Let $\mathcal{P}$ be a zig-zag collection supported by $I_{1},\dots,I_{\ell}$. Then:
  1. (1)
    $\mathrm{HP}_{K[\mathcal{P}]}(t)=(1-t)^{\ell}\prod_{i=1}^{\ell}\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)=\frac{h_{1}(t)\cdot h_{2}(t)\cdots h_{\ell}(t)}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|}},$
    where $h_{i}(t)$ is the $h$-polynomial of $K[\mathcal{P}_{i}].$
  2. (2)
    $K[\mathcal{P}]$ is Cohen-Macaulay with Krull dimension $|V(\mathcal{P})|-|\mathcal{P}|$.
Proof.
(1) By Lemma 3.4 we obtain that $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})=\otimes_{i=1}^{\ell}S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$. Observe that for all $i\in[\ell]$ then $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})=S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})\otimes_{K}K[x_{v_{i}}]$. So we argue that $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=(1-t)\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)$ and we obtain our claim on $\mathrm{HP}_{K[\mathcal{P}]}(t)$.
(2) By [28, Theorem 2.2] we know that $K[\mathcal{P}_{i}]$ is a Cohen-Macaulay domain of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$, so by Proposition 1.6 also $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$. Therefore, by Proposition 1.9 we obtain that $S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1$ and as consequence $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Cohen-Macaulay of dimension $\sum_{i=1}^{\ell}(|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1)=|V(\mathcal{P})|-|\mathcal{P}|$. By Proposition 1.6 the same property holds for $K[\mathcal{P}]$. ∎
Example 3.7.
We provide here an example of a non zig-zag collection of cells, whose coordinate ring is Cohen-Macaulay. Let $\mathcal{P}$ be the collection of cells in Figure 12.
Figure 12. A not zig-zag collection of cells
Denote by $<$ the lexicographic order on $S_{\mathcal{P}}$ induced by the following total order of the variables:
$x_{34}>x_{42}>x_{45}>x_{44}>x_{54}>x_{73}>x_{61}>x_{65}>x_{57}>x_{47}>x_{37}>x_{16}>x_{15}>x_{14}>x_{22}>x_{72}>$
$>x_{66}>x_{64}>x_{63}>x_{62}>x_{56}>x_{55}>x_{53}>x_{52}>x_{51}>x_{46}>x_{43}>x_{36}>x_{33}>x_{32}>x_{27}>$
$>x_{26}>x_{25}>x_{24}>x_{23}>x_{13}$
By some computations we get that the set of the generators of $I_{\mathcal{P}}$ forms the reduced Gröbner basis of $I_{\mathcal{P}}$ with respect to $<$ and the initial ideal of $I_{\mathcal{P}}$ is generated by the monomials:
$x_{42}x_{54},x_{45}x_{54},x_{34}x_{63},x_{44}x_{63},x_{54}x_{63},x_{73}x_{62},x_{65}x_{56},x_{42}x_{55},x_{34}x_{53},x_{42}x_{53},x_{45}x_{53},x_{44}x_{53},x_{61}x_{52},x_{57}x_{46},$
$x_{34}x_{43},x_{57}x_{36},x_{47}x_{36},x_{22}x_{33},x_{57}x_{26},x_{47}x_{26},x_{37}x_{26},x_{16}x_{25},x_{16}x_{24},x_{15}x_{24},x_{16}x_{23},x_{15}x_{23},x_{14}x_{23}.$
Hence $I_{\mathcal{P}}$ is radical but it is not prime since $[(2,2),(3,3)],[(3,3),(6,4)],[(6,2),(7,3)],[(5,1),(6,2)],$ $[(4,2),(5,5)],[(5,5),(6,6)],[(2,6),(5,7)],[(1,3),(2,6)]$ is a zig-zag walk of $\mathcal{P}$. Using Macaulay2 ([4, 12]) we check that $S_{\mathcal{P}}/\mathrm{in}_{<}(I_{\mathcal{P}})$ is a Cohen-Macaulay ring. Hence, from Proposition 1.6, we have that $K[\mathcal{P}]$ is Cohen-Macaulay.
We point out that we needed to check the Cohen-Macaulayness of $S_{\mathcal{P}}/\mathrm{in}_{<}(I_{\mathcal{P}})$ and not of $S_{\mathcal{P}}/I_{\mathcal{P}}$ using Macaulay2, because the time of computation to check the same property for $S_{\mathcal{P}}/I_{\mathcal{P}}$ is too long.
Corollary 3.8.
Let $\mathcal{P}$ be a zig-zag collection supported by $I_{1},\dots,I_{\ell}$. Then
$\mathrm{HP}_{K[\mathcal{P}]}(t)=\frac{h(t)}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|}},$
where $h(t)$ is the switching rook polynomial of $\mathcal{P}$. Moreover $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
Proof.
For $i\in[\ell]$ we denote by $h_{i}(t)$ the $h$-polynomial of $K[\mathcal{P}_{i}]$, so from [29, Theorem 3.5] we have that $h_{i}(t)$ is the switching rook-polynomial of $\mathcal{P}_{i}$ for all $i\in[\ell]$. Recall that Theorem 3.6 states in particular that $h(t)=\prod_{i=1}^{\ell}h_{i}(t)$. From Lemma 1.1 we have that $\prod_{i=1}^{\ell}h_{i}(t)$ is the switching rook polynomial of $\cup_{i=1}^{\ell}\mathcal{P}_{i}$, that is $h(t)$ is the switching rook polynomial of $\mathcal{P}$. Moreover, since $K[\mathcal{P}]$ is Cohen-Macaulay from Theorem 3.6, then it follows that $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$. ∎
Moreover, if $I$ is an interval of ${\mathbb{Z}}^{2}$ and $\mathcal{P}_{I}$ is the attached polyomino then we say that $\mathcal{P}_{I}$ is a square if $I=[a,a+n(1,1)]$ for $a\in{\mathbb{Z}}^{2}$ and for $n>0.$ If $n=1$, then $I$ is a cell of ${\mathbb{Z}}^{2}$.
Corollary 3.9.
Let $\mathcal{P}$ be a zig-zag collection supported by $I_{1},\dots,I_{\ell}$.
  1. (1)
    If $K[\mathcal{P}]$ is Gorenstein, then $\mathcal{P}_{i}$ is a square for all $i\in[\ell]$.
  2. (2)
    If $\mathcal{P}_{i}$ is a cell for all $i\in[\ell]$, then $K[\mathcal{P}]$ is Gorenstein.
Proof.
(1) Let $h(t)=\sum_{k=1}^{s}h_{k}t^{k}$ be the $h$-polynomial of $K[\mathcal{P}]$. From Corollary 3.8 we have that $h(t)$ is the switching rook polynomial of $\mathcal{P}$ and $s=r(\mathcal{P})$. Since $K[\mathcal{P}]$ is Gorenstein, it follows from [37, Corollary 5.3.10] that $h_{k}=h_{s-k}$ for all $k\in[s]$. In particular, we have that $h_{s}=1$. Suppose by contradiction that there exists $h\in[\ell]$ such that $\mathcal{P}_{h}$ is not a square. Up to reflections, rotations or translations of $\mathcal{P}$, we may assume that $\mathcal{P}_{h}=\mathcal{P}_{[(1,1),(m.n)]}$, where $m>n$. Observe that $r(\mathcal{P})=n-1$. We define a canonical configuration $\mathcal{T}_{1}$ of $n-1$ rooks in $\mathcal{P}_{h}$, placing a rook in the cell of $\mathcal{P}_{h}$ having lower left corner $(j,j)$ for all $j\in[n-1]$. Moreover, since $m>n$, we can define another canonical configuration $\mathcal{T}_{2}$ of $n-1$ rooks in $\mathcal{P}_{h}$, placing a rook in the cell of $\mathcal{P}_{h}$ having lower left corner $(j,j)$ for all $j\in[n-2]$ and a rook in that one with $(n,n-1)$ as lower left corner. Denote by $\mathcal{C}_{i}$ a canonical configuration of $r(\mathcal{P}_{i})$ rooks in $\mathcal{P}_{i}$, for all $i\in[s]\setminus\{h\}$, and set $\mathcal{C}=\cup_{i\in[s]\setminus\{h\}}\mathcal{C}_{i}$. It is easy to see that $\mathcal{C}\cup\mathcal{T}_{1}$ and $\mathcal{C}\cup\mathcal{T}_{2}$ are two canonical configuration of $r(\mathcal{P})$ rooks in $\mathcal{P}$, so $h_{s}>1$. This is a contradiction because $h_{s}=1$. In conclusion $\mathcal{P}_{i}$ is a square for all $i\in[\ell]$.
(2) Observe that $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ can be viewed as an edge ring of a graph $G$ on the vertex set $V(\mathcal{P})$, where $\{a,b\}\in E(G)$ if either $a,b$ are the anti-diagonal corners of an inner interval of $I_{2k}$ or they are the diagonal corners of an inner interval of $I_{2k-1}$, for $k\in[\ell/2]$. If $\mathcal{P}_{i}$ is a cell for all $i\in[\ell]$, $G$ is a disjoint union of edges, so $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Gorenstein by [20, Corollary 9.3.3]. Therefore, it follows that $K[\mathcal{P}]$ is Gorenstein by (2) of Proposition 1.5. ∎
Remark 3.10.
If there exists $j\in[\ell]$ such that $\mathcal{P}_{j}$ is a square different from a cell, then $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ cannot be Gorenstein because $G$ is a chordal graph but it is not a disjoint union of edges. Anyway, we tested different examples with Macaulay2 that shows $K[\mathcal{P}]$ is still Gorenstein. So, we ask if the converse of (1) is true in general.
4. Cohen-Macaulay property and Hilbert series for closed path polyominoes having zig-zag walks
From now onward, we assume that $\mathcal{P}$ is a non-prime closed path polyomino. By [3, Theorem 6.2], $\mathcal{P}$ contains a zig-zag walk. See Figure 13 for an example of a closed path with zig-zag walks.
Figure 13. Closed path with zig-zag walks.
{tikzpicture}
{tikzpicture}
Figure 14. Possible “changes of direction” in a closed path having a zig-zag walk.
Discussion 4.1.
If $\mathcal{P}$ is a closed path polyomino having zig-zag walks, by [3, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration or ladders with at least three steps. So, any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 14a or 14b (up to reflections or rotations). Furthermore, by [6] we know that there exists a monomial order such that the Gröbner basis is quadratic and squarefree. The monomial order that provides this property can be built following [6, Algorithm 4.1]. Following this construction, we obtain that for all inner intervals containing the vertex $m$ as in Figure 14a, (up to reflection and rotation), the initial term of the related inner 2-minors contain the variable $x_{m}$ in its support. From now onward we denote by $M_{\mathcal{P}}$ the set of all vertices $m$ in $\mathcal{P}$ as in Figure 14a, up to rotations and reflections, and by $C_{m}$ the cell labeled with $X$ with reference to the same figure.
Set-up 4.2.
Let $\mathcal{P}$ be a closed path polyomino. With reference to Discussion 4.1, assume $M_{\mathcal{P}}=\{m_{1},\ldots,m_{r}\}$ for some $r\in\mathbb{N}$. Then we define for all $i\in[r]$:
  1. (1)
    $\mathcal{Q}_{0}=\mathcal{P}$ and $\mathcal{Q}_{i}=\mathcal{P}\setminus\{C_{m_{1}},\ldots,C_{m_{i}}\}$;
  2. (2)
    $\mathcal{Q}^{\prime}_{i}=\mathcal{Q}_{i-1}\setminus\{C\in\mathcal{P}\mid C\ \mbox{belongs to a maximal cell interval containing}\ C_{m_{i}}\}.$
Remark 4.3.
Let $\mathcal{P}$ be a closed path with a zig-zag walk and $\prec$ be the monomial order on $S_{\mathcal{P}}$ defined by [6, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$. For all $i\in[r]$, with reference to Set-up 4.2, since $V(\mathcal{Q}_{i})\subset V(\mathcal{P})$ and $V(\mathcal{Q}_{i}^{\prime})\subset V(\mathcal{P})$, the monomial order $\prec$ can be considered also in $S_{\mathcal{Q}_{i}}$ and $S_{\mathcal{Q}^{\prime}_{i}}$. In particular, $\prec$ trivially preserves the same initial terms in the ideals $I_{\mathcal{P}}$, $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}^{\prime}_{i}}$. Furthermore, by [7, Proposition 4.1], $\prec$ provides a quadratic squarefree Gröbner basis also for $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}_{i}^{\prime}}$. So, with abuse of notation, we refer to $\mathrm{in}(I_{\mathcal{Q}_{i}})$, $\mathrm{in}(I_{\mathcal{Q}_{i-1}})$ and $\mathrm{in}(I_{\mathcal{Q}^{\prime}_{i}})$ considering the monomial order $\prec$.
Example 4.4.
In Figure 15 we show an example of a closed path $\mathcal{P}$ and the related collection of cells $\mathcal{Q}_{r}$ and $\mathcal{Q}_{2}^{\prime}$, with reference to Set-up 4.2.
(a) $\mathcal{P}$
(b) $\mathcal{Q}_{r}$
(c) $\mathcal{Q}_{2}^{\prime}$
Figure 15. An example of a polyomino $\mathcal{P}$ and the related $\mathcal{Q}_{r}$, $\mathcal{Q}_{4}$ and $\mathcal{Q}_{2}^{\prime}$.
Lemma 4.5.
Let $\mathcal{P}$ be closed path polyomino. In the framework of Set-up 4.2 we have:
$\mathrm{HP}_{K[\mathcal{P}]}(t)=\mathrm{HP}_{K[\mathcal{Q}_{r}]}(t)+\sum_{i=1}^{r}\frac{t}{1-t}\mathrm{HP}_{K[\mathcal{Q}^{\prime}_{i}]}(t)=\mathrm{HP}_{K[\mathcal{Q}_{r}]}(t)+\frac{t}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|}}\left(\sum_{i=1}^{r}r_{\mathcal{Q}^{\prime}_{i}}(t)\right)$
Moreover $\mathrm{depth}(K[\mathcal{P}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{r}]),|V(\mathcal{P})|-|\mathcal{P}|\}$.
Proof.
Let $\prec$ be the monomial order in $S_{\mathcal{P}}$ defined by [6, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$. For all $i\in[r]$, we focus on the initial ideals of $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}_{i}^{\prime}}$ as explained in Remark 4.3. In particular, let $i\in[r]$ and consider the following exact sequence:
$0\longrightarrow\frac{S_{\mathcal{Q}_{i-1}}}{(\mathrm{in}(I_{\mathcal{Q}_{i-1}}):x_{m_{i}})}\longrightarrow\frac{S_{\mathcal{Q}_{i-1}}}{\mathrm{in}(I_{\mathcal{Q}_{i-1}})}\longrightarrow\frac{S_{\mathcal{Q}_{i-1}}}{(\mathrm{in}(I_{\mathcal{Q}_{i-1}}),x_{m_{i}})}\longrightarrow 0$
Arguing as in the proof of [7, Proposition 4.3], we can prove that:
  • $\frac{S_{\mathcal{Q}_{i-1}}}{(\mathrm{in}(I_{\mathcal{Q}_{i-1}}),x_{m_{i}})}\cong\frac{S_{\mathcal{Q}_{i}}}{\mathrm{in}(I_{\mathcal{Q}_{i}})}$
  • $\frac{S_{\mathcal{Q}_{i-1}}}{(\mathrm{in}(I_{\mathcal{Q}_{i-1}}):x_{m_{i}})}\cong\frac{S_{\mathcal{Q}^{\prime}_{i}}}{\mathrm{in}(I_{\mathcal{Q}^{\prime}_{i}})}\otimes K[x_{m_{i}}]$.
So by Propositions 1.5 and 1.7 it follows that for all $i\in\{1\,\ldots,r\}$ it is verified:
$\mathrm{HP}_{K[\mathcal{Q}_{i-1}]}(t)=\mathrm{HP}_{K[\mathcal{Q}_{i}]}(t)+\frac{t}{1-t}\mathrm{HP}_{K[\mathcal{Q}^{\prime}_{i}]}(t)$
and also, by Propositions 1.6, 1.9 and 1.10, we obtain $\mathrm{depth}(K[\mathcal{Q}_{i-1}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{i}]),\mathrm{depth}(K[\mathcal{Q}_{i}^{\prime}])+1\}$. In particular, we obtain the first equality of $\mathrm{HP}_{K[\mathcal{P}]}(t)$.
Moreover, in the framework of Set-up 4.2, we have $|V(\mathcal{Q}_{i-1})|-|\mathcal{Q}_{i-1}|=|V(\mathcal{Q}_{i})|-|\mathcal{Q}_{i}|$ and $|V(\mathcal{Q}_{i-1})|-|\mathcal{Q}_{i-1}|=|V(\mathcal{Q}^{\prime}_{i})|-|\mathcal{Q}^{\prime}_{i}|+1$ for all $i\in\{1,\ldots,r\}$. So $|V(\mathcal{P})|-|\mathcal{P}|=|V(\mathcal{Q}^{\prime}_{i})|-|\mathcal{Q}^{\prime}_{i}|+1$ for all $i\in\{1,\ldots,r\}$. Therefore, considering that $\mathcal{Q}^{\prime}_{i}$ is a simple collection of cells, then $K[\mathcal{Q}^{\prime}_{i}]$ is a normal Cohen-Macaulay domain of dimension $|V(\mathcal{Q}^{\prime}_{i})|-|\mathcal{Q}^{\prime}_{i}|$ for each $i\in[r]$ (by Proposition 1.3), in particular $\operatorname{depth}(K[\mathcal{Q}_{i}^{\prime}])=|V(\mathcal{Q}^{\prime}_{i})|-|\mathcal{Q}^{\prime}_{i}|$. Hence, $\mathrm{depth}(K[\mathcal{Q}_{i-1}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{i}]),|V(\mathcal{P})|-|\mathcal{P}|\}$, obtaining the bound on $\operatorname{depth}(K[\mathcal{P}])$. Finally, the second equality in the expression of $\mathrm{HP}_{K[\mathcal{P}](t)}$, follows by Corollary 2.5. ∎
Corollary 4.6.
Let $\mathcal{P}$ be a non-prime closed path polyomino not containing the configuration in Figure 14b. Then $K[\mathcal{P}]$ is Cohen-Macaulay of dimension $|V(\mathcal{P})|-|\mathcal{P}|$.
Proof.
If $\mathcal{P}$ is a non-prime closed path polyomino not containing the configuration in Figure 14b then $\mathcal{Q}_{r}$ is a zig-zag collection and by Theorem 3.6 we have that $K[\mathcal{Q}_{r}]$ is Cohen-Macaulay of dimension $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{P})|-|\mathcal{P}|$. Hence $\mathrm{depth}(K[\mathcal{P}])\geq|V(\mathcal{P})|-|\mathcal{P}|$ but we also know that $\mathrm{dim}(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$ (see [10, Theorem 3.4]), so $K[\mathcal{P}]$ is Cohen-Macaulay of dimension $|V(\mathcal{P})|-|\mathcal{P}|$. ∎
Observe that the collection of cells $\mathcal{Q}_{r}$ in Lemma 4.5 does not contain any configuration as in Figure 14a. Now we focus on such a kind of configuration. Having in mind Corollary 4.6, we have to consider the case $\mathcal{Q}_{r}$ contains configurations as in Figure 14b.
Figure 16. Configurations examined in Discussion 4.7 and Set-up 4.8
Discussion 4.7.
Let $\mathcal{Q}_{r}$ be a collection of cells obtained as in Set-up 4.2. Recall that $\mathcal{Q}_{r}$ is obtained from a closed path $\mathcal{P}$ having zig-zag walks, by removing all cells $X$ in each sub-configurations of $\mathcal{P}$ as in Figure 14a (up to reflections and rotations). Observe that, since $\mathcal{P}$ has not $L$-configurations and ladders of at least three steps, then $\mathcal{Q}_{r}$ can be built as a union of opportune reflections and rotations of the configurations in Figure 16. That is, we can express $\mathcal{Q}_{r}=\bigcup_{k=1}^{q}\mathcal{C}_{k}$, where $\mathcal{C}_{k}$ is a configuration as in Figure 16a or Figure 16b or Figure 16c (up to rotations and reflections), for all $k\in\{1,\ldots,q\}$, for some positive integer $q$. In particular, with reference to the highlighted points in the figures, we build $\mathcal{Q}_{r}$ by gluing those configurations with this procedure: setting the first configuration $\mathcal{C}_{k}$, the second one, say $\mathcal{C}_{l}$ with $l=k+1$, is glued considering an opportune reflection or rotation of it, in such a way that the vertices $c_{k},d_{k}$ and $e_{k}$ in orange of the first configuration are overlapped, respectively, to the vertices $c_{l-1},d_{l-1}$ and $e_{l-1}$ in blue of the second configuration. See Figure 17 for an example of this construction.
Figure 17. Example of gluing of the configurations in Figure 16a and Figure 16b, where the overlapped part is in green.
Set-up 4.8.
Let $\mathcal{Q}_{r}$ be a collection of cells obtained as in Set-up 4.2. As explained in Discussion 4.7, set $\mathcal{Q}_{r}=\bigcup_{k=1}^{q}\mathcal{C}_{k}$ for some positive integer $q$, where $\mathcal{C}_{k}$ is a configuration as in Figure 16a or Figure 16b or Figure 16c (up to rotations and reflections), for all $k\in[q]$. Without loss of generality, we can assume that $\mathcal{C}_{1}$ is a configuration as in Figure 16a or Figure 16b and that has the same orientation provided in those figures. For all $k\in[q]$, denote by $E_{k},B_{i,k},A_{i,k},D_{k}$ and by $c_{k-1},d_{k-1},e_{k-1},c_{k},d_{k},e_{k},a_{i,k}^{(j_{i})},b_{i,k}^{(j_{i})}$, for $j_{i}\in\{1,2\}$, respectively the cells and the vertices in $\mathcal{C}_{k}$ as in Figure 16. Moreover, we set $n_{k}$ as the index such that $a_{n_{k},k}^{(1)}$ and $a_{n_{k},k}^{(2)}$ belong to the same edge interval of $c_{k}$. By the structure of $\mathcal{Q}_{r}$ we have $c_{q}=c_{0}$, $d_{q}=d_{0}$ and $e_{q}=e_{0}$, where $e_{0},d_{0},c_{0}$ are the vertices of $\mathcal{C}_{1}$ in blue.
[…]
For all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$ define:
  1. (1)
    $\mathcal{R}_{i,k}=\mathcal{Q}_{r}\setminus(\{B_{j,h}\mid h<k,j\in[n_{h}]\}\cup\{B_{j,k}\mid j\leq i\})$.
  2. (2)
    $\mathcal{R}^{\prime}_{i,k}=\mathcal{R}_{i,k}\setminus\{C\in\mathcal{Q}_{r}\mid C\ \mbox{belongs to a maximal cell interval containing}\ B_{i,k}\}.$
In what follows, we often refer to the framework explained in Set-up 4.8, and we made extensive use of the notations introduced there, also with the help of figures.
Example 4.9.
We provide here an example of the framework described in in Set-up 4.8, with the help of a figure. In Figure 18 we show an example of a collection of cells of kind $\mathcal{Q}_{r}$ and some of the other related collection of cells. In this case $\mathcal{Q}_{r}=\bigcup_{k=1}^{q}\mathcal{C}_{k}$ with $q=8$ and $J_{\mathcal{Q}_{r}}=\{1,3,7\}$. Observe that the last collection of cells that we build from the described procedure is $\mathcal{R}_{2,7}$, and it is a zig-zag collection.
(a) $\mathcal{Q}_{r}$
(b) $\mathcal{R}_{2,7}$
(c) $\mathcal{R}_{1,3}^{\prime}$
Figure 18. An example of a collection of cells $\mathcal{Q}_{r}$ and the related collections $\mathcal{R}_{2,7}$ and $\mathcal{R}_{1,3}^{\prime}$.
With reference to Set-up 4.8, for the following lemma, observe that if $\prec$ is a monomial order on $S_{\mathcal{Q}_{r}}$ then it can be viewed also as a monomial order in $S_{\mathcal{R}_{i,k}}$ and $S_{\mathcal{R}_{i,k}^{\prime}}$ for all $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$.
Lemma 4.10.
Let $\mathcal{Q}_{r}$ be a collection of cells obtained as in Set-up 4.2. Using notations of Set-up 4.8, then there exists a term order $\prec$ on $S_{\mathcal{Q}_{r}}$ satisfying the following:
  1. (1)
    The ideals $I_{\mathcal{Q}_{r}},I_{\mathcal{R}_{i,k}}$ and $I_{\mathcal{R}_{i,k}^{\prime}}$, for all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$, have a squarefree quadratic Gröbner basis with respect to $\prec$.
  2. (2)
    For all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$, if $x_{b_{i,k}^{(2)}}\in\operatorname{supp}(x_{u}x_{v})$ for some inner 2-minor $f=\pm(x_{u}x_{v}-x_{w}x_{z})\in I_{\mathcal{Q}_{r}}$, then $\operatorname{in}(f)=x_{u}x_{v}$.
Proof.
Denote $s=\max J_{\mathcal{Q}_{r}}$. Observe that $\mathcal{R}_{n_{s},s}$ is a zig-zag collection, so we denote by $I_{1},\ldots,I_{p}$ the set of maximal intervals of such a zig-zag collection, for some integer $p$. We can assume, without loss of generality, that $I_{1}$ is related to the cell interval $[E_{1},B_{1,1}]\setminus\{B_{1,1}\}$, with reference to the notations explained in Set-up 4.8 and to Figures 16a and 16b, assuming also that $I_{1}$ have the same orientation of these figures.
By Lemma 3.4, we know that there exists a term order $\prec_{\mathcal{R}_{n_{s},s}}$ for which $I_{\mathcal{R}_{n_{s},s}}$ has a quadratic Gröbner basis and satisfies:
  • (C1)
    If $i\in[p]$ is odd, the initial monomials of all inner 2-minors related to $I_{i}$ are the monomials related to anti-diagonal corners.
  • (C2)
    If $i\in[p]$ is even, the initial monomials of all inner 2-minors related to $I_{i}$ are the monomials related to diagonal corners.
Now, we focus in $\mathcal{Q}_{r}$ on the subconfiguration $\mathcal{C}_{k}$ for $k\in[q]$, with reference to Set-up 4.8. Recall that each configuration $\mathcal{C}_{k}$ is obtained up to reflection and rotations of the configurations in Figure 16. Let $\mathcal{C}_{k}^{\prime}=\mathcal{C}_{k}\setminus\{B_{j,k}\mid j\in\{1,\ldots,n_{k}\}\}$. Therefore, $\mathcal{R}_{n_{s},s}=\bigcup_{k=1}^{q}\mathcal{C}_{k}^{\prime}$ and the intervals $I_{j}$ are, up to reflections or rotations, of kind $[E_{k},B_{i,k}[$ or $[A_{i,k},B_{i+1,k}[$ or $[A_{n_{k},k},D_{k}]$, as in Figure 16a. For all $k\in[q]$, considering the order $\prec_{R_{n_{s},s}}$, in the configuration $\mathcal{C}_{k}^{\prime}$ the following holds:
  • (R1)
    Assume $k\in J_{\mathcal{R}_{n_{s},s}}$. For all $i\in[n_{k}]$, if $x_{b_{i,k}^{(1)}}\in\operatorname{supp}(x_{u}x_{v})$ with $f=\pm(x_{u}x_{v}-x_{w}x_{z})\in I_{\mathcal{R}_{n_{s},s}}$ then $\operatorname{in}(f)=x_{u}x_{v}$.
  • (R2)
    Consider the vertex $d_{k-1}$ with the convention $d_{0}=d_{q}$. If $x_{d_{k-1}}\in\operatorname{supp}(x_{u}x_{v})$ with $f=\pm(x_{u}x_{v}-x_{w}x_{z})\in I_{\mathcal{R}_{n_{s},s}}$ such that $u,v,w,z\in V([E_{k},B_{1,k}])$ (or $u,v,w,z\in V([E_{k},D_{k}])$ if $k\notin J_{\mathcal{Q}_{r}}$), then $\operatorname{in}(f)=x_{u}x_{v}$.
  • (R3)
    Consider the vertex $d_{k}$. Then $x_{d_{k}}\notin\operatorname{supp}(\operatorname{in}(f))$ for every inner interval $f=x_{u}x_{v}-x_{w}x_{z}\in I_{\mathcal{R}_{n_{s},s}}$ such that $u,v,w,z\in V([A_{n_{k},k},D_{k}])$ (or $u,v,w,z\in V([E_{k},D_{k}])$ if $k\notin J_{\mathcal{Q}_{r}}$).
  • (R4)
    Assume $k\in J_{\mathcal{Q}_{r}}$. Set $V_{1,k}=V([E_{k},B_{1,k}]\setminus\{B_{1,k})\})$, $V_{i,k}=V([A_{i-1,k},B_{i,k}]\setminus\{B_{i,k})\})$ for all $i\in[n_{k}]\setminus\{1,n_{k}\}$, and $V_{n_{k},k}=V([A_{n_{k},k},D_{k}])$. For all $i\in[n_{k}-1]$, set $V_{i,k}\cap V_{i+1,k}=\{h_{i,k}\}$. By the construction of $\prec_{\mathcal{R}_{n_{s},s}}$ (see in particular Remark 3.5), for all $u\in V_{i,k}$ and for all $v\in V_{i+1,k}\setminus\{h_{i,k}\}$ we have $x_{u}\prec_{\mathcal{R}_{n_{s},s}}x_{v}$.
Furthermore, by the proof of Lemma 3.4, the order $\prec_{\mathcal{R}_{n_{s},s}}$ can be defined as the lexicographic order induced by a total on the variables $x_{a}$ with $a\in V(\mathcal{R}_{n_{s},s})$. Now we want to define a monomial order $\prec_{\mathcal{Q}_{r}}$ on $S_{\mathcal{Q}_{r}}$ that extends the order $\prec_{\mathcal{R}_{n_{s},s}}$ defined in $S_{\mathcal{R}_{n_{s},s}}$. Let $Y=\{b_{i,k}^{(2)}\in V(\mathcal{Q}_{r})\mid k\in J_{\mathcal{Q}_{r}},i\in[n_{k}]\}$ and observe that $V(\mathcal{Q}_{r})=Y\cup V(\mathcal{R}_{n_{s},s})$ (disjoint union). Set $<_{Y}$ to be any total order on the set $Y$. Let $<_{\mathcal{Q}_{r}}$ be the total order on the variables of $S_{\mathcal{Q}_{r}}$ defined in the following way:
$x_{a}<_{\mathcal{Q}_{r}}x_{b}\Longleftrightarrow\left\{\begin{array}[]{l}a,b\in V(\mathcal{R}_{n_{s},s})\text{ and }x_{a}\prec_{\mathcal{R}_{n_{s},s}}x_{b}\\ a\in V(\mathcal{Q}_{r})\text{ and }b\in Y\\ a,b\in Y\text{ and }a<_{Y}b\end{array}\right.$
Set $\prec_{\mathcal{Q}_{r}}$ the lexicographic order on $S_{\mathcal{Q}_{r}}$ induced by the total order $<_{\mathcal{Q}_{r}}$. Trivially, the monomial order $\prec_{\mathcal{Q}_{r}}$ satisfies claim (2). We want to prove also that claim (1) holds. Firstly, we prove it for the ideal $I_{\mathcal{Q}_{r}}$.
Let $f,g$ be inner 2-minors of $I_{\mathcal{Q}_{r}}$ and denote by $S(f,g)$ the $S$-polynomial of $f$ and $g$. It is not difficult to see that if $f,g\in I_{\mathcal{R}_{n_{s},s}}$, since $S(f,g)$ reduces to 00 in $I_{\mathcal{R}_{n_{s},s}}$, then $S(f,g)$ reduces to 00 also in $I_{\mathcal{Q}_{r}}$. So, suppose $f\notin I_{\mathcal{R}_{n_{s},s}}$. Hence the inner interval related to $f$ contains the cell $B_{i,k}$ for some $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$. Then, referring to Figure 19, up to reflection an rotations $f$ is one of the following:
  • (a)
    $f_{j,i,k}=\pm(x_{b_{i,k}^{(2)}}x_{h_{j}}-x_{v_{j}}x_{h_{0}})$ for $j\in[\ell]$ with $\operatorname{in}(f_{j,i,k})=x_{b_{i,k}^{(2)}}x_{h_{j}}$.
  • (b)
    $f_{b,i,k}=\pm(x_{b_{i,k}^{(2)}}x_{a_{i,k}^{(1)}}-x_{a_{i,k}^{(2)}}x_{b_{i,k}^{(1)}})$ with $\operatorname{in}(f_{b,i,k})=x_{b_{i,k}^{(2)}}x_{a_{i,k}^{(1)}}$.
Figure 19. In the figures above $\ell\geq 3$ and the cells $B_{i-1,k}$, $B_{i,k}$ and $B_{1,k}$ above belong to $\mathcal{Q}_{r}$ but not to $\mathcal{R}_{n_{s},s}$
For the polynomial $g$, we need to examine only the case $\gcd(\operatorname{in}(f),\operatorname{in}(g))\neq 1$. Then it suffices to examine the following cases:
  • Case 1)
    $f=f_{j,i,k}$ with $j\in[\ell]$ and $g=\pm(x_{v_{j_{1}}}x_{h_{j_{2}}}-x_{v_{j_{2}}}x_{h_{j_{1}}})$ with $j_{1},j_{2}\in[\ell]$ or $g=f_{l,i,k}$ with $l\in[\ell]\setminus\{j\}$. In this case $S(f,g)$ reduces to 00 by [6, Remark 1].
  • Case 2)
    $f=f_{1,i,k}$ and $g=f_{b,i,k}$. Also in this case $S(f,g)$ reduces to 00 by [6, Remark 1].
  • Case 3)
    $f=f_{j,i,k}$ with $j\in[\ell]\setminus\{1\}$ and $g=f_{b,i,k}$. In this case, by [6, Lemma 2 (condition 1)], we obtain that $S(f,g)$ reduces to 00 if both the following properties holds:
    • 3a)
      $x_{a_{i,k}^{(1)}},x_{h_{0}},x_{v_{j}}\prec_{\mathcal{Q}_{r}}x_{a_{i,k}^{(2)}}$.
    • 3b)
      $x_{h_{1}},x_{v_{j}}\prec_{\mathcal{Q}_{r}}x_{h_{j}}$ or $x_{h_{1}},x_{v_{j}}\prec_{\mathcal{Q}_{r}}x_{v_{1}}$.
    We prove 3a). Consider that, by observation (R4), $x_{v_{j}}\prec_{\mathcal{Q}_{r}}x_{a_{i,k}^{(2)}}$, $x_{h_{1}}\prec_{\mathcal{Q}_{r}}x_{a_{i,k}^{(1)}}$ and $x_{h_{1}}\prec x_{h_{0}}$. Moreover, considering $g^{\prime}=\pm(x_{a_{i,k}^{(2)}}x_{h_{1}}-x_{a_{i,k}^{(1)}}x_{h_{0}})$, then $\operatorname{in}(g^{\prime})=x_{a_{i,k}^{(2)}}x_{h_{1}}$ by observation (R1) and conditions (C1)-(C2) of the monomial order $\prec_{\mathcal{R}_{n_{s},s}}$.
    Assuming that $x_{a_{i,k}^{(2)}}\prec_{\mathcal{Q}_{r}}x_{a_{i,k}^{(1)}}$ or $x_{a_{i,k}^{(2)}}\prec_{\mathcal{Q}_{r}}x_{h_{0}}$, we obtain $\operatorname{in}(g^{\prime})=x_{a_{i,k}^{(1)}}x_{h_{0}}$, that is a contradiction. Hence $x_{a_{i,k}^{(1)}}\prec_{\mathcal{Q}_{r}}x_{a_{i,k}^{(2)}}$ and $x_{h_{0}}\prec_{\mathcal{Q}_{r}}x_{a_{i,k}^{(2)}}$. So 3a) is proved.
    Let us prove that 3b) holds. Set $g^{\prime}=\pm(x_{v_{1}}x_{h_{j}}-x_{v_{j}}x_{h_{1}})$ and observe that $\operatorname{in}(g^{\prime})=x_{v_{1}}x_{h_{j}}$, by observation (R1) provided at the beginning. If $x_{h_{j}}\prec_{\mathcal{Q}_{r}}x_{v_{j}}$ and $x_{v_{1}}\prec_{\mathcal{Q}_{r}}x_{v_{j}}$ then $\operatorname{in}(g^{\prime})=x_{v_{j}}x_{h_{1}}$, a contradiction. We obtain the same if $x_{h_{j}}\prec_{\mathcal{Q}_{r}}x_{h_{1}}$ and $x_{v_{1}}\prec_{\mathcal{Q}_{r}}x_{h_{1}}$. Assume $x_{h_{j}}\prec_{\mathcal{Q}_{r}}x_{v_{j}}$ and $x_{v_{1}}\prec_{\mathcal{Q}_{r}}x_{h_{1}}$, then comparing $x_{h_{j}}$ and $x_{v_{1}}$ with respect to $\prec_{\mathcal{Q}_{r}}$, we obtain that $x_{h_{j}},x_{v_{1}}\prec_{\mathcal{Q}_{r}}x_{v_{j}}$ or $x_{h_{j}},x_{v_{1}}\prec_{\mathcal{Q}_{r}}x_{h_{1}}$, that is, $\operatorname{in}(g^{\prime})=x_{v_{j}}x_{h_{1}}$, a contradiction. We obtain the same if $x_{h_{j}}\prec_{\mathcal{Q}_{r}}x_{h_{1}}$ and $x_{v_{1}}\prec_{\mathcal{Q}_{r}}x_{v_{j}}$. So, property 3b) holds.
  • Case 4)
    $f=f_{1,i,k}$ and $g=\pm(x_{h_{1}}x_{u}-x_{a_{i,k}^{(1)}}x_{w})$ with $u,w\in V([A_{i,k},B_{i+1,k}]$ if $i\neq n_{k}$, or $u,w\in V([A_{i,k},D_{k}]$ if $i=n_{k}$, and $\operatorname{in}(g)=x_{h_{1}}x_{u}$. In this case, recall that $x_{z}\prec_{\mathcal{Q}_{r}}x_{b_{i,k}^{(2)}}$ for all $z\notin Y$. Hence, if $u\notin Y$ or $u\in Y$ with $x_{u}\prec_{\mathcal{Q}_{r}}x_{b_{i,k}^{(2)}}$ then $S(f,g)$ reduces to 00 by [6, Lemma 4 (condition 3)], while if $u\in Y$ and $x_{b_{i,k}^{(2)}}\prec_{\mathcal{Q}_{r}}x_{u}$ then $S(f,g)$ reduces to 00 by [6, Lemma 4 (condition 1)].
  • Case 5)
    $f=f_{\ell,i,k}$ and $g=\pm(x_{h_{\ell}}x_{b_{i-1,k}^{(2)}}-x_{b_{i-1,k}^{(1)}}x_{h_{\ell-1}})$. In this case observe that $\operatorname{in}(g)=x_{h_{\ell}}x_{b_{i-1,k}^{(2)}}$ and we can argue that $S(f,g)$ reduces to 00 using the same argument of Case 4).
The previous cases allow us to conclude that $I_{\mathcal{Q}_{r}}$ has a quadratic Gröbner basis with respect to $\prec_{\mathcal{Q}_{r}}$. It is not difficult to see that the same arguments hold also to prove that $I_{\mathcal{R}_{i,k}}$ and $I_{\mathcal{R}_{i,k}^{\prime}}$ have a squarefree quadratic Gröbner basis with respect to $\prec_{\mathcal{Q}_{r}}$, for all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$. So we can conclude. ∎
The following Lemma, which we need later, is formulated as a general result on a collection of cells. For this reason, only for this result, we use an independent notation, referring to Figure 20.
Lemma 4.11.
Let $\mathcal{P}$ be a collection of cells having a subconfiguration as in Figure 20.
Figure 20. Configuration examined in Lemma 4.11.
With reference to that figure, denote $\mathcal{Q}=\mathcal{P}\setminus\{A_{1}\}$ and $\mathcal{R}=\mathcal{P}\setminus\{A_{1},\ldots,A_{r},B\}$ and assume $\{a,c,a_{1},\ldots,a_{r},b_{1},\ldots,b_{r-1}\}\cap V(\mathcal{R})=\emptyset$. Suppose that there exists a term order $\prec$ such that $I_{\mathcal{P}}$, $I_{\mathcal{Q}}$ and $I_{\mathcal{R}}$ have a quadratic Gröbner basis with respect to $\prec$ and assume $x_{a}x_{c},x_{a}x_{b_{i}}\in\operatorname{in}(I_{\mathcal{P}})$ for all $i\in[r]$. Then
  1. (1)
    $\mathrm{HP}_{K[\mathcal{P}]}(t)=\mathrm{HP}_{K[\mathcal{Q}]}(t)+\frac{t}{(1-t)^{r}}\mathrm{HP}_{K[\mathcal{R}]}(t)$.
  2. (2)
    $\operatorname{depth}(K[\mathcal{P}])\geq\min\{\operatorname{depth}(K[\mathcal{Q}]),\operatorname{depth}(K[\mathcal{R}])+r\}$
Furthermore, if $\mathcal{R}$ is simple then:
  • $\mathrm{HP}_{K[\mathcal{P}]}(t)=\mathrm{HP}_{K[\mathcal{Q}]}(t)+\frac{t\cdot h_{K[\mathcal{R}]}(t)}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|}}$.
  • $\operatorname{depth}(K[\mathcal{P}])\geq\min\{\operatorname{depth}(K[\mathcal{Q}]),|V(\mathcal{P})|-|\mathcal{P}|\}$.
Proof.
Consider the following exact sequence:
$0\longrightarrow\frac{S_{\mathcal{P}}}{(\mathrm{in}(I_{\mathcal{P}}):x_{a})}\longrightarrow\frac{S_{\mathcal{P}}}{\mathrm{in}(I_{\mathcal{P}})}\longrightarrow\frac{S_{\mathcal{P}}}{(\mathrm{in}(I_{\mathcal{P}}),x_{a})}\longrightarrow 0$
Since $I_{\mathcal{P}},I_{\mathcal{Q}}$ and $I_{\mathcal{R}}$ have quadratic Gröbner basis with respect to the same term order $\prec$ and considering our assumptions, we have:
  • $\operatorname{in}(I_{\mathcal{P}})=\operatorname{in}(I_{\mathcal{Q}})+(x_{a}x_{c})+\sum_{i=1}^{r}(x_{a}x_{b_{i}})$.
  • $\operatorname{in}(I_{\mathcal{P}})=\operatorname{in}(I_{\mathcal{R}})+(x_{a}x_{c})+\sum_{i=1}^{r}(x_{a}x_{b_{i}})+\max_{\prec}(\{x_{a_{i}}x_{b_{j}},x_{a_{j}}x_{b_{i}}\mid x_{a_{i}}x_{b_{j}}-x_{a_{j}}x_{b_{i}}\in I_{\mathcal{P}},i,j\in[r],i<j\})+\max_{\prec}(\{x_{c}x_{z_{j}},x_{b_{1}}x_{w_{j}}\mid x_{c}x_{z_{j}}-x_{b_{1}}x_{w_{j}}\in I_{\mathcal{P}}\}).$
As consequence, we obtain:
  • $(\operatorname{in}(I_{\mathcal{P}}),x_{a})=(\operatorname{in}(I_{\mathcal{Q}}),x_{a})$ and in particular $\frac{S_{\mathcal{P}}}{(\operatorname{in}(I_{\mathcal{P}}),x_{a})}\cong\frac{S_{\mathcal{Q}}}{\operatorname{in}(I_{\mathcal{Q}})}$.
  • $\operatorname{in}(I_{\mathcal{P}}):x_{a}=\operatorname{in}(I_{\mathcal{R}})+(x_{c})+\sum_{i=1}^{r}(b_{i})$, and in particular $\frac{S_{\mathcal{P}}}{(\mathrm{in}(I_{\mathcal{P}}):x_{a})}\cong\frac{S_{\mathcal{R}}}{(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})}\otimes K[x_{a},x_{a_{1}},\ldots,x_{a_{r}}]$
Observe that, since by assumptions $I_{\mathcal{P}}$ has a quadratic Gröbner basis with respect to $\prec$, $x_{b_{r}}$ does not divide any generator of $\operatorname{in}(I_{\mathcal{P}})$ different by $x_{a_{i}}x_{b_{r}}$, for $i\in[r-1]$. In fact, assume $f=x_{b_{r}}x_{v}-x_{z}x_{w}\in I_{\mathcal{P}}$ for some $v,w,z\in V(\mathcal{P})$ such that $v\in V(\mathcal{P})\setminus\{a,a_{1},\ldots,a_{r}\}$ and $\operatorname{in}(f)=x_{b_{r}}x_{v}$. So, considering $g=x_{a}x_{b_{r}}-x_{b}x_{a_{r}}\in I_{\mathcal{P}}$, we have $\gcd(\operatorname{in}(f),\operatorname{in}(g))\neq 0$. Without loss of generality, assume $z$ is in the same horizontal edge interval of $b_{r}$, while $w$ is in the same vertical edge interval of $b_{r}$. As consequence, assuming the $S$-polynomial of $f$ and $g$ reduces to zero, by [6, Lemma 6] we have $[z,a_{r}]$ or $[w,a]$ is an inner interval, a contradiction.
Therefore, if $y$ is another indeterminate and denoting by $S^{\prime}$ the polynomial ring such that $S^{\prime}\otimes K[x_{b_{r}}]\cong S_{\mathcal{R}}$, we can express
$\frac{S_{\mathcal{R}}}{\operatorname{in}(I_{\mathcal{R}})}\cong\frac{S^{\prime}}{\operatorname{in}(I_{\mathcal{R}})}\otimes K[x_{b_{r}}]\cong\frac{S_{\mathcal{R}}}{(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})}\otimes K[y]$
Hence, by Proposition 1.8, we have $\mathrm{HP}_{S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})}(t)=(1-t)\cdot\mathrm{HP}_{S_{\mathcal{R}}/\operatorname{in}(I_{\mathcal{R}})}(t)$.
Therefore, by Proposition 1.5 and Proposition 1.7 we obtain $\mathrm{HP}_{K[\mathcal{P}]}(t)=\mathrm{HP}_{K[\mathcal{R}]}(t)+\frac{t}{(1-t)^{r}}\mathrm{HP}_{K[\mathcal{R}]}(t)$, that is the first claim.
For the second claim, by Proposition 1.10 we have $\operatorname{depth}(S_{\mathcal{P}}/\operatorname{in}(I_{\mathcal{P}}))\geq\{\operatorname{depth}(S_{\mathcal{Q}}/\operatorname{in}(I_{\mathcal{Q}})),\operatorname{depth}(S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})\otimes K[x_{a},x_{a_{1}},\ldots,x_{a_{r}}])\}$. Hence, by Propositions 1.6 and 1.9 we obtain $\operatorname{depth}(K[\mathcal{P}])\geq\min\{\operatorname{depth}(K[\mathcal{Q}]),\operatorname{depth}(K[\mathcal{R}])+r\}$.
Finally, for the last two claims, if $\mathcal{R}$ is a simple collection of cells, by Proposition 1.3 we have that $\dim(K[\mathcal{R}])=\operatorname{depth}(K[\mathcal{R}])=|V(\mathcal{R})|-|\mathcal{R}|$. Moreover $|V(\mathcal{R})|=|V(\mathcal{P})|-(2r+1)$ and $|\mathcal{R}|=|\mathcal{P}|-(r+1)$, that is, $|V(\mathcal{R})|-|\mathcal{R}|+r=|V(\mathcal{P})|-|\mathcal{P}|$. From this fact we can easily conclude. ∎
Lemma 4.12.
Let $\mathcal{P}$ be a non-prime closed path polyomino and $\mathcal{Q}_{r}$ be the collection of cells as in Set-up 4.2. Consider the framework of Set-up 4.8. Then $\mathrm{depth}(K[\mathcal{Q}_{r}])\geq|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$.
Moreover, if $s=\max J_{\mathcal{Q}_{r}}$, then:
$\mathrm{HP}_{K[\mathcal{Q}_{r}]}(t)=\mathrm{HP}_{K[\mathcal{R}_{n_{s},s}]}(t)+\frac{t}{(1-t)^{|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|}}\left(\sum_{k\in J}\sum_{i=1}^{n_{k}}h_{K[\mathcal{R}^{\prime}_{i,k}]}(t)\right)$
Proof.
Let $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$ and denote $h=b_{i,k}^{(2)}$. Consider, respectively, the following collection of cells and cell interval:
$\mathcal{S}_{i,k}:=\left\{\begin{array}[]{ll}\mathcal{R}_{i-1,k}&\mbox{if $i\neq 1$}\\ \mathcal{R}_{n_{k-1},k-1}&\mbox{if $i=1$ and $k\neq\min J_{\mathcal{Q}_{r}}$}\\ \mathcal{Q}_{r}&\mbox{otherwise}\end{array}\right.\qquad\mbox{and}\qquad\mathcal{E}_{i,k}:=\left\{\begin{array}[]{ll}[A_{i-1,k};B_{i,k}]&\mbox{if}\ i\neq 1\\ \left[E_{k};B_{1,k}\right]&\mbox{if $i=1$}\end{array}\right.$
Observe that, up to rotations and reflections, the configuration of cells given by $\mathcal{E}_{i,k}\cup\{A_{i,k}\}$ determine a sub-configuration of $\mathcal{S}_{i,k}$ as in Figure 20. Moreover, observe that $\mathcal{S}_{i,k}\setminus\{B_{i,k}\}=\mathcal{R}_{i,k}$ and $\mathcal{S}_{i,k}\setminus(\mathcal{E}_{i,k}\cup\{A_{i,k}\})=\mathcal{R}_{i,k}^{\prime}$. Furthermore, by Lemma 4.10, we have that $\mathcal{S}_{i,k}$ satisfies all conditions of Lemma 4.11 and $\mathcal{R}_{i,k}^{\prime}$ is a simple collection of cells. Therefore $\mathrm{HP}_{K[\mathcal{S}_{i,k}]}(t)=\mathrm{HP}_{K[\mathcal{R}_{i,k}]}(t)+\frac{t\cdot h_{K[\mathcal{R}^{\prime}_{i,k}]}(t)}{(1-t)^{|V(\mathcal{S}_{i,k})|-|\mathcal{S}_{i,k}|}}$. and $\operatorname{depth}(K[\mathcal{S}_{i,k}])\geq\min\{\operatorname{depth}(K[\mathcal{R}_{i,k}]),|V(\mathcal{S}_{i,k})|-|\mathcal{S}_{i,k}|\}$.
So, denoting $J_{\mathcal{Q}_{r}}=\{k_{1},\ldots,k_{s}\}$, starting from $\mathcal{Q}_{r}$ we obtain:
$\mathrm{HP}_{K[\mathcal{Q}_{r}]}(t)$ $=\mathrm{HP}_{K[\mathcal{R}_{1,k_{1}}]}(t)+\frac{t\cdot h_{K[\mathcal{R}_{1,k_{1}}^{\prime}]}(t)}{(1-t)^{|V(\mathcal{Q}_{r})-|\mathcal{Q}_{r}|}}=$
$=\mathrm{HP}_{K[\mathcal{R}_{2,k_{1}}]}(t)+\frac{t}{(1-t)^{|V(\mathcal{Q}_{r})-|\mathcal{Q}_{r}|}}\left(h_{K[\mathcal{R}_{1,k_{1}}^{\prime}]}(t)+h_{K[\mathcal{R}_{2,k_{1}}^{\prime}]}(t)\right)=$
==$\cdots$== ⋯ =
$=\mathrm{HP}_{K[\mathcal{R}_{n_{k_{s}},k_{s}}]}(t)+\frac{t}{(1-t)^{|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|}}\left(\sum_{k\in J_{\mathcal{Q}_{r}}}\sum_{i=1}^{n_{k}}h_{K[\mathcal{R}^{\prime}_{i,k}]}(t)\right)$
Furthermore,
$\operatorname{depth}(K[\mathcal{Q}_{r}])\geq$ $\min\{\operatorname{depth}(K[\mathcal{R}_{1,k_{1}}]),|V(Q_{r})|-|\mathcal{Q}_{r}|\}.$
$\operatorname{depth}(K[\mathcal{R}_{1,k_{1}}])\geq$ $\min\{\operatorname{depth}(K[\mathcal{R}_{2,k_{1}}]),|V(\mathcal{R}_{1,k_{1}})|-|\mathcal{R}_{1,k_{1}}|\}.$
$\cdots$
$\operatorname{depth}(K[\mathcal{S}_{n_{k_{s}},k_{s}}])\geq$ $\min\{\operatorname{depth}(K[\mathcal{R}_{n_{k_{s}},k_{s}}]),|V(\mathcal{S}_{n_{k_{s}},k_{s}})|-|\mathcal{S}_{n_{k_{s}},k_{s}}|\}.$
Observing that $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{R}_{i,k})|-|\mathcal{R}_{i,k}|$ for all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$ (in fact $\mathcal{S}_{i,k}$ differs from $\mathcal{R}_{i,k}$ only for a vertex and cells), we obtain $\operatorname{depth}(K[\mathcal{Q}_{r}])\geq\min\{\operatorname{depth}(K[\mathcal{R}_{n_{k_{s}},k_{s}}]),|V(Q_{r})|-|\mathcal{Q}_{r}|\}$. Finally, observe that $\mathcal{R}_{n_{k_{s}},k_{s}}$ is a zig-zag collection. Hence, by Theorem 3.6, we have that $K[\mathcal{R}_{n_{k_{s}},k_{s}}]$ is a Cohen-Macaulay domain of dimension $|V(\mathcal{R}_{n_{k_{s}},k_{s}})|-|\mathcal{R}_{n_{k_{s}},k_{s}}|=|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$, obtaining $\operatorname{depth}(K[\mathcal{Q}_{r}])\geq|V(Q_{r})|-|\mathcal{Q}_{r}|$
Theorem 4.13.
Let $\mathcal{P}$ be a closed path with a zig-zag walk. Then $K[\mathcal{P}]$ is Cohen-Macaulay of dimension $|V(\mathcal{P})|-|\mathcal{P}|$.
Proof.
By Lemma 4.5 we have $\mathrm{depth}(K[\mathcal{P}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{r}]),|V(\mathcal{P})|-|\mathcal{P}|\}$ and by Lemma 4.12 $\mathrm{depth}(K[\mathcal{Q}_{r}])\geq|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$. Moreover, by the proof of Lemma 4.5 we have also $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{P})|-|\mathcal{P}|$. Hence, $\operatorname{depth}(K[\mathcal{P}])\geq|V(\mathcal{P})|-|\mathcal{P}|$. By [10, Theorem 3.4] we also know that $\dim(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$, so $K[\mathcal{P}]$ is Cohen-Macaulay with the given dimension. ∎
Remark 4.14.
Observe that the result in Theorem 4.13 holds also for all collection of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ described in Set-up 4.2 and Set-up 4.8. In fact, if $\mathcal{R}$ is one of such configurations, by [3, Lemma 3.3] and by construction, we can argue that there exists a cell $A$ in $\mathcal{R}$ such that $\mathcal{R}\setminus\{A\}$ is simple. So, by [8, Proposition 4.9] (and [37, Corollary 3.1.7]) we have $\dim(K[\mathcal{R}])=|V(\mathcal{R})|-|\mathcal{R}|$. Therefore, we can obtain this claim using the same arguments developed in this section.
Theorem 4.15.
Let $\mathcal{P}$ a closed path with a zig-zag walk. Then the $h$-polynomial of $K[\mathcal{P}]$ is the rook polynomial of $\mathcal{P}$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
Proof.
We start by showing how the claim holds for the collection of cells $\mathcal{Q}_{r}$ as in Set-up 4.2. If $\mathcal{P}$ is a non-prime closed path polyomino not containing the configuration in Figure 14b then $\mathcal{Q}_{r}$ is a zig-zag collection, so $h_{K[\mathcal{Q}_{r}]}(t)$ is the rook polynomial by Corollary 3.8. Assume $\mathcal{P}$ contains the configuration in Figure 14b. Then, with reference to Set-up 4.8 and the proof of Lemma 4.12, for $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$ consider the collections of cells $\mathcal{S}_{i,k}$ and $\mathcal{R}_{i,k}^{\prime}$. In Lemma 4.12 we proved that
$\mathrm{HP}_{K[\mathcal{S}_{i,k}]}(t)=\mathrm{HP}_{K[\mathcal{R}_{i,k}]}(t)+\frac{t\cdot h_{K[\mathcal{R}^{\prime}_{i,k}]}(t)}{(1-t)^{|V(\mathcal{S}_{i,k})|-|\mathcal{S}_{i,k}|}}.$
Moreover, for all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$, observe that $\mathcal{R}_{i,k}^{\prime}$ is a simple collection of cells . So, by Corollary 2.5, $h_{K[\mathcal{R}^{\prime}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{R}^{\prime}_{i,k}$. Furthermore, for all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$, by the definition of $\mathcal{S}_{i,k}$ and by the proof of Lemma 4.12 we have $|V(\mathcal{S}_{i,k})|-|\mathcal{S}_{i,k}|=|V(\mathcal{R}_{i,k})|-|\mathcal{R}_{i,k}|=|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$, and by Remark 4.14 we have $\dim(K[\mathcal{R}_{i,k}])=|V(\mathcal{R}_{i,k})|-|\mathcal{R}_{i,k}|=|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$. As a consequence, we have:
$\mathrm{HP}_{K[\mathcal{S}_{i,k}]}(t)=\frac{h_{K[\mathcal{R}_{i,k}]}(t)+t\cdot h_{K[\mathcal{R}^{\prime}_{i,k}]}(t)}{(1-t)^{|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|}}.$
Let $s_{1}=\max J_{\mathcal{Q}_{r}}$ and $s_{2}=\min J_{\mathcal{Q}_{r}}$. Following the collections of cells involved, observe that, $\mathcal{R}_{n_{s},s}$ is a zig-zag collection and and $\mathcal{S}_{1,s_{1}}=\mathcal{Q}_{r}$. In particular, by Corollary 3.8, $h_{K[\mathcal{R}_{n_{s},s}]}(t)$ is the rook polynomial of $\mathcal{R}_{n_{s},s}$. Therefore, if $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$, it suffices to prove the following claim: if $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{R}_{i,k}$ then $h_{K[\mathcal{S}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{S}_{i,k}$.
So, assume $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{R}_{i,k}$. For all non-negative integer $k$, denote by $r^{(1)}_{k}$ and $r^{(2)}_{k}$, respectively, the coefficient of $t^{k}$ in the polynomials $h_{K[\mathcal{R}_{i,k}]}(t)$ and $h_{K[\mathcal{R}_{i,k}^{\prime}]}(t)$. Denote by $r_{k}=r^{(1)}_{k}+r^{(2)}_{k-1}$ (with the convention $r_{-1}^{(2)}=0$). Observe that $r_{k}$ is the coefficient of $t^{k}$ in $h_{K[\mathcal{S}_{i,k}]}(t)$. Hence, it suffices to show that $r_{k}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$. Recall that, as explained in the proof of Lemma 4.12, we can refer to Figure 20 and express $\mathcal{R}_{i,k}=\mathcal{S}_{i,k}\setminus\{A_{1}\}$ and $\mathcal{R}_{i,k}^{\prime}=\mathcal{S}_{i,k}\setminus(\{B,A_{1},\ldots,A_{r}\})$. So, consider that we can write $r_{k}=r_{1}+r_{2}$, where $r_{1}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$ having no rook placed in $A_{1}$ and $r_{2}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$ with a rook placed in $A_{1}$. It is not difficult to argue that $r_{1}$ is also equal to the number of $k$-rook configurations of $\mathcal{R}_{i,k}$, while $r_{2}$ is equal the number of $(k-1)$-rook configurations of $\mathcal{R}_{i,k}^{\prime}$. Therefore, $r_{1}=r_{k}^{(1)}$ and $r_{2}=r_{k-1}^{(2)}$, obtaining that $r_{k}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$. As consequence, $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{S}_{i,k}$, proving our claim.
So, the $h$-polynomial of $K[\mathcal{Q}_{r}]$ is the rook polynomial of $\mathcal{Q}_{r}$. In order to obtain our results for $\mathcal{P}$, let us consider, with reference to Set-up 4.2, the collections of cells $\mathcal{Q}_{i}$ and $\mathcal{Q}_{i}^{\prime}$ for all $i\in[r]$. In Lemma 4.5 we showed that $|V(\mathcal{P})|-|\mathcal{P}|=|V(\mathcal{Q}_{i})|-|\mathcal{Q}_{i}|=|V(\mathcal{Q}_{i}^{\prime})|-|\mathcal{Q}_{i}^{\prime}|+1$ for all $i\in[r]$ and we also proved the following:
$\mathrm{HP}_{K[\mathcal{Q}_{i-1}]}(t)=\mathrm{HP}_{K[\mathcal{Q}_{i}]}(t)+\frac{t}{1-t}\mathrm{HP}_{K[\mathcal{Q}^{\prime}_{i}]}(t)$
Moreover, for all $i\in[r]$ consider that $\mathcal{Q}_{i}^{\prime}$ is a simple collection of cells, in particular $\dim(K[\mathcal{Q}_{i}^{\prime}])=|V(\mathcal{Q}_{i}^{\prime})|-|\mathcal{Q}_{i}^{\prime}|$ and by Corollary 2.5 we have that $h_{K[\mathcal{Q}_{i}^{\prime}]}$ is the rook polynomial of $\mathcal{Q}_{i}^{\prime}$. So, considering also Remark 4.14, by the previous expression we obtain:
$\mathrm{HP}_{K[\mathcal{Q}_{i-1}]}(t)=\frac{h_{K[\mathcal{Q}_{i}]}(t)+t\cdot h_{K[\mathcal{Q}_{i}^{\prime}]}}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|}}$
Following the collections of cells involved and observing that $\mathcal{Q}_{0}=\mathcal{P}$, we can obtain our result if for all $i\in[r]$ we prove the following claim: if $h_{K[\mathcal{Q}_{i}]}(t)$ is the rook polynomial of $\mathcal{Q}_{i}$ then $h_{K[\mathcal{Q}_{i-1}]}(t)$ is the rook polynomial of $\mathcal{Q}_{i-1}$. But it is not difficult to see that this fact can be proved by the same argument used before, so we can conclude. ∎
Remark 4.16.
It is not difficult to see that the claim provided for $\mathcal{P}$ in Theorem 4.15, following its proof, holds also for all collection of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ described in Set-up 4.2 and Set-up 4.8.
Proposition 4.17.
Let $\mathcal{P}$ be a closed path with a zig-zag walk. Then $K[\mathcal{P}]$ is not Gorenstein.
Proof.
Denote by $h(t)=\sum_{k=1}^{s}h_{k}t^{k}$ the $h$-polynomial of $K[\mathcal{P}]$. From Theorem 4.15, $h(t)$ is the rook polynomial of $\mathcal{P}$ and $s=r(\mathcal{P})$. Suppose by contradiction that $K[\mathcal{P}]$ is Gorenstein, so it follows from [37, Corollary 5.3.10] that $h_{k}=h_{s-k}$ for all $k\in\{0,\dots,s\}$. In particular we have that $h_{s-1}=h_{1}=|\mathcal{P}|$ and also $h_{s}=h_{0}=1$. We examine the following three cases.
  1. (1)
    Let us start showing that $|I|\leq 3$ for every maximal cell interval $I$ of $\mathcal{P}$. In fact, suppose there exists a maximal interval $I=[A,B]$ with $|I|\geq 4$. Since $|I|\geq 4$, then we can consider two distinct cells $C,D\in I\setminus\{A,B\}$. Note that every $s$-rook configuration of $\mathcal{P}$ has a rook in a cell of $I\setminus\{A,B\}$. Let $\mathcal{T}$ be an $s$-rook configuration of $\mathcal{P}$ with a rook placed in $C$. Therefore, we obtain a new $s$-rook configuration of $\mathcal{P}$, moving the rook in $C$ into $D$, so $h_{s}\geq 2>h_{0}=1$, that is a contradiction. Hence, $|I|\leq 3$ for every maximal interval $I$ of $\mathcal{P}$, as claimed.
  2. (2)
    Now, suppose that $\mathcal{P}$ contains a sub-polyomino as in Figure 14b. Arguing as done before, there exists an $s$-rook configuration $\mathcal{U}$ in $\mathcal{P}$ having a rook placed in $W$ and another one in $Z$. Moreover, we can get from $\mathcal{U}$ another $s$-rook configuration in $\mathcal{P}$, moving the rook in $\mathcal{U}$ from $W$ to $X$. Therefore the same previous contradiction arises.
  3. (3)
    From the previous cases it follows that every maximal interval of $\mathcal{P}$ has rank less or equal to $3$ and $\mathcal{P}$ does not contain any sub-polyomino as in Figure 14b. Hence the only possibility is that $\mathcal{P}$ is as in Figure 21. In such a case, Macaulay2 computations show that the reduced Hilbert series of $\mathcal{P}$ is given by
    $\frac{1+16\,t+100\,t^{2}+308\,t^{3}+486\,t^{4}+376\,t^{5}+134\,t^{6}+20\,t^{7}+t^{8}}{\left(1-t\right)^{16}}$
    so $h_{7}=20\neq 16=h_{1}$, which is a contradiction.
    Figure 21. A closed path.
In all three cases we get a contradiction, so we conclude that $K[\mathcal{P}]$ cannot be Gorenstein. ∎
In the following result we gather, as a summary, the main algebraic properties of $K[\mathcal{P}]$ when $\mathcal{P}$ is a closed path polyomino.
Theorem 4.18.
Let $\mathcal{P}$ be a closed path polyomino. Then:
  1. (1)
    $K[\mathcal{P}]$ is Cohen-Macaulay with Krull dimension $|V(\mathcal{P})|-|\mathcal{P}|$; moreover, if $\mathcal{P}$ does not contain any zig-zag walk, then $K[\mathcal{P}]$ is also a normal domain.
  2. (2)
    Let
    $\mathrm{HP}_{K[\mathcal{P}]}(t)=\frac{h(t)}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|}}$
    be the reduced Hilbert series of $K[\mathcal{P}]$. Then $h(t)$ is the rook polynomial of $K[\mathcal{P}]$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
  3. (3)
    $K[\mathcal{P}]$ is Gorenstein if and only if $\mathcal{P}$ consists of maximal blocks of rank three.
Proof.
We need to discuss just two cases. In the first one assume that $\mathcal{P}$ has a zig-zag walk. Then (1) and (2) follow from Theorems 4.13 and 4.15. Concerning (3), we know that $K[\mathcal{P}]$ is not Gorenstein from Proposition 4.17. Moreover, by [3, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration, so any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 14a or 14b (up to reflections or rotations). Hence, we argue that if $\mathcal{P}$ has a zig-zag walk then it contains maximal blocks of length two. This means that (3) holds if $\mathcal{P}$ has zig-zag walks. Now, in the second case suppose that $\mathcal{P}$ does not contain a zig-zag walk. The claims (1) and (2) are already proved in [7, Theorem 5.5] and (3) in [7, Theorem 5.7]. In conclusion, the theorem is completely proved. ∎
5. Some open questions
We conclude this paper with some considerations and insights for future research. About the Cohen-Macaulayness of the inner $2$-minor ideals of collections of cells, it is known that if $\mathcal{P}$ is a simple collection of cells then $K[\mathcal{P}]$ is Cohen-Macaulay (see Proposition 1.3). Other classes of non-simple polyominoes whose coordinate ring is Cohen-Macaulay are discussed in [6], [21] and [26], although in these cases the examined classes consist just of prime polyominoes. In that cases, it is sufficient to study the Gröbner basis of the polyomino ideal by two well-known theorems of Sturmfels ([36]) and of Hochster ([2]). One of the main goal of this work is to investigate some classes of non-simple collections of cells whose coordinate ring is not a domain but it is Cohen-Macaulay. In particular, we proved that the zig-zag collections and the closed paths with zig-zag walks are Cohen-Macaulay, as well as the collections of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ introduced respectively in Set-up 4.2 and Set-up 4.8. Also, the coordinate ring of the collection of cells in Example 3.7 is Cohen-Macaulay. Nowadays, an example of a collection of cells having a not Cohen-Macaulay coordinate ring is still unknown. Therefore, the following general question naturally arises.
Question 5.1.
Let $\mathcal{P}$ be a collection of cells. Then $K[\mathcal{P}]$ is Cohen-Macaulay?
If $\mathcal{P}$ is a zig-zag collection we proved in Corollary 3.9(1) a necessary condition for $K[\mathcal{P}]$ to be Gorenstein. Motivated by some examples tested in Macaulay2, we ask if that condition is also sufficient, as mentioned in Remark 3.10. So, we formulate the following question:
Question 5.2.
Let $\mathcal{P}$ be a zig-zag collection supported by the intervals $I_{1},\dots,I_{\ell}$. Let $\mathcal{P}_{i}$ be the collection of cells related to $I_{i}$ and suppose that $\mathcal{P}_{i}$ is a square for all $i\in[\ell]$. Then, is $K[\mathcal{P}]$ Gorenstein?
Finally, in Theorem 4.18 we provide for the class of closed path polyominoes a description of the main algebraic properties of the related coordinate ring, characterizing them combinatorially. These properties are: be a domain, Hilbert-Poincaré series, regularity, be Cohen-Macaulay and Gorenstein. Recently in [32], the properties of Levelness and Pseudo-Gorenstein have been investigated in polyomino ideals for the first time, in particular for the class of path polyominoes (that are simple polyominoes). So another possible question is the following:
Question 5.3.
Let $\mathcal{P}$ be a closed path polyomino. Is it possible to characterize, by the combinatorics of $\mathcal{P}$, when $K[\mathcal{P}]$ is Level and Pseudo-Gorenstein?

### v3

@@
-On algebraic properties of some non-prime ideals of collections of cells
+On Cohen-Macaulay non-prime collections of cells
 Abstract.
-In this work, we consider a family of coordinate rings obtained as a quotient ring of a binomial ideal associated to a collection of cells.
-The algebraic properties of Cohen-Macaulayness, Gorensteiness and the Hilbert-Poincaré series have been studied in literature only for some classes of collections of cells whose related binomial ideals are prime.
-We study these algebraic properties for some classes of collections of cells whose coordinate rings are not domain, considering a new strategy involving the combinatorics of the collections of cells and a recent result of Conca and Varbaro on the radical initial ideals.
+In this paper we investigate Cohen-Macaulayness, Gorensteinness and the Hilbert-Poincaré series for some classes of non-prime collections of cells.
+In particular, we show that all closed path polyominoes are Cohen-Macaulay and we characterize those that are Gorenstein.
 Key words and phrases:
@@
 Stanley ([34]).
-In 2012 a new topic emerged in this field by the work of A.
+In 2012 a new topic emerged in this field due to the work of A.
 A.
-Qureshi.
-In [28] she establishes a connection between collections of cells and Commutative Algebra, assigning to every collection of cells $\mathcal{P}$ the ideal of the inner 2-minors of $\mathcal{P}$, in a suitable polynomial ring $S_{\mathcal{P}}$.
-This ideal $I_{\mathcal{P}}$ is called the inner 2-minor ideal of $\mathcal{P}$ and $K[\mathcal{P}]=S_{\mathcal{P}}/I_{\mathcal{P}}$ is said the coordinate ring of $\mathcal{P}$.
+Qureshi ([29]).
+She established a connection between collections of cells and Commutative Algebra, assigning to every collection $\mathcal{P}$ of cells the ideal of the inner 2-minors of $\mathcal{P}$ in a suitable polynomial ring $S_{\mathcal{P}}$.
+This ideal $I_{\mathcal{P}}$ is called the inner 2-minor ideal of $\mathcal{P}$ and $K[\mathcal{P}]=S_{\mathcal{P}}/I_{\mathcal{P}}$ is said to be the coordinate ring of $\mathcal{P}$.
 In particular, if $\mathcal{P}$ is a polyomino, that is a collection of cells where the cells are joined edge by edge, then $I_{\mathcal{P}}$ is called the polyomino ideal of $\mathcal{P}$.
-The aim of the research consists in the studying of the main algebraic properties of $K[\mathcal{P}]$ depending on the shape of $\mathcal{P}$.
-Since 2012, many algebraic properties of the polyomino ideals have been studied and many open questions have arisen.
-For example, one of the most interesting is the classification of the polyominoes whose coordinate ring is a domain.
-As usual, we say that a collection of cells is prime if the related ideal is prime.
-In [18, 19, 30], it is proved that if $\mathcal{P}$ is simple, which means $\mathcal{P}$ has no hole, then $K[\mathcal{P}]$ is a domain.
-Nowadays, the study is devoted to non-simple polyominoes and, although there are several articles where it is discussed, a complete characterization is still unknown.
-We invite the reader to consult [3, 5, 18, 19, 21, 25, 26, 30, 33].
-Also, many other algebraic properties, like when $K[\mathcal{P}]$ is Cohen-Macaulay or Gorenstein, are known only for some specific polyominoes and they are discussed just for prime classes of collections of cells so far.
-Regarding the Cohen-Macaulayness, the demonstrative strategy is based on two well-known theorems of Sturmfels ([36, Chapter 13]) and Hochster ([2, Theorem 6.3.5]), which state that a toric ring whose defining ideal admits a squarefree initial ideal is a normal Cohen-Macaulay ring.
-See for instance [6, 21, 26, 30].
-No class of collections of cells or polyominoes, whose coordinate ring is a Cohen-Macaulay ring but not a domain, has been discussed.
-Motivated by these considerations, we investigate the Cohen-Macaulayness and the Hilbert-Poincaré series of the zig-zag collections, a class of non-prime collections of cells coming from the so-called zig-zag walks (see [25, Definition 3.2]), and of the closed paths, a kind of non-simple polyominoes introduced in [3].
-Taking advantage of some recent results shown in [9], we provide a demonstrative technique involving the combinatorics of the collections of cells, like the combinatorial interpretation of the Krull dimension and the rook polynomial of a polyomino.
+The aim of the research is to study the main algebraic properties of $K[\mathcal{P}]$ depending on the geometry of $\mathcal{P}$.
+The most investigated include primality, Cohen-Macaulayness and Gorensteinness.
+As usual, we say that a collection of cells is prime (respectively, Cohen-Macaulay or Gorenstein) if the related binomial ideal is prime (respectively, Cohen-Macaulay or Gorenstein).
+Regarding the classification of the primality for polyominoes, it is proven in [19, 20, 31] that if $\mathcal{P}$ is simple, which means that $\mathcal{P}$ has no hole, then $K[\mathcal{P}]$ is a normal Cohen-Macaulay domain.
+Nowadays, the study is devoted to non-simple polyominoes but a complete characterization of primality, Cohen-Macaulay and Gorenstein properties is still unknown, despite the efforts of many mathematicians (see [2, 4, 6, 8, 19, 20, 22, 26, 27, 30, 31, 32, 33]).
+Other very interesting combinatorial problems on polyomino ideals have been investigated as shown in [9, 11, 14, 16, 24].
+Cohen-Macaulayness and Goresteiness have so far been discussed only for prime collections of cells.
+In all the papers where Cohen-Macaulayness is studied ([7, 22, 25, 27, 31]), the demonstrative strategy is based on two well-known theorems of Sturmfels ([36, Chapter 13]) and Hochster ([3, Theorem 6.3.5]), which state that a toric ring whose defining ideal admits a squarefree initial ideal is a normal Cohen-Macaulay ring.
+For Goresteiness, a well-known result of Stanley ([35]) allows to characterize the prime polyominoes which are Gorenstein, by checking if the $h$-polynomial is palindromic.
+In such a context, the so-called rook polynomial plays a crucial role.
 The rook polynomial is a polynomial whose $k$-th coefficient is the number of ways of placing $k$ non-attacking rooks in a polyomino.
 The maximum number of non-attacking rooks is called the rook number of $\mathcal{P}$.
-The connection of the Castelnuovo-Mumford regularity of $K[\mathcal{P}]$ with the rook number is shown in [14] for the class of $L$-convex polyominoes.
-In general, the study of the $h$-polynomial in terms of the rook polynomial is a very interesting issue and this research has produced a relevant number of articles ([7, 11, 22, 24, 29, 31]).
-The rook polynomial also plays a crucial role in the study of the Gorensteiness of $K[\mathcal{P}]$ when it is a domain.
-A well-known result of Stanley ([35]) allows us to characterize the Gorenstein coordinate rings which are domain, checking if the $h$-polynomial is palindromic, which means counting the number of suitable configurations of non-attacking rooks.
-It is hard to characterize non-prime polyominoes whose coordinate ring is Gorenstein and such a class has never appeared so far.
-Some partial results are discussed in our paper (look at Corollaries 2.5, 3.9 and Theorem 4.18) and we point out that an approach using the initial ideal is not useful (Remark 3.10).
-Finally, for other very interesting combinatorial problems on polyomino ideals, we invite the reader to look at [1, 8, 10, 13, 15, 23, 32].
+The connection of the Castelnuovo-Mumford regularity of $K[\mathcal{P}]$ with the rook number was shown in [15] for the first time.
+So far, the study of the $h$-polynomial in terms of the rook polynomial has produced a relevant number of articles, including [8, 12, 23, 25, 30, 32], and it could help to characterize the prime polyominoes that are Gorenstein.
+Indeed, if the $h$-polyonomial of a prime polyomino $\mathcal{P}$ is the rook polynomial, then we can check the palindromicity counting the number of suitable configurations of non-attacking rooks in $\mathcal{P}$.
+On the other hand, discussing non-prime polyominoes which are Gorenstein seems very hard and such classes have not appeared so far.
+In this context a natural question arises: what can we say about the Cohen-Macaulay and the Gorenstein properties for non-prime collections of cells?
+Motivated by this question, we begin to investigate these properties for non-prime collections of cells.
+From [26, Corollary 3.6] we know that if a collection of cells contains the so-called zig-zag walks (see [26, Definition 3.2]), then the related binomial ideal is not prime.
+This suggests to consider as a first step a class of non-prime collections of cells made up of zig-zag walks, which we called zig-zag collections.
+As conjectured in [26, Conjecture 4.6], every non-prime collection of cells should contain a zig-zag collection, so the latter could be useful for gaining information on non-prime collections of cells, by applying a decomposition of them into suitable zig-zag collections.
+This method is used to study the before mentioned properties for the closed paths, a kind of non-simple polyominoes introduced for the first time in [4], in the non-prime case.
+In particular, for this class of polyominoes, we state the following result at the end of this paper:
+Theorem 4.18 Let $\mathcal{P}$ be a closed path polyomino. Then:
+  1. (1)
+    $K[\mathcal{P}]$ is Cohen-Macaulay with Krull dimension $|V(\mathcal{P})|-|\mathcal{P}|$; moreover, if $\mathcal{P}$ does not contain any zig-zag walk, then $K[\mathcal{P}]$ is also a normal domain.
+  2. (2)
+    The $h$-polyonomial of $K[\mathcal{P}]$ is the rook polynomial of $\mathcal{P}$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
+  3. (3)
+    $K[\mathcal{P}]$ is Gorenstein if and only if $\mathcal{P}$ consists of maximal blocks of rank three.
 The article is structured as follows.
-Section 1 is devoted to provide some basic definitions and properties of collections of cells and the associated coordinate ring and to recall some useful known results of Commutative Algebra.
-In Section 2 we prove that the $h$-polynomial of a simple collection of cells is equal to the product of the $h$-polynomial of each connected component (Theorem 2.2), exploiting the monomial order in [27] and [5, Theorem 3.3].
-As a direct consequence, we obtain in Corollary 2.5 that the $h$-polynomial of a simple thin collection of cells coincides with the rook-polynomial.
+Section 1 is devoted to provide some basic definitions and properties of collections of cells and the associated coordinate rings.
+In Section 2 we prove that the $h$-polynomial of a simple collection of cells is equal to the product of the $h$-polynomial of each connected component (Theorem 2.2), exploiting the monomial order in [28] and [6, Theorem 3.3].
+As a direct consequence, we can generalize [26, Theorem 3.13], obtaining in Corollary 2.5 that the $h$-polynomial of a simple thin collection of cells coincides with the rook-polynomial.
 In Section 3 we introduce a new class of non-prime collections of cells, named zig-zag collections.
-We show that their coordinate rings are Cohen-Macaulay (Theorem 3.6), we compute the Hilbert-Poincaré series in terms of the switching rook polynomial (Corollary 3.8) and we eventually discuss the Gorenstein property (Corollary  3.9).
-The results achieved in these sections are used in Section 4 to prove our main result, which states that the coordinate ring of a non-prime closed path is Cohen-Macaulay (Theorem 4.13).
-In the end, Theorem 4.18 summarizes the results about the primality and the Cohen-Macaulayness of this kind of polyominoes and provides a complete characterization of the Gorenstein property in terms of the length of the maximal blocks.
-We conclude the paper with Section 5, giving some open questions for future works.
-1. Collections of cells and known results on graded $K$-algebras
+Taking advantage of some recent results shown in [10], we show that their coordinate rings are Cohen-Macaulay (Theorem 3.7), we compute the Hilbert-Poincaré series in terms of the switching rook polynomial (Corollary 3.8) and we eventually discuss the Gorenstein property (Corollary  3.9).
+The results achieved in the previous sections are collected in Section 4 to prove Theorem 4.18, which is our main result.
+1. Collections of cells and their associated coordinate rings
 This section is devoted to introduce several preliminaries.
@@
 A polyomino $\mathcal{P}$ is a non-empty, finite collection of cells in $\mathbb{Z}^{2}$ where any two cells of $\mathcal{P}$ are connected in $\mathcal{P}$ (see Figure 1 (a)) Let $\mathcal{P}$ be a non-empty, finite collection of cells in $\mathbb{Z}^{2}$, $\mathcal{P}$ is called weakly connected if for any two cells $C$ and $D$ in $\mathcal{P}$ there exists a sequence of cells $\mathcal{C}:C=C_{1},\dots,C_{m}=D$ of $\mathcal{P}$ such that $V(C_{i})\cap V(C_{i+1})\neq\emptyset$ for all $i=1,\dots,m-1$.
-A subset of cells $\mathcal{P}^{\prime}$ of $\mathcal{P}$ is called a connected component of $\mathcal{P}$ if $\mathcal{P}^{\prime}$ is a polyomino and it is maximal with respect to the set inclusion, that is if $A\in\mathcal{P}\setminus\mathcal{P}^{\prime}$ then $\mathcal{P}^{\prime}\cup\{A\}$ is not a polyomino.
+A subset of cells $\mathcal{P}^{\prime}$ of $\mathcal{P}$ is called a connected component of $\mathcal{P}$ if $\mathcal{P}^{\prime}$ is a polyomino and it is maximal with respect to the set inclusion, that is, if $A\in\mathcal{P}\setminus\mathcal{P}^{\prime}$ then $\mathcal{P}^{\prime}\cup\{A\}$ is not a polyomino.
 For instance, see Figure 1 (b).
 Observe trivially that every polyomino is a weakly connected collection of cells.
-(a) A polyomino.
-(b) A collection of cells.
-Figure 1. The collection of cells in the right has three connected components.
+Figure 1. A non-simple polyomino, a non-simple collection of cells, a simple thin polyomino and a simple thin collection of cells.
 A sub-polyomino of $\mathcal{P}$ is a polyomino whose cells belong to $\mathcal{P}$.
-A weakly connected collection of cells is said to be thin if it does not contain the square tetromino (that is, the square obtained as a union of four distinct cells).
+A weakly connected collection of cells is said to be thin if it does not contain the square tetromino (that is, the square obtained as a union of four distinct cells); see Figures 1c and 1d.
 We say that $\mathcal{P}$ is simple if for any two cells $C$ and $D$ not in $\mathcal{P}$ there exists a path of cells not in $\mathcal{P}$ from $C$ to $D$.
 A finite collection of cells $\mathcal{H}$ not in $\mathcal{P}$ is an hole of $\mathcal{P}$ if any two cells of $\mathcal{H}$ are connected in $\mathcal{H}$ and $\mathcal{H}$ is maximal with respect to set inclusion.
-For example, the collections of cells in Figure 1 are not simple.
+For example, the collections of cells in Figures 1a, 1b and 2 are not simple.
 Each hole of $\mathcal{P}$ is a simple polyomino, and $\mathcal{P}$ is simple if and only if it has no hole.
@@
     for all $i,j\in\{1,\dots,\ell\}$ with $i\neq j$, there exists no inner interval $J$ of $\mathcal{P}$ such that $z_{i}$, $z_{j}$ belong to $J$.
-Figure 2. An example of a zig-zag walk of $\mathcal{P}$.
+(a) An example of a zig-zag walk of $\mathcal{P}$.
+(b) An example of a $6$-rook configuration in $\mathcal{P}$.
+Figure 2. Non-simple polyominoes.
+See Figure 2a for an example.
 Consider two cells $A$ and $B$ of $\mathbb{Z}^{2}$ with $a=(i,j)$ and $b=(k,l)$ as the lower left corners of $A$ and $B$ with $a\leq b$.
@@
 A $k$-rook configuration in $\mathcal{P}$ is a configuration of $k$ rooks which are arranged in $\mathcal{P}$ in non-attacking positions.
-Figure 3 shows a 6-rook configuration.
-Figure 3. An example of a $6$-rook configuration in $\mathcal{P}$.
+Figure 2b shows a 6-rook configuration.
 The rook number $r(\mathcal{P})$ is the maximum number of rooks that can be arranged in $\mathcal{P}$ in non-attacking positions.
-For all $k\in\{0,\ldots,r(\mathcal{P})\}$, we denote by $\mathcal{R}_{k}$ the set of all $k$-rook configurations in $\mathcal{P}$, setting conventionally $R_{0}=\{\emptyset\}$.
+For all $k\in\{0,\ldots,r(\mathcal{P})\}$, we denote by $\mathcal{R}_{k}$ the set of all $k$-rook configurations in $\mathcal{P}$, using the conventional setting $R_{0}=\{\emptyset\}$.
 We also set $r_{k}=|\mathcal{R}_{k}|$ for all $k\in\{0,\dots,r(\mathcal{P})\}$.
@@
 The operation of replacing $R_{1}$ and $R_{2}$ by $R_{1}^{\prime}$ and $R_{2}^{\prime}$ is called switch of $R_{1}$ and $R_{2}$.
-This induce the following equivalence relation $\sim$ on $\mathcal{R}_{k}$: let $F_{1},F_{2}\in\mathcal{R}_{k}$, then $F_{1}\sim F_{2}$ if $F_{2}$ can be obtained from $F_{1}$ after some switches.
+This induces the following equivalence relation $\sim$ on $\mathcal{R}_{k}$: let $F_{1},F_{2}\in\mathcal{R}_{k}$, then $F_{1}\sim F_{2}$ if $F_{2}$ can be obtained from $F_{1}$ after some switches.
 We define the quotient set ~$\tilde{\mathcal{R}}_{k}=\mathcal{R}_{k}/\sim$.
@@
 We can continue these arguments until $\mathcal{P}_{n}$, getting that $\prod_{i=1}^{n}\tilde{r}_{i}(t)$ is the switching rook polynomial of $\mathcal{P}$.
-Assume that $\mathcal{P}$ consists of two connected components $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Denote by $\tilde{r}(t)$ the switching rook polynomial of $\mathcal{P}$ and we set
-$\tilde{r}_{1}(t)=\sum_{i=0}^{r(\mathcal{P}_{1})}r_{i}^{(1)}t^{i}\quad\text{and}\quad\tilde{r}_{2}(t)=\sum_{j=0}^{r(\mathcal{P}_{2})}r_{j}^{(2)}t^{j}.$
-Hence
+Assume that $\mathcal{P}$ consists of two connected components $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$.
+Denote by $\tilde{r}(t)$ the switching rook polynomial of $\mathcal{P}$ and we set $\tilde{r}_{1}(t)=\sum_{i=0}^{r(\mathcal{P}_{1})}r_{i}^{(1)}t^{i}$ and $\tilde{r}_{2}(t)=\sum_{j=0}^{r(\mathcal{P}_{2})}r_{j}^{(2)}t^{j}.$ Hence
 $\tilde{r}_{1}(t)\tilde{r}_{2}(t)=\Bigg{(}\sum_{i=0}^{r(\mathcal{P}_{1})}r_{i}^{(1)}t^{j}\Bigg{)}\Bigg{(}\sum_{j=0}^{r(\mathcal{P}_{2})}r_{j}^{(2)}t^{j}\Bigg{)}=\sum_{k=0}^{r(\mathcal{P}_{1})+r(\mathcal{P}_{2})}c_{k}t^{k}.$
@@
 For this reason, whenever the term “collection of cells” is introduced in the paper, we mean a weakly connected collection of cells.
-For utility, we state the following result about the Cohen-Macaulay property of the simple collection of cells, that in literature we found stated in this form only in the case $\mathcal{P}$ is polyomino (see [18, Theorem 2.1] and [19, Corollary 3.3]).
+For utility, we state the following result about the Cohen-Macaulay property of the simple collection of cells, that in the literature we found stated in this form only in the case $\mathcal{P}$ is polyomino (see [19, Theorem 2.1] and [20, Corollary 3.3]).
 For the sake of completeness, we also provide its proof.
@@
 Proof.
-By [5, Theorem 3.3] we know that $I_{\mathcal{P}}=J_{\mathcal{P}}$ where $J_{\mathcal{P}}$ is the toric ring of a bipartite weakly chordal graph.
+By [6, Theorem 3.3] we know that $I_{\mathcal{P}}=J_{\mathcal{P}}$ where $J_{\mathcal{P}}$ is the toric ring of a weakly chordal bipartite graph.
 In particular $I_{\mathcal{P}}$ is prime and, by [28], $I_{\mathcal{P}}$ has a squarefree quadratic Gröbner basis.
-A toric ring whose toric ideal admits a squarefree initial ideal is normal, see [16, Corollary 4.26] and by a theorem of Hochster (see [2, Theorem 6.3.5]), a normal toric ring is Cohen–Macaulay.
-Finally, by [28, Theorem 3.5] and [28, Corollary 3.6] we know that $I_{\mathcal{P}}=I_{\Lambda}$ for a saturated lattice $\Lambda$ such that $\operatorname{rank}_{\mathbb{Z}}(\Lambda)=|\mathcal{P}|$.
-Hence $\operatorname{height}(I_{\mathcal{P}})=|\mathcal{P}|$ (see [16, Problem 3.12]) and so $\dim K[\mathcal{P}]=|V(\mathcal{P})|-|\mathcal{P}|$ (see also [37, Corollary 3.1.7]).
+A toric ring whose toric ideal admits a squarefree initial ideal is normal, see [17, Corollary 4.26] and by a theorem of Hochster (see [3, Theorem 6.3.5]), a normal toric ring is Cohen–Macaulay.
+Finally, by [29, Theorem 3.5] and [29, Corollary 3.6] we know that $I_{\mathcal{P}}=I_{\Lambda}$ for a saturated lattice $\Lambda$ such that $\operatorname{rank}_{\mathbb{Z}}(\Lambda)=|\mathcal{P}|$.
+Hence $\operatorname{height}(I_{\mathcal{P}})=|\mathcal{P}|$ (see [17, Problem 3.12]) and so $\dim K[\mathcal{P}]=|V(\mathcal{P})|-|\mathcal{P}|$ (see also [37, Corollary 3.1.7]).
 ∎
@@
     For all $i\in\{1,\dots,n\}$ and for all $j\notin\{i-2,i-1,i,i+1,i+2\}$ then $V(A_{i})\cap V(A_{j})=\emptyset$, where $A_{-1}=A_{n-1}$, $A_{0}=A_{n}$, $A_{n+1}=A_{1}$ and $A_{n+2}=A_{2}$.
-Figure 4. A closed path with two $L$-configurations and a ladder of four steps.
+Figure 3. Examples of closed paths.
 A path of five cells $C_{1},C_{2},C_{3},C_{4},C_{5}$ of $\mathcal{P}$ is called an L-configuration if the two sequences $C_{1},C_{2},C_{3}$ and $C_{3},C_{4},C_{5}$ go in two orthogonal directions.
 A set =$\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1,\dots,n}$ of maximal horizontal (or vertical) blocks of rank at least two, with $V(\mathcal{B}_{i})\cap V(\mathcal{B}_{i+1})=\{a_{i},b_{i}\}$ and $a_{i}\neq b_{i}$ for all $i=1,\dots,n-1$, is called a ladder of $n$ steps if $[a_{i},b_{i}]$ is not on the same edge interval of $[a_{i+1},b_{i+1}]$ for all $i=1,\dots,n-2$.
-For instance, in Figure 4 there is a closed path having two $L$-configurations and a ladder of four steps.
-It is proved in [3, Section 6] that a closed path has no zig-zag walks if and only if it contains an $L$-configuration or a ladder of at least three steps.
-Furthermore, in [3, Theorem 6.2] the authors characterize the primality of a closed path $\mathcal{P}$ proving that $I_{\mathcal{P}}$ is prime if and only if $\mathcal{P}$ does not contain any zig-zag walk.
-1.4. Some results in Commutative Algebra.
-In this subsection we recall some well-known results in Commutative Algebra that we use extensively throughout the paper.
-We recall, for completeness, that $\operatorname{depth}(M)\leq\dim(M)$ holds for every $R$-module $M$ over a Noetherian ring $R$, and $M$ is Cohen-Macaulay when equality holds.
+For instance, in Figure 3a there is a closed path having two $L$-configurations and a ladder of four steps.
+It is proved in [4, Section 6] that a closed path has no zig-zag walks if and only if it contains an $L$-configuration or a ladder of at least three steps.
+Furthermore, in [4, Theorem 6.2] the authors characterize the primality of a closed path $\mathcal{P}$ proving that $I_{\mathcal{P}}$ is prime if and only if $\mathcal{P}$ does not contain any zig-zag walk.
+2. Hilbert-Poincaré series of simple collections of cells
+In this section we present some results on simple collections of cells.
+From [7, Theorem 3.3] and [28] we know that, if $\mathcal{P}$ is a simple collection of cells, the set of generators of $I_{\mathcal{P}}$ forms the reduced Gröbner basis of $I_{\mathcal{P}}$ with respect to a suitable monomial order.
+The monomial order provided in [28] is fundamental for our purposes because it allows us to write $S_{\mathcal{P}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}})$ as $K$-tensor product of suitable $K$-algebras coming from the connected components.
 About some notations we use throughout the paper, if $S$ is a polynomial ring, $\prec$ is a monomial order on it and $f\in S$ is a polynomial, we denote by $\operatorname{in}_{\prec}(f)$, or $\operatorname{in}(f)$ if the order is clear in the context, the initial monomial of $f$ with respect to $\prec$.
 As usual, if $S=K[x_{1},\ldots,x_{n}]$ and $m$ is a monomial, then $\operatorname{supp}(m)=\{x_{i}\in S\mid x_{i}\text{ divides }m\}$.
-We start by providing some basics on the Hilbert function and the Hilbert-Poincaré series of a graded $K$-algebra $R/I$.
-Let $R$ be a graded $K$-algebra and $I$ be an homogeneous ideal of $R$.
-Then $R/I$ has a natural structure of graded $K$-algebra as $\bigoplus_{k\in\mathbb{N}}(R/I)_{k}$.
-The numerical function $\mathrm{H}_{R/I}:\mathbb{N}\to\mathbb{N}$ with $\mathrm{H}_{R/I}(k)=\dim_{K}(R/I)_{k}$ is called the Hilbert function of $R/I$.
-The formal power series $\mathrm{HP}_{R/I}(t)=\sum_{k\in\mathbb{N}}\mathrm{H}_{R/I}(k)t^{k}$ is called the Hilbert-Poincaré series of $R/I$.
-It is known by Hilbert-Serre theorem that there exists a polynomial $h(t)\in\mathbb{Z}[t]$ with $h(1)\neq 0$ such that $\mathrm{HP}_{R/I}(t)=\frac{h(t)}{(1-t)^{d}}$, where $d$ is the Krull dimension of $R/I$.
-The polynomial $h(t)$ is called h-polynomial of $R/I$.
-Moreover, if $R/I$ is Cohen-Macaulay then $\mathrm{reg}(R/I)=\deg h(t)$.
-Recall also that if $S=K[x_{1},\ldots,x_{n}]$ then $\mathrm{HP}_{S}(t)=\frac{1}{(1-t)^{n}}$.
-In some parts of the paper, if $K[\mathcal{P}]$ is the coordinate ring of a collection of cells, we use the notation $h_{K[\mathcal{P}]}$ for the $h$-polyonimial of the Hilbert-Poincaré series of $K[\mathcal{P}]$.
-Proposition 1.5.
-If $I$ is a homogeneous ideal of a graded ring $R$ then:
-  1. (1)
-    $R/I$ and $R/\mathrm{in}(I)$ have the same Hilbert function for all monomial orders (see [37, Corollary 3.3.15]);
-  2. (2)
-    If $R/\mathrm{in}(I)$ is Cohen–Macaulay (resp. Gorenstein), then $R/I$ is Cohen–Macaulay (resp. Gorenstein) (see [37, Proposition 9.6.17]).
-From [9] we deduce that the converse for the Cohen-Macaulayness in (2) of Proposition 1.5 holds for a homogeneous ideal whose initial ideal is squarefree.
-Proposition 1.6 ([9]).
-Let $I$ be a homogeneous ideal in a polynomial ring $S$ and suppose $\mathrm{in}(I)$ is radical with respect to some term order.
-Then $\mathrm{reg}(S/I)=\mathrm{reg}(S/\mathrm{in}(I))$ and $\mathrm{depth}(S/I)=\mathrm{depth}(S/\mathrm{in}(I))$.
-Furthermore, $S/I$ is Cohen-Macaulay if and only if $S/\mathrm{in}(I)$ is Cohen-Macaulay.
-For the following result, see Lemma 5.1.2 and Lemma 5.1.3 of [37].
-Proposition 1.7.
-Let $R$ be a graded $K$-algebra and $I$ be a graded ideal of $R$. Let $q$ be an homogeneous element of $R$ of degree $m$ and let
-$0\longrightarrow R/(I:q)\longrightarrow R/I\longrightarrow R/(I,q)\longrightarrow 0$
-be a short exact sequence. Then $\mathrm{HP}_{R/I}(t)=\mathrm{HP}_{R/(I,q)}(t)+t^{m}\mathrm{HP}_{R/(I:q)}(t)$.
-For the following result, see [37, Lemma 5.1.11].
-Proposition 1.8.
-Let $A$ and $B$ be standard graded $K$-algebras over a field $K$.
-Then $\mathrm{HP}_{A\otimes_{K}B}(t)=\mathrm{HP}_{A}(t)\cdot\mathrm{HP}_{B}(t)$.
-In particular $\mathrm{dim}(A\otimes_{K}B)=\mathrm{dim}(A)+\mathrm{dim}(B)$.
-For the following result, see Lemma 3.1.34, Lemma 3.1.35 and Proposition 11.5.8 of [37].
-Proposition 1.9.
-Let $A$ and $B$ be standard graded $K$-algebras over a field $K$.
-Then $\mathrm{depth}(A\otimes_{K}B)=\mathrm{depth}(A)+\mathrm{depth}(B)$.
-In particular $A\otimes_{K}B$ is Cohen-Macaulay (Gorenstein) if and only if $A$ and $B$ are Cohen-Macaulay (Gorenstein).
-For the following result, see [2, Proposition 1.2.9].
-Proposition 1.10.
-Let $0\longrightarrow N_{1}\longrightarrow N_{2}\longrightarrow N_{3}\longrightarrow 0$ be a short exact sequence of finitely generated $R$-module over a Noetherian ring $R$. Then:
-  • $\mathrm{depth}(N_{2})\geq\min\{\mathrm{depth}(N_{3}),\mathrm{depth}(N_{1})\}$.
-  • $\mathrm{depth}(N_{1})\geq\min\{\mathrm{depth}(N_{2}),\mathrm{depth}(N_{3})+1\}$.
-  • $\mathrm{depth}(N_{3})\geq\min\{\mathrm{depth}(N_{1})-1,\mathrm{depth}(N_{2})\}$.
-2. Hilbert-Poincaré series of simple collections of cells
-In this section we present some results on simple collections of cells.
-From [6, Theorem 3.3] and [27] we know that, if $\mathcal{P}$ is a simple collection of cells, the set of generators of $I_{\mathcal{P}}$ forms the reduced Gröbner basis of $I_{\mathcal{P}}$ with respect to a suitable monomial order.
-The monomial order provided in [27] is fundamental for our purposes because it allows us to write $S_{\mathcal{P}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}})$ as $K$-tensor product of suitable $K$-algebras coming from the connected components.
 We start proving the following Lemma.
@@
 ∎
+Let $R$ be a graded $K$-algebra and $I$ be a homogeneous ideal of $R$.
+The formal power series $\mathrm{HP}_{R/I}(t)=\sum_{k\in\mathbb{N}}\dim_{K}(R/I)_{k}t^{k}$ is called the Hilbert-Poincaré series of $R/I$.
+It is known by Hilbert-Serre theorem that there exists a polynomial $h(t)\in\mathbb{Z}[t]$ with $h(1)\neq 0$ such that $\mathrm{HP}_{R/I}(t)=\frac{h(t)}{(1-t)^{d}}$, where $d$ is the Krull dimension of $R/I$.
+The polynomial $h(t)$ is called h-polynomial of $R/I$.
+Moreover, if $R/I$ is Cohen-Macaulay then $\mathrm{reg}(R/I)=\deg h(t)$.
+Recall also that if $S=K[x_{1},\ldots,x_{n}]$ then $\mathrm{HP}_{S}(t)=\frac{1}{(1-t)^{n}}$.
+In some parts of the paper, if $K[\mathcal{P}]$ is the coordinate ring of a collection of cells, we use the notation $h_{K[\mathcal{P}]}$ for the $h$-polyonimial of the Hilbert-Poincaré series of $K[\mathcal{P}]$.
+Once we set the previous notation, let us prove the main result of this section.
 Theorem 2.2.
@@
 Let $\mathcal{P}$ be a simple collection of cells.
-According to the references [5, Theorem 3.3] and [27], $I_{\mathcal{P}}$ possesses a quadratic Gröbner basis with respect to a certain monomial order.
+According to references [6, Theorem 3.3] and [28], $I_{\mathcal{P}}$ possesses a quadratic Gröbner basis with respect to a certain monomial order.
 We fix such a monomial order, then for any pair of inner 2-minors $f$ and $g$, their $S$-polynomial $S(f,g)$ reduces to 0.
@@
 Observe that for all $i\in[n]$ we have $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})=S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})\otimes_{K}K[x_{v}\mid v\in D_{\mathcal{P}_{i}}]$.
-In particular, by Proposition 1.8, $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=(1-t)^{d_{i}}\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)$, and since $\mathcal{P}_{i}$ is a simple polyomino, by Proposition 1.3 we have $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=\frac{h_{\mathcal{P}_{i}}(t)}{(1-t)^{|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-d_{i}}}$.
+In particular, by [37, Lemma 5.1.11], $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=(1-t)^{d_{i}}\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)$, and since $\mathcal{P}_{i}$ is a simple polyomino, by Proposition 1.3 we have $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=\frac{h_{\mathcal{P}_{i}}(t)}{(1-t)^{|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-d_{i}}}$.
 Note that $\sum_{i=1}^{n}|V(\mathcal{P}_{i})|=|V(\mathcal{P})|+|A_{\mathcal{P}}|$ and $\sum_{i=1}^{n}|\mathcal{P}_{i}|=|\mathcal{P}|$.
-Therefore, by Proposition 1.8, we have:
+Therefore, by [37, Lemma 5.1.11], we have:
 $\mathrm{HP}_{K[\mathcal{P}]}(t)=\frac{\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|+|A_{\mathcal{P}}|+|C_{\mathcal{P}}|-\sum_{i=1}^{n}d_{i}}}$
@@
 By the previous result, we argue that, if $\mathcal{P}$ is a simple collection of cells, the study of the $h$-polynomial of $K[\mathcal{P}]$ in terms of the switching rook polynomial of $\mathcal{P}$ can be reduced to investigate the Hilbert-Poincaré series of its connected components.
-In particular, if for all simple polyominoes [29, Conjecture 3.2] is true, then the same holds for all simple collections of cells.
+In particular, if for all simple polyominoes [30, Conjecture 3.2] is true, then the same holds for all simple collections of cells.
 Corollary 2.3.
@@
 Let $\mathcal{P}$ be a simple collection of cells where every connected component $\mathcal{P}_{i}$ is a parallelogram polyomino.
-Then, by [29, Theorem 3.5], $h_{K[\mathcal{P}_{i}]}(t)$ is the switching rook polynomial of $\mathcal{P}_{i}$ for all $i\in[n]$, so $h_{K[\mathcal{P}]}$ is the switching rook polynomial of $\mathcal{P}$ by Corollary 2.3.
-As a consequence, we generalize [31, Theorem 3.13 and Theorem 4.2] to simple thin collections of cells (see Figure 6 for an example). This is useful in the next sections.
+Then, by [30, Theorem 3.5], $h_{K[\mathcal{P}_{i}]}(t)$ is the switching rook polynomial of $\mathcal{P}_{i}$ for all $i\in[n]$, so $h_{K[\mathcal{P}]}$ is the switching rook polynomial of $\mathcal{P}$ by Corollary 2.3.
+As a consequence, we generalize [32, Theorem 3.13 and Theorem 4.2] to simple thin collections of cells (see Figure 1c and 1d for an example). This is useful in the next sections.
 We recall the definition of S-property given in [32].
@@
 $\mathcal{P}$ has the S-property if every maximal interval of $\mathcal{P}$ has only one single cell.
-Figure 6. A simple thin polyomino and a simple thin collection of cells.
 Corollary 2.5.
@@
 Proof.
-As a consequence of Corollary 2.3 and [31, Theorem 3.13], we get that $h_{K[\mathcal{P}]}$ is the rook polynomial of $\mathcal{P}$.
+As a consequence of Corollary 2.3 and [32, Theorem 3.13], we get that $h_{K[\mathcal{P}]}$ is the rook polynomial of $\mathcal{P}$.
 Moreover, $\operatorname{reg}(K[\mathcal{P}])=r(\mathcal{P})$ since $K[\mathcal{P}]$ is a Cohen-Macaulay ring.
-Now we prove the result about the Gorensteiness.
+Now we prove the result about Gorensteinness.
 Let $h_{K[\mathcal{P}]}(t)=r_{\mathcal{P}}(t)=\sum_{k=0}^{s}r_{k}t^{k}$, where $s=r(\mathcal{P})$.
-Since $K[\mathcal{P}]$ is a Cohen-Macaulay domain, it is known from [35] that $K[\mathcal{P}]$ is Gorenstein if and only if $r_{i}=r_{s-i}$ for all $i=0,\dots,s$, that is $h_{K[\mathcal{P}]}(t)$ is palindromic.
+Since $K[\mathcal{P}]$ is a Cohen-Macaulay domain, it is known from [35] that $K[\mathcal{P}]$ is Gorenstein if and only if $r_{i}=r_{s-i}$ for all $i=0,\dots,s$, that is, $h_{K[\mathcal{P}]}(t)$ is palindromic.
 Assume that $\mathcal{P}$ consists of $n$ connected components and denote by $h_{K[\mathcal{P}_{i}]}(t)$ the $h$-polynomial of $K[\mathcal{P}_{i}]$ for all $i\in[n]$.
@@
 Since $\mathcal{P}$ has the $S$-property, then each connected component $\mathcal{P}_{i}$ has the $S$-property as well.
-By [31, Theorem 4.2], $h_{K[\mathcal{P}_{i}]}(t)$ is palindromic for all $i\in[n]$, so $h_{K[\mathcal{P}]}(t)$ is palindromic as well, since $h_{K[\mathcal{P}]}(t)=\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}(t)$ and the product of palindromic polynomials is palindromic.
+By [32, Theorem 4.2], $h_{K[\mathcal{P}_{i}]}(t)$ is palindromic for all $i\in[n]$, so $h_{K[\mathcal{P}]}(t)$ is palindromic as well, since $h_{K[\mathcal{P}]}(t)=\prod_{i=1}^{n}h_{K[\mathcal{P}_{i}]}(t)$ and the product of palindromic polynomials is palindromic.
 Hence it follows by [35] that $K[\mathcal{P}]$ is Gorenstein.
@@
 For all $i\in[n]$, let $h_{K[\mathcal{P}_{i}]}(t)=\sum_{j=0}^{s_{i}}r_{j}^{(i)}t^{j}$ be the $h$-polynomial of $K[\mathcal{P}_{i}]$, where $s_{i}=r(\mathcal{P}_{i})$ and observe that $r_{1}^{(i)}=|\mathcal{P}_{i}|$.
-Looking at the proof of case (b)$\Rightarrow$(c) of [31, Theorem 4.2], for all $k\in[n]$ such that $\mathcal{P}_{k}$ has not the $S$-property we have $r_{s_{k}}>1$ or $r_{s_{k}-1}>|\mathcal{P}_{k}|$.
-Therefore, it follows easily that $r_{s}>1$ or $r_{s-1}>\sum_{i=1}^{n}|\mathcal{P}_{i}|=|\mathcal{P}|$, that is a contradiction, since $h_{K[\mathcal{P}]}(t)$ is palindromic.
+Looking at the proof of case (b)$\Rightarrow$(c) of [32, Theorem 4.2], for all $k\in[n]$ such that $\mathcal{P}_{k}$ has not the $S$-property we have $r_{s_{k}}>1$ or $r_{s_{k}-1}>|\mathcal{P}_{k}|$.
+Therefore, it easily follows that $r_{s}>1$ or $r_{s-1}>\sum_{i=1}^{n}|\mathcal{P}_{i}|=|\mathcal{P}|$, that is a contradiction, since $h_{K[\mathcal{P}]}(t)$ is palindromic.
 ∎
@@
 (b) A non zig-zag collection.
-Figure 7. The collection of cells in the right is not a zig-zag collection but contains a zig-zag walk.
+Figure 5. Collections of cells containing a zig-zag walk.
 Remark 3.2.
@@
 Moreover, $I_{\mathcal{P}}$ is not prime because $I_{1},\dots,I_{\ell}$ is a zig-zag walk of $\mathcal{P}$.
-For more details we refer to [25, Section 3], where the arguments given for polyominoes can be extended easily to collections of cells.
-Our first step is to study the Gröbner basis of the inner 2-minors ideal of a zig-zag collection.
+For more details we refer to [26, Section 3], where the arguments given for polyominoes can be extended easily to collections of cells.
+Our first step is to study the Gröbner basis of the inner 2-minor ideal of a zig-zag collection.
 For this aim, we introduce four total orders on $\mathbb{Z}^{2}$.
@@
 For $i\in\{1,2,3,4\}$, denote by $<_{\mathrm{lex}}^{(i)}$ the lexicographic order on $K[x_{v}\mid v\in V(\mathcal{P})]$, induced by the following total order on the variables: $x_{a}<_{\mathrm{lex}}^{(i)}x_{b}$ if and only if $a<^{(i)}b$ for $a,b\in V(\mathcal{P})$.
-If $f$ is an inner 2-minor in $I_{\mathcal{P}}$, observe that for $i\in\{1,3\}$ the initial monomial of $f$ with respect to $<_{\mathrm{lex}}^{(i)}$ is related to the anti-diagonal corners of the inner interval associated.
+If $f$ is an inner 2-minor in $I_{\mathcal{P}}$, observe that for $i\in\{1,3\}$ the initial monomial of $f$ with respect to $<_{\mathrm{lex}}^{(i)}$ is related to the anti-diagonal corners of the associated inner interval.
 While for $i\in\{2,4\}$ the initial monomial of $f$ is related to the diagonal corners.
-Moreover, [28, Theorem 4.1] and the arguments of its proof hold also considering the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{2,4\}$.
-The same can be considered for [28, Remark 4.2] with respect to the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{1,3\}$.
-Similar considerations can be made using the reverse lexicographic order, with reference to [26, Proposition 3.2].
-For our aim, we state the result below, which can be proved by the same arguments of the mentioned results in [28] or also using the lemmas in [6, Section 3].
+Moreover, [29, Theorem 4.1] and the arguments of its proof hold also considering the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{2,4\}$.
+The same can be considered for [29, Remark 4.2] with respect to the monomial orders $<_{\mathrm{lex}}^{(i)}$ for $i\in\{1,3\}$.
+Similar considerations can be made using the reverse lexicographic order, with reference to [27, Proposition 3.2].
+For our aim, we state the result below, which can be proved by the same arguments of the mentioned results in [29] or also using the lemmas in [7, Section 3].
 Proposition 3.3.
@@
 In the first step we give a total order on $(I_{1}\setminus\{v_{1}\})\cup I_{2}$.
-We use $<^{(3)}$ and $<^{(2)}$ respectively to order the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$; look at Figure 9b, where the arrows denote the sense of the orders of the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$ respectively.
+We use $<^{(3)}$ and $<^{(2)}$, respectively, to order the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$; look at Figure 7b, where the arrows denote the sense of the orders of the vertices in $I_{1}\setminus\{v_{1}\}$ and $I_{2}\setminus\{v_{2}\}$ respectively.
 Moreover, we set that every vertex of $\mathcal{P}$ in $I_{1}\setminus\{v_{1}\}$ is smaller than each one in $I_{2}\setminus\{v_{2}\}$.
@@
   1. (1)
-    If $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 10a and 10d or as in Figures 11a, 11b and 11d, then we use $<^{(1)}$ and $<^{(2)}$ to order the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ respectively.
+    If $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 8a and 8d or as in Figures 9a, 9b and 9d, then we use $<^{(1)}$ and $<^{(2)}$ to order the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ respectively.
   2. (2)
-    In the case that $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are as in Figure 10b, then we define an order of the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ using $<^{(3)}$ and $<^{(2)}$ respectively.
+    In the case that $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are as in Figure 8b, then we define an order of the vertices in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$ using $<^{(3)}$ and $<^{(2)}$, respectively.
   3. (3)
-    When $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are in the position of Figure 10c, then use respectively $<^{(1)}$ and $<^{(4)}$ to make order in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$.
+    When $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are in the position of Figure 8c, then use respectively $<^{(1)}$ and $<^{(4)}$ to make an order in $I_{2k-1}\setminus\{v_{2k-1}\}$ and $I_{2k}\setminus\{v_{2k}\}$.
   4. (4)
@@
 We prove that $<_{\mathrm{lex}}$ provides the desired claim.
-We start proving (1).
+We start by proving (1).
 Let $f=x_{a}x_{b}-x_{c}x_{d}$ and $g=x_{p}x_{q}-x_{r}x_{s}$ be two generators of $I_{\mathcal{P}}$ attached to the inner intervals $[a,b]$ and $[p,q]$ of $\mathcal{P}$, respectively.
@@
 Let $k\in[\ell/2]$.
-We may consider just the case when $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 10b (with the convention that $I_{0}=I_{\ell}$); indeed, all the other situations described in Figures 10 and 11 can be proved similarly.
+We may consider just the case when $I_{2k-2}$, $I_{2k-1}$ and $I_{2k}$ are arranged as in Figures 8b (with the convention that $I_{0}=I_{\ell}$); indeed, all the other situations described in Figures 8 and 9 can be proved similarly.
 If $[a,b],[p,q]\subseteq I_{2k-1}$ or $[a,b],[p,q]\subseteq I_{2k}$ then $S(f,g)$ reduces to 00 with respect to $G(\mathcal{P})$ by Proposition 3.3.
@@
 In particular, the claim (2) follows from the considerations done for claim (1).
-Finally, it is known that if the initial ideal of a homogeneous ideal $I$ of $K[x_{1},\dots,x_{n}]$, with respect to a monomial order, is squarefree then $I$ is radical (see [17, Corollary 2.2]).
+Finally, it is known that if the initial ideal of a homogeneous ideal $I$ of $K[x_{1},\dots,x_{n}]$, with respect to a monomial order, is squarefree then $I$ is radical (see [18, Corollary 2.2]).
 Since $\mathop{\rm in}\nolimits_{<_{\mathrm{lex}}}(I_{\mathcal{P}})$ is squarefree, then $I_{\mathcal{P}}$ is radical.
@@
 We highlight that, by the proof of Lemma 3.4, the order $\prec$ is the lexicographic order induced by a total order on $V(\mathcal{P})$ and it satisfies the following property: if $a\in\mathcal{E}_{i}$ and $b\in\mathcal{E}_{j}$ with $i<j$ then $x_{a}\prec x_{b}$.
-This fact is useful further.
-For shortness we will write $\mathrm{in}(I_{\mathcal{P}})$ and $\mathrm{in}(I_{\mathcal{P}_{i}})$ for $\mathrm{in}_{\prec}(I_{\mathcal{P}})$ and $\mathrm{in}_{<_{i}}(I_{\mathcal{P}_{i}})$ respectively.
+This observation will be useful in later arguments.
+For shortness, we will write $\mathrm{in}(I_{\mathcal{P}})$ and $\mathrm{in}(I_{\mathcal{P}_{i}})$ for $\mathrm{in}_{\prec}(I_{\mathcal{P}})$ and $\mathrm{in}_{<_{i}}(I_{\mathcal{P}_{i}})$ respectively.
 Moreover, following the construction in Lemma 3.4, without loss of generality we can always assume that $v_{1}$ is an anti-diagonal corner in $I_{1}$.
@@
 Set also $\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)=\frac{h_{i}(t)}{(1-t)^{n_{i}}}$, in particular we know that $n_{i}=|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$.
+Finally, before proving the main result of this section, we recall a beautiful result achieved in [10], which will be crucial also in the other parts of this work.
+Proposition 3.6 ([10]).
+Let $I$ be a homogeneous ideal in a polynomial ring $S$ and suppose that $\mathrm{in}(I)$ is radical with respect to some term order.
+Then $\mathrm{reg}(S/I)=\mathrm{reg}(S/\mathrm{in}(I))$ and $\mathrm{depth}(S/I)=\mathrm{depth}(S/\mathrm{in}(I))$.
+Furthermore, $S/I$ is Cohen-Macaulay if and only if $S/\mathrm{in}(I)$ is Cohen-Macaulay.
 Theorem 3.7.
@@
 So we argue that $\mathrm{HP}_{S^{\prime}_{\mathcal{P}_{i}}/\mathop{\rm in}\nolimits(I_{\mathcal{P}_{i}})}(t)=(1-t)\mathrm{HP}_{K[\mathcal{P}_{i}]}(t)$ and we obtain our claim on $\mathrm{HP}_{K[\mathcal{P}]}(t)$.
-(2) By [28, Theorem 2.2] we know that $K[\mathcal{P}_{i}]$ is a Cohen-Macaulay domain of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$, so by Proposition 1.6 also $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$.
-Therefore, by Proposition 1.9 we obtain that $S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1$ and as consequence $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Cohen-Macaulay of dimension $\sum_{i=1}^{\ell}(|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1)=|V(\mathcal{P})|-|\mathcal{P}|$.
+(2) By [29, Theorem 2.2] we know that $K[\mathcal{P}_{i}]$ is a Cohen-Macaulay domain of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$, so by Proposition 3.6 also $S_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|$.
+Therefore, by [37, Lemma 3.1.35] we obtain that $S^{\prime}_{\mathcal{P}_{i}}/\mathrm{in}(I_{\mathcal{P}_{i}})$ is Cohen-Macaulay of dimension $|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1$ and, as a consequence, $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Cohen-Macaulay of dimension $\sum_{i=1}^{\ell}(|V(\mathcal{P}_{i})|-|\mathcal{P}_{i}|-1)=|V(\mathcal{P})|-|\mathcal{P}|$.
 By Proposition 3.6 the same property holds for $K[\mathcal{P}]$.
 ∎
-Example 3.7.
-We provide here an example of a non zig-zag collection of cells, whose coordinate ring is Cohen-Macaulay. Let $\mathcal{P}$ be the collection of cells in Figure 12.
-Figure 12. A not zig-zag collection of cells
-Denote by $<$ the lexicographic order on $S_{\mathcal{P}}$ induced by the following total order of the variables:
-$x_{34}>x_{42}>x_{45}>x_{44}>x_{54}>x_{73}>x_{61}>x_{65}>x_{57}>x_{47}>x_{37}>x_{16}>x_{15}>x_{14}>x_{22}>x_{72}>$
-$>x_{66}>x_{64}>x_{63}>x_{62}>x_{56}>x_{55}>x_{53}>x_{52}>x_{51}>x_{46}>x_{43}>x_{36}>x_{33}>x_{32}>x_{27}>$
-$>x_{26}>x_{25}>x_{24}>x_{23}>x_{13}$
-By some computations we get that the set of the generators of $I_{\mathcal{P}}$ forms the reduced Gröbner basis of $I_{\mathcal{P}}$ with respect to $<$ and the initial ideal of $I_{\mathcal{P}}$ is generated by the monomials:
-$x_{42}x_{54},x_{45}x_{54},x_{34}x_{63},x_{44}x_{63},x_{54}x_{63},x_{73}x_{62},x_{65}x_{56},x_{42}x_{55},x_{34}x_{53},x_{42}x_{53},x_{45}x_{53},x_{44}x_{53},x_{61}x_{52},x_{57}x_{46},$
-$x_{34}x_{43},x_{57}x_{36},x_{47}x_{36},x_{22}x_{33},x_{57}x_{26},x_{47}x_{26},x_{37}x_{26},x_{16}x_{25},x_{16}x_{24},x_{15}x_{24},x_{16}x_{23},x_{15}x_{23},x_{14}x_{23}.$
-Hence $I_{\mathcal{P}}$ is radical but it is not prime since $[(2,2),(3,3)],[(3,3),(6,4)],[(6,2),(7,3)],[(5,1),(6,2)],$ $[(4,2),(5,5)],[(5,5),(6,6)],[(2,6),(5,7)],[(1,3),(2,6)]$ is a zig-zag walk of $\mathcal{P}$.
-Using Macaulay2 ([4, 12]) we check that $S_{\mathcal{P}}/\mathrm{in}_{<}(I_{\mathcal{P}})$ is a Cohen-Macaulay ring.
-Hence, from Proposition 1.6, we have that $K[\mathcal{P}]$ is Cohen-Macaulay.
-We point out that we needed to check the Cohen-Macaulayness of $S_{\mathcal{P}}/\mathrm{in}_{<}(I_{\mathcal{P}})$ and not of $S_{\mathcal{P}}/I_{\mathcal{P}}$ using Macaulay2, because the time of computation to check the same property for $S_{\mathcal{P}}/I_{\mathcal{P}}$ is too long.
 Corollary 3.8.
@@
 Proof.
-For $i\in[\ell]$ we denote by $h_{i}(t)$ the $h$-polynomial of $K[\mathcal{P}_{i}]$, so from [29, Theorem 3.5] we have that $h_{i}(t)$ is the switching rook-polynomial of $\mathcal{P}_{i}$ for all $i\in[\ell]$.
+For $i\in[\ell]$ we denote by $h_{i}(t)$ the $h$-polynomial of $K[\mathcal{P}_{i}]$, so from [30, Theorem 3.5] we have that $h_{i}(t)$ is the switching rook-polynomial of $\mathcal{P}_{i}$ for all $i\in[\ell]$.
 Recall that Theorem 3.7 states in particular that $h(t)=\prod_{i=1}^{\ell}h_{i}(t)$.
@@
 Denote by $\mathcal{C}_{i}$ a canonical configuration of $r(\mathcal{P}_{i})$ rooks in $\mathcal{P}_{i}$, for all $i\in[s]\setminus\{h\}$, and set $\mathcal{C}=\cup_{i\in[s]\setminus\{h\}}\mathcal{C}_{i}$.
-It is easy to see that $\mathcal{C}\cup\mathcal{T}_{1}$ and $\mathcal{C}\cup\mathcal{T}_{2}$ are two canonical configuration of $r(\mathcal{P})$ rooks in $\mathcal{P}$, so $h_{s}>1$.
+It is easy to see that $\mathcal{C}\cup\mathcal{T}_{1}$ and $\mathcal{C}\cup\mathcal{T}_{2}$ are two canonical configurations of $r(\mathcal{P})$ rooks in $\mathcal{P}$, so $h_{s}>1$.
 This is a contradiction because $h_{s}=1$.
@@
 (2) Observe that $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ can be viewed as an edge ring of a graph $G$ on the vertex set $V(\mathcal{P})$, where $\{a,b\}\in E(G)$ if either $a,b$ are the anti-diagonal corners of an inner interval of $I_{2k}$ or they are the diagonal corners of an inner interval of $I_{2k-1}$, for $k\in[\ell/2]$.
-If $\mathcal{P}_{i}$ is a cell for all $i\in[\ell]$, $G$ is a disjoint union of edges, so $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Gorenstein by [20, Corollary 9.3.3].
-Therefore, it follows that $K[\mathcal{P}]$ is Gorenstein by (2) of Proposition 1.5.
+If $\mathcal{P}_{i}$ is a cell for all $i\in[\ell]$, $G$ is a disjoint union of edges, so $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ is Gorenstein by [21, Corollary 9.3.3].
+Therefore, it follows that $K[\mathcal{P}]$ is Gorenstein by [37, Proposition 9.6.17].
 ∎
@@
 If there exists $j\in[\ell]$ such that $\mathcal{P}_{j}$ is a square different from a cell, then $S_{\mathcal{P}}/\mathrm{in}(I_{\mathcal{P}})$ cannot be Gorenstein because $G$ is a chordal graph but it is not a disjoint union of edges.
-Anyway, we tested different examples with Macaulay2 that shows $K[\mathcal{P}]$ is still Gorenstein.
+Anyway, we tested different examples with Macaulay2 ([5, 13]) that shows $K[\mathcal{P}]$ is still Gorenstein.
 So, we ask if the converse of (1) is true in general.
-4. Cohen-Macaulay property and Hilbert series for closed path polyominoes having zig-zag walks
-From now onward, we assume that $\mathcal{P}$ is a non-prime closed path polyomino.
-By [3, Theorem 6.2], $\mathcal{P}$ contains a zig-zag walk.
-See Figure 13 for an example of a closed path with zig-zag walks.
-Figure 13. Closed path with zig-zag walks.
-{tikzpicture}
-{tikzpicture}
+4. Cohen-Macaulay property and Hilbert-Poincaré series of closed paths with zig-zag walks
+In this section we investigate Cohen-Macaulayness and the Hilbert-Poincaré series of non-prime closed path polyominoes.
+From now onward, we assume that $\mathcal{P}$ is a non-prime closed path.
+By [4, Theorem 6.2], $\mathcal{P}$ contains a zig-zag walk.
+See Figure 3b for an example of a closed path with zig-zag walks.
+Discussion 4.1.
+If $\mathcal{P}$ is a closed path polyomino having zig-zag walks, by [4, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration or ladders with at least three steps.
+So, any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 10a or 10b (up to reflections or rotations).
+$m$$X$
+$W$$X$$Y$$Z$
 Figure 10. Possible “changes of direction” in a closed path having a zig-zag walk.
-Discussion 4.1.
-If $\mathcal{P}$ is a closed path polyomino having zig-zag walks, by [3, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration or ladders with at least three steps.
-So, any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 14a or 14b (up to reflections or rotations).
 Furthermore, by [7] we know that there exists a monomial order such that the Gröbner basis is quadratic and squarefree.
-The monomial order that provides this property can be built following [6, Algorithm 4.1].
+The monomial order that provides this property can be built following [7, Algorithm 4.1].
 Following this construction, we obtain that for all inner intervals containing the vertex $m$ as in Figure 10a, (up to reflection and rotation), the initial term of the related inner 2-minors contain the variable $x_{m}$ in its support.
-From now onward we denote by $M_{\mathcal{P}}$ the set of all vertices $m$ in $\mathcal{P}$ as in Figure 14a, up to rotations and reflections, and by $C_{m}$ the cell labeled with $X$ with reference to the same figure.
+From now onward, we denote by $M_{\mathcal{P}}$ the set of all vertices $m$ in $\mathcal{P}$ as in Figure 10a, up to rotations and reflections, and by $C_{m}$ the cell labeled with $X$ with reference to the same figure.
 Set-up 4.2.
@@
 Remark 4.3.
-Let $\mathcal{P}$ be a closed path with a zig-zag walk and $\prec$ be the monomial order on $S_{\mathcal{P}}$ defined by [6, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$.
+Let $\mathcal{P}$ be a closed path with a zig-zag walk and $\prec$ be the monomial order on $S_{\mathcal{P}}$ defined by [7, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$.
 For all $i\in[r]$, with reference to Set-up 4.2, since $V(\mathcal{Q}_{i})\subset V(\mathcal{P})$ and $V(\mathcal{Q}_{i}^{\prime})\subset V(\mathcal{P})$, the monomial order $\prec$ can be considered also in $S_{\mathcal{Q}_{i}}$ and $S_{\mathcal{Q}^{\prime}_{i}}$.
 In particular, $\prec$ trivially preserves the same initial terms in the ideals $I_{\mathcal{P}}$, $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}^{\prime}_{i}}$.
-Furthermore, by [7, Proposition 4.1], $\prec$ provides a quadratic squarefree Gröbner basis also for $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}_{i}^{\prime}}$.
+Furthermore, by [8, Proposition 4.1], $\prec$ provides a quadratic squarefree Gröbner basis also for $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}_{i}^{\prime}}$.
 So, with abuse of notation, we refer to $\mathrm{in}(I_{\mathcal{Q}_{i}})$, $\mathrm{in}(I_{\mathcal{Q}_{i-1}})$ and $\mathrm{in}(I_{\mathcal{Q}^{\prime}_{i}})$ considering the monomial order $\prec$.
@@
 Proof.
-Let $\prec$ be the monomial order in $S_{\mathcal{P}}$ defined by [6, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$.
+Let $\prec$ be the monomial order in $S_{\mathcal{P}}$ defined by [7, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$.
 For all $i\in[r]$, we focus on the initial ideals of $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}_{i}^{\prime}}$ as explained in Remark 4.3.
@@
 $0\longrightarrow\frac{S_{\mathcal{Q}_{i-1}}}{(\mathrm{in}(I_{\mathcal{Q}_{i-1}}):x_{m_{i}})}\longrightarrow\frac{S_{\mathcal{Q}_{i-1}}}{\mathrm{in}(I_{\mathcal{Q}_{i-1}})}\longrightarrow\frac{S_{\mathcal{Q}_{i-1}}}{(\mathrm{in}(I_{\mathcal{Q}_{i-1}}),x_{m_{i}})}\longrightarrow 0$
-Arguing as in the proof of [7, Proposition 4.3], we can prove that:
+Arguing as in the proof of [8, Proposition 4.3], we can prove that:
   • $\frac{S_{\mathcal{Q}_{i-1}}}{(\mathrm{in}(I_{\mathcal{Q}_{i-1}}),x_{m_{i}})}\cong\frac{S_{\mathcal{Q}_{i}}}{\mathrm{in}(I_{\mathcal{Q}_{i}})}$
   • $\frac{S_{\mathcal{Q}_{i-1}}}{(\mathrm{in}(I_{\mathcal{Q}_{i-1}}):x_{m_{i}})}\cong\frac{S_{\mathcal{Q}^{\prime}_{i}}}{\mathrm{in}(I_{\mathcal{Q}^{\prime}_{i}})}\otimes K[x_{m_{i}}]$.
-So by Propositions 1.5 and 1.7 it follows that for all $i\in\{1\,\ldots,r\}$ it is verified:
+So by [37, Corollary 3.3.15] and [37, Lemmas 5.1.2 and 5.1.3] it follows that for all $i\in\{1\,\ldots,r\}$ it is verified:
 $\mathrm{HP}_{K[\mathcal{Q}_{i-1}]}(t)=\mathrm{HP}_{K[\mathcal{Q}_{i}]}(t)+\frac{t}{1-t}\mathrm{HP}_{K[\mathcal{Q}^{\prime}_{i}]}(t)$
-and also, by Propositions 1.6, 1.9 and 1.10, we obtain $\mathrm{depth}(K[\mathcal{Q}_{i-1}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{i}]),\mathrm{depth}(K[\mathcal{Q}_{i}^{\prime}])+1\}$.
+and also, by Propositions 3.6, [37, Lemma 3.1.34] and [3, Proposition 1.2.9], we obtain $\mathrm{depth}(K[\mathcal{Q}_{i-1}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{i}]),\mathrm{depth}(K[\mathcal{Q}_{i}^{\prime}])+1\}$.
 In particular, we obtain the first equality of $\mathrm{HP}_{K[\mathcal{P}]}(t)$.
@@
 If $\mathcal{P}$ is a non-prime closed path polyomino not containing the configuration in Figure 10b then $\mathcal{Q}_{r}$ is a zig-zag collection and by Theorem 3.7 we have that $K[\mathcal{Q}_{r}]$ is Cohen-Macaulay of dimension $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{P})|-|\mathcal{P}|$.
-Hence $\mathrm{depth}(K[\mathcal{P}])\geq|V(\mathcal{P})|-|\mathcal{P}|$ but we also know that $\mathrm{dim}(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$ (see [10, Theorem 3.4]), so $K[\mathcal{P}]$ is Cohen-Macaulay of dimension $|V(\mathcal{P})|-|\mathcal{P}|$.
+Hence $\mathrm{depth}(K[\mathcal{P}])\geq|V(\mathcal{P})|-|\mathcal{P}|$ but we also know that $\mathrm{dim}(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$ (see [11, Theorem 3.4]), so $K[\mathcal{P}]$ is Cohen-Macaulay of dimension $|V(\mathcal{P})|-|\mathcal{P}|$.
 ∎
@@
 That is, we can express $\mathcal{Q}_{r}=\bigcup_{k=1}^{q}\mathcal{C}_{k}$, where $\mathcal{C}_{k}$ is a configuration as in Figure 12a or Figure 12b or Figure 12c (up to rotations and reflections), for all $k\in\{1,\ldots,q\}$, for some positive integer $q$.
-In particular, with reference to the highlighted points in the figures, we build $\mathcal{Q}_{r}$ by gluing those configurations with this procedure: setting the first configuration $\mathcal{C}_{k}$, the second one, say $\mathcal{C}_{l}$ with $l=k+1$, is glued considering an opportune reflection or rotation of it, in such a way that the vertices $c_{k},d_{k}$ and $e_{k}$ in orange of the first configuration are overlapped, respectively, to the vertices $c_{l-1},d_{l-1}$ and $e_{l-1}$ in blue of the second configuration.
+In particular, with reference to the points highlighted in the figures, we build $\mathcal{Q}_{r}$ by gluing those configurations with this procedure: setting the first configuration $\mathcal{C}_{k}$, the second one, say $\mathcal{C}_{l}$ with $l=k+1$, is glued considering an opportune reflection or rotation of it, in such a way that the vertices $c_{k},d_{k}$ and $e_{k}$ in orange of the first configuration are overlapped, respectively, to the vertices $c_{l-1},d_{l-1}$ and $e_{l-1}$ in blue of the second configuration.
 See Figure 13 for an example of this construction.
@@
     $\mathcal{R}^{\prime}_{i,k}=\mathcal{R}_{i,k}\setminus\{C\in\mathcal{Q}_{r}\mid C\ \mbox{belongs to a maximal cell interval containing}\ B_{i,k}\}.$
-In what follows, we often refer to the framework explained in Set-up 4.8, and we made extensive use of the notations introduced there, also with the help of figures.
+In what follows, we often refer to the framework explained in Set-up 4.8, and we made extensive use of the notation introduced there, also with the help of figures.
 Example 4.9.
@@
 Observe that $\mathcal{R}_{n_{s},s}$ is a zig-zag collection, so we denote by $I_{1},\ldots,I_{p}$ the set of maximal intervals of such a zig-zag collection, for some integer $p$.
-We can assume, without loss of generality, that $I_{1}$ is related to the cell interval $[E_{1},B_{1,1}]\setminus\{B_{1,1}\}$, with reference to the notations explained in Set-up 4.8 and to Figures 16a and 16b, assuming also that $I_{1}$ have the same orientation of these figures.
+We can assume, without loss of generality, that $I_{1}$ is related to the cell interval $[E_{1},B_{1,1}]\setminus\{B_{1,1}\}$, with reference to the notations explained in Set-up 4.8 and to Figures 12a and 12b, assuming also that $I_{1}$ have the same orientation of these figures.
 By Lemma 3.4, we know that there exists a term order $\prec_{\mathcal{R}_{n_{s},s}}$ for which $I_{\mathcal{R}_{n_{s},s}}$ has a quadratic Gröbner basis and satisfies:
@@
 Trivially, the monomial order $\prec_{\mathcal{Q}_{r}}$ satisfies claim (2).
-We want to prove also that claim (1) holds.
+We also want to prove that claim (1) holds.
 Firstly, we prove it for the ideal $I_{\mathcal{Q}_{r}}$.
@@
 Hence the inner interval related to $f$ contains the cell $B_{i,k}$ for some $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$.
-Then, referring to Figure 19, up to reflection an rotations $f$ is one of the following:
+Then, referring to Figure 15, up to reflection and rotations, $f$ is one of the following:
   • (a)
@@
     $f=f_{j,i,k}$ with $j\in[\ell]$ and $g=\pm(x_{v_{j_{1}}}x_{h_{j_{2}}}-x_{v_{j_{2}}}x_{h_{j_{1}}})$ with $j_{1},j_{2}\in[\ell]$ or $g=f_{l,i,k}$ with $l\in[\ell]\setminus\{j\}$.
-In this case $S(f,g)$ reduces to 00 by [6, Remark 1].
+In this case $S(f,g)$ reduces to 00 by [7, Remark 1].
   • Case 2)
-    $f=f_{1,i,k}$ and $g=f_{b,i,k}$. Also in this case $S(f,g)$ reduces to 00 by [6, Remark 1].
+    $f=f_{1,i,k}$ and $g=f_{b,i,k}$. In this case also, $S(f,g)$ reduces to 00 by [7, Remark 1].
   • Case 3)
-    $f=f_{j,i,k}$ with $j\in[\ell]\setminus\{1\}$ and $g=f_{b,i,k}$. In this case, by [6, Lemma 2 (condition 1)], we obtain that $S(f,g)$ reduces to 00 if both the following properties holds:
+    $f=f_{j,i,k}$ with $j\in[\ell]\setminus\{1\}$ and $g=f_{b,i,k}$. In this case, by [7, Lemma 2 (condition 1)], we obtain that $S(f,g)$ reduces to 00 if both the following properties holds:
     • 3a)
@@
 In this case, recall that $x_{z}\prec_{\mathcal{Q}_{r}}x_{b_{i,k}^{(2)}}$ for all $z\notin Y$.
-Hence, if $u\notin Y$ or $u\in Y$ with $x_{u}\prec_{\mathcal{Q}_{r}}x_{b_{i,k}^{(2)}}$ then $S(f,g)$ reduces to 00 by [6, Lemma 4 (condition 3)], while if $u\in Y$ and $x_{b_{i,k}^{(2)}}\prec_{\mathcal{Q}_{r}}x_{u}$ then $S(f,g)$ reduces to 00 by [6, Lemma 4 (condition 1)].
+Hence, if $u\notin Y$ or $u\in Y$ with $x_{u}\prec_{\mathcal{Q}_{r}}x_{b_{i,k}^{(2)}}$ then $S(f,g)$ reduces to 00 by [7, Lemma 4 (condition 3)], while if $u\in Y$ and $x_{b_{i,k}^{(2)}}\prec_{\mathcal{Q}_{r}}x_{u}$ then $S(f,g)$ reduces to 00 by [7, Lemma 4 (condition 1)].
   • Case 5)
@@
   • $\operatorname{in}(I_{\mathcal{P}})=\operatorname{in}(I_{\mathcal{R}})+(x_{a}x_{c})+\sum_{i=1}^{r}(x_{a}x_{b_{i}})+\max_{\prec}(\{x_{a_{i}}x_{b_{j}},x_{a_{j}}x_{b_{i}}\mid x_{a_{i}}x_{b_{j}}-x_{a_{j}}x_{b_{i}}\in I_{\mathcal{P}},i,j\in[r],i<j\})+\max_{\prec}(\{x_{c}x_{z_{j}},x_{b_{1}}x_{w_{j}}\mid x_{c}x_{z_{j}}-x_{b_{1}}x_{w_{j}}\in I_{\mathcal{P}}\}).$
-As consequence, we obtain:
+As a consequence, we obtain:
   • $(\operatorname{in}(I_{\mathcal{P}}),x_{a})=(\operatorname{in}(I_{\mathcal{Q}}),x_{a})$ and in particular $\frac{S_{\mathcal{P}}}{(\operatorname{in}(I_{\mathcal{P}}),x_{a})}\cong\frac{S_{\mathcal{Q}}}{\operatorname{in}(I_{\mathcal{Q}})}$.
@@
 Without loss of generality, assume $z$ is in the same horizontal edge interval of $b_{r}$, while $w$ is in the same vertical edge interval of $b_{r}$.
-As consequence, assuming the $S$-polynomial of $f$ and $g$ reduces to zero, by [6, Lemma 6] we have $[z,a_{r}]$ or $[w,a]$ is an inner interval, a contradiction.
+As a consequence, assuming that the $S$-polynomial of $f$ and $g$ reduces to zero, by [7, Lemma 6] we have $[z,a_{r}]$ or $[w,a]$ is an inner interval, a contradiction.
 Therefore, if $y$ is another indeterminate and denoting by $S^{\prime}$ the polynomial ring such that $S^{\prime}\otimes K[x_{b_{r}}]\cong S_{\mathcal{R}}$, we can express
 $\frac{S_{\mathcal{R}}}{\operatorname{in}(I_{\mathcal{R}})}\cong\frac{S^{\prime}}{\operatorname{in}(I_{\mathcal{R}})}\otimes K[x_{b_{r}}]\cong\frac{S_{\mathcal{R}}}{(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})}\otimes K[y]$
-Hence, by Proposition 1.8, we have $\mathrm{HP}_{S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})}(t)=(1-t)\cdot\mathrm{HP}_{S_{\mathcal{R}}/\operatorname{in}(I_{\mathcal{R}})}(t)$.
-Therefore, by Proposition 1.5 and Proposition 1.7 we obtain $\mathrm{HP}_{K[\mathcal{P}]}(t)=\mathrm{HP}_{K[\mathcal{R}]}(t)+\frac{t}{(1-t)^{r}}\mathrm{HP}_{K[\mathcal{R}]}(t)$, that is the first claim.
-For the second claim, by Proposition 1.10 we have $\operatorname{depth}(S_{\mathcal{P}}/\operatorname{in}(I_{\mathcal{P}}))\geq\{\operatorname{depth}(S_{\mathcal{Q}}/\operatorname{in}(I_{\mathcal{Q}})),\operatorname{depth}(S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})\otimes K[x_{a},x_{a_{1}},\ldots,x_{a_{r}}])\}$.
-Hence, by Propositions 1.6 and 1.9 we obtain $\operatorname{depth}(K[\mathcal{P}])\geq\min\{\operatorname{depth}(K[\mathcal{Q}]),\operatorname{depth}(K[\mathcal{R}])+r\}$.
+Hence, by [37, Lemma 5.1.11], we have $\mathrm{HP}_{S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})}(t)=(1-t)\cdot\mathrm{HP}_{S_{\mathcal{R}}/\operatorname{in}(I_{\mathcal{R}})}(t)$.
+Therefore, by [37, Corollary 3.3.15] and [37, Lemmas 5.1.2 and 5.1.3] we obtain $\mathrm{HP}_{K[\mathcal{P}]}(t)=\mathrm{HP}_{K[\mathcal{R}]}(t)+\frac{t}{(1-t)^{r}}\mathrm{HP}_{K[\mathcal{R}]}(t)$, that is the first claim.
+For the second claim, by [3, Proposition 1.2.9] we have $\operatorname{depth}(S_{\mathcal{P}}/\operatorname{in}(I_{\mathcal{P}}))\geq\{\operatorname{depth}(S_{\mathcal{Q}}/\operatorname{in}(I_{\mathcal{Q}})),\operatorname{depth}(S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})\otimes K[x_{a},x_{a_{1}},\ldots,x_{a_{r}}])\}$.
+Hence, by Propositions 3.6 and [37, Lemma 3.1.34] we obtain $\operatorname{depth}(K[\mathcal{P}])\geq\min\{\operatorname{depth}(K[\mathcal{Q}]),\operatorname{depth}(K[\mathcal{R}])+r\}$.
 Finally, for the last two claims, if $\mathcal{R}$ is a simple collection of cells, by Proposition 1.3 we have that $\dim(K[\mathcal{R}])=\operatorname{depth}(K[\mathcal{R}])=|V(\mathcal{R})|-|\mathcal{R}|$.
@@
 and $\operatorname{depth}(K[\mathcal{S}_{i,k}])\geq\min\{\operatorname{depth}(K[\mathcal{R}_{i,k}]),|V(\mathcal{S}_{i,k})|-|\mathcal{S}_{i,k}|\}$.
-So, denoting $J_{\mathcal{Q}_{r}}=\{k_{1},\ldots,k_{s}\}$, starting from $\mathcal{Q}_{r}$ we obtain:
+So, denote $J_{\mathcal{Q}_{r}}=\{k_{1},\ldots,k_{s}\}$.
+Starting from $\mathcal{Q}_{r}$ we obtain:
 $\mathrm{HP}_{K[\mathcal{Q}_{r}]}(t)$ $=\mathrm{HP}_{K[\mathcal{R}_{1,k_{1}}]}(t)+\frac{t\cdot h_{K[\mathcal{R}_{1,k_{1}}^{\prime}]}(t)}{(1-t)^{|V(\mathcal{Q}_{r})-|\mathcal{Q}_{r}|}}=$
@@
 $\operatorname{depth}(K[\mathcal{S}_{n_{k_{s}},k_{s}}])\geq$ $\min\{\operatorname{depth}(K[\mathcal{R}_{n_{k_{s}},k_{s}}]),|V(\mathcal{S}_{n_{k_{s}},k_{s}})|-|\mathcal{S}_{n_{k_{s}},k_{s}}|\}.$
-Observing that $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{R}_{i,k})|-|\mathcal{R}_{i,k}|$ for all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$ (in fact $\mathcal{S}_{i,k}$ differs from $\mathcal{R}_{i,k}$ only for a vertex and cells), we obtain $\operatorname{depth}(K[\mathcal{Q}_{r}])\geq\min\{\operatorname{depth}(K[\mathcal{R}_{n_{k_{s}},k_{s}}]),|V(Q_{r})|-|\mathcal{Q}_{r}|\}$.
+Observing that $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{R}_{i,k})|-|\mathcal{R}_{i,k}|$ for all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$ (in fact, $\mathcal{S}_{i,k}$ differs from $\mathcal{R}_{i,k}$ only for a vertex and cells), we obtain $\operatorname{depth}(K[\mathcal{Q}_{r}])\geq\min\{\operatorname{depth}(K[\mathcal{R}_{n_{k_{s}},k_{s}}]),|V(Q_{r})|-|\mathcal{Q}_{r}|\}$.
 Finally, observe that $\mathcal{R}_{n_{k_{s}},k_{s}}$ is a zig-zag collection.
@@
 By Lemma 4.5 we have $\mathrm{depth}(K[\mathcal{P}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{r}]),|V(\mathcal{P})|-|\mathcal{P}|\}$ and by Lemma 4.12 $\mathrm{depth}(K[\mathcal{Q}_{r}])\geq|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$.
-Moreover, by the proof of Lemma 4.5 we have also $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{P})|-|\mathcal{P}|$.
+Moreover, by the proof of Lemma 4.5 we also have $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{P})|-|\mathcal{P}|$.
 Hence, $\operatorname{depth}(K[\mathcal{P}])\geq|V(\mathcal{P})|-|\mathcal{P}|$.
-By [10, Theorem 3.4] we also know that $\dim(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$, so $K[\mathcal{P}]$ is Cohen-Macaulay with the given dimension.
+By [11, Theorem 3.4] we also know that $\dim(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$, so $K[\mathcal{P}]$ is Cohen-Macaulay with the given dimension.
 ∎
@@
 Observe that the result in Theorem 4.13 holds also for all collection of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ described in Set-up 4.2 and Set-up 4.8.
-In fact, if $\mathcal{R}$ is one of such configurations, by [3, Lemma 3.3] and by construction, we can argue that there exists a cell $A$ in $\mathcal{R}$ such that $\mathcal{R}\setminus\{A\}$ is simple.
-So, by [8, Proposition 4.9] (and [37, Corollary 3.1.7]) we have $\dim(K[\mathcal{R}])=|V(\mathcal{R})|-|\mathcal{R}|$.
+In fact, if $\mathcal{R}$ is one of such configurations, by [4, Lemma 3.3] and by construction, we can argue that there exists a cell $A$ in $\mathcal{R}$ such that $\mathcal{R}\setminus\{A\}$ is simple.
+So, by [9, Proposition 4.9] (and [37, Corollary 3.1.7]) we have $\dim(K[\mathcal{R}])=|V(\mathcal{R})|-|\mathcal{R}|$.
 Therefore, we can obtain this claim using the same arguments developed in this section.
 Theorem 4.15.
-Let $\mathcal{P}$ a closed path with a zig-zag walk. Then the $h$-polynomial of $K[\mathcal{P}]$ is the rook polynomial of $\mathcal{P}$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
+Let $\mathcal{P}$ be a closed path with a zig-zag walk. Then the $h$-polynomial of $K[\mathcal{P}]$ is the rook polynomial of $\mathcal{P}$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
 Proof.
@@
 Let $s_{1}=\max J_{\mathcal{Q}_{r}}$ and $s_{2}=\min J_{\mathcal{Q}_{r}}$.
-Following the collections of cells involved, observe that, $\mathcal{R}_{n_{s},s}$ is a zig-zag collection and and $\mathcal{S}_{1,s_{1}}=\mathcal{Q}_{r}$.
+Following the collections of cells involved, observe that $\mathcal{R}_{n_{s},s}$ is a zig-zag collection and and $\mathcal{S}_{1,s_{1}}=\mathcal{Q}_{r}$.
 In particular, by Corollary 3.8, $h_{K[\mathcal{R}_{n_{s},s}]}(t)$ is the rook polynomial of $\mathcal{R}_{n_{s},s}$.
@@
 Therefore, $r_{1}=r_{k}^{(1)}$ and $r_{2}=r_{k-1}^{(2)}$, obtaining that $r_{k}$ is the number of $k$-rook configurations of $\mathcal{S}_{i,k}$.
-As consequence, $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{S}_{i,k}$, proving our claim.
+As a consequence, $h_{K[\mathcal{R}_{i,k}]}(t)$ is the rook polynomial of $\mathcal{S}_{i,k}$, proving our claim.
 So, the $h$-polynomial of $K[\mathcal{Q}_{r}]$ is the rook polynomial of $\mathcal{Q}_{r}$.
@@
 Following the collections of cells involved and observing that $\mathcal{Q}_{0}=\mathcal{P}$, we can obtain our result if for all $i\in[r]$ we prove the following claim: if $h_{K[\mathcal{Q}_{i}]}(t)$ is the rook polynomial of $\mathcal{Q}_{i}$ then $h_{K[\mathcal{Q}_{i-1}]}(t)$ is the rook polynomial of $\mathcal{Q}_{i-1}$.
-But it is not difficult to see that this fact can be proved by the same argument used before, so we can conclude.
+However, it is not difficult to see that this fact can be proved by the same argument used before, so we can conclude.
 ∎
@@
   3. (3)
-    From the previous cases it follows that every maximal interval of $\mathcal{P}$ has rank less or equal to $3$ and $\mathcal{P}$ does not contain any sub-polyomino as in Figure 14b.
-Hence the only possibility is that $\mathcal{P}$ is as in Figure 21.
-In such a case, Macaulay2 computations show that the reduced Hilbert series of $\mathcal{P}$ is given by
-    $\frac{1+16\,t+100\,t^{2}+308\,t^{3}+486\,t^{4}+376\,t^{5}+134\,t^{6}+20\,t^{7}+t^{8}}{\left(1-t\right)^{16}}$
-    so $h_{7}=20\neq 16=h_{1}$, which is a contradiction.
-    Figure 21. A closed path.
+    From the previous cases it follows that every maximal interval of $\mathcal{P}$ has a rank less than or equal to $3$ and $\mathcal{P}$ does not contain any sub-polyomino as in Figure 10b.
+Hence the only possibility is that $\mathcal{P}$ is as in Figure 3c.
+In such a case, Macaulay2 ([5, 13]) computations show that $h_{K[\mathcal{P}]}(t)=1+16\,t+100\,t^{2}+308\,t^{3}+486\,t^{4}+376\,t^{5}+134\,t^{6}+20\,t^{7}+t^{8},$ so $h_{7}=20\neq 16=h_{1}$, which is a contradiction.
 In all three cases we get a contradiction, so we conclude that $K[\mathcal{P}]$ cannot be Gorenstein. ∎
@@
   2. (2)
-    Let
-    $\mathrm{HP}_{K[\mathcal{P}]}(t)=\frac{h(t)}{(1-t)^{|V(\mathcal{P})|-|\mathcal{P}|}}$
-    be the reduced Hilbert series of $K[\mathcal{P}]$. Then $h(t)$ is the rook polynomial of $K[\mathcal{P}]$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
+    The $h$-polyonomial of $K[\mathcal{P}]$ is the rook polynomial of $\mathcal{P}$ and $\mathrm{reg}(K[\mathcal{P}])=r(\mathcal{P})$.
   3. (3)
@@
 Concerning (3), we know that $K[\mathcal{P}]$ is not Gorenstein from Proposition 4.17.
-Moreover, by [3, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration, so any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 14a or 14b (up to reflections or rotations).
+Moreover, by [4, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration, so any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 10a or 10b (up to reflections or rotations).
 Hence, we argue that if $\mathcal{P}$ has a zig-zag walk then it contains maximal blocks of length two.
@@
 Now, in the second case suppose that $\mathcal{P}$ does not contain a zig-zag walk.
-The claims (1) and (2) are already proved in [7, Theorem 5.5] and (3) in [7, Theorem 5.7].
+The claims (1) and (2) are already proved in [8, Theorem 5.5] and (3) in [8, Theorem 5.7].
 In conclusion, the theorem is completely proved.
 ∎
-5. Some open questions
-We conclude this paper with some considerations and insights for future research.
-About the Cohen-Macaulayness of the inner $2$-minor ideals of collections of cells, it is known that if $\mathcal{P}$ is a simple collection of cells then $K[\mathcal{P}]$ is Cohen-Macaulay (see Proposition 1.3).
-Other classes of non-simple polyominoes whose coordinate ring is Cohen-Macaulay are discussed in [6], [21] and [26], although in these cases the examined classes consist just of prime polyominoes.
-In that cases, it is sufficient to study the Gröbner basis of the polyomino ideal by two well-known theorems of Sturmfels ([36]) and of Hochster ([2]).
-One of the main goal of this work is to investigate some classes of non-simple collections of cells whose coordinate ring is not a domain but it is Cohen-Macaulay.
-In particular, we proved that the zig-zag collections and the closed paths with zig-zag walks are Cohen-Macaulay, as well as the collections of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ introduced respectively in Set-up 4.2 and Set-up 4.8.
-Also, the coordinate ring of the collection of cells in Example 3.7 is Cohen-Macaulay.
+We proved that the zig-zag collections and the closed paths with zig-zag walks are Cohen-Macaulay, as well as the collections of cells $\mathcal{Q}_{i}$ and $\mathcal{R}_{i,k}$ introduced respectively in Set-up 4.2 and Set-up 4.8.
 Nowadays, an example of a collection of cells having a not Cohen-Macaulay coordinate ring is still unknown.
-Therefore, the following general question naturally arises.
-Question 5.1.
-Let $\mathcal{P}$ be a collection of cells. Then $K[\mathcal{P}]$ is Cohen-Macaulay?
-If $\mathcal{P}$ is a zig-zag collection we proved in Corollary 3.9(1) a necessary condition for $K[\mathcal{P}]$ to be Gorenstein.
-Motivated by some examples tested in Macaulay2, we ask if that condition is also sufficient, as mentioned in Remark 3.10.
-So, we formulate the following question:
-Question 5.2.
-Let $\mathcal{P}$ be a zig-zag collection supported by the intervals $I_{1},\dots,I_{\ell}$.
-Let $\mathcal{P}_{i}$ be the collection of cells related to $I_{i}$ and suppose that $\mathcal{P}_{i}$ is a square for all $i\in[\ell]$.
-Then, is $K[\mathcal{P}]$ Gorenstein?
-Finally, in Theorem 4.18 we provide for the class of closed path polyominoes a description of the main algebraic properties of the related coordinate ring, characterizing them combinatorially.
-These properties are: be a domain, Hilbert-Poincaré series, regularity, be Cohen-Macaulay and Gorenstein.
-Recently in [32], the properties of Levelness and Pseudo-Gorenstein have been investigated in polyomino ideals for the first time, in particular for the class of path polyominoes (that are simple polyominoes).
-So another possible question is the following:
-Question 5.3.
-Let $\mathcal{P}$ be a closed path polyomino. Is it possible to characterize, by the combinatorics of $\mathcal{P}$, when $K[\mathcal{P}]$ is Level and Pseudo-Gorenstein?
+Therefore, from [19, 20, 31] and from the results of this work, the following general question naturally arises.
+Question 4.19.
+Let $\mathcal{P}$ be a collection of cells. Then, is $K[\mathcal{P}]$ Cohen-Macaulay?
+Acknowledgement.
+The third author is supported by Scientific and Technological Research Council of Turkey TÜBİTAK under the Grant No: 122F128, and he is thankful to TÜBİTAK for their support.
+The first and the third author acknowledge support of INDAM-GNSAGA.
+The second author is supported by Scientific and Technological Research Council of Turkey TÜBİTAK under the Grant No: 118F169, and he is thankful to TÜBİTAK for their support.
+The second author also acknowledges the support of Erasmus funding during his stay at the University of Messina.
+Special thanks to Professor Rosanna Utano for her support and hospitality during this period.